### v2

# Understanding the Challenges and Promises of Developing Generative AI Apps: An Empirical Study

Abstract.
The release of ChatGPT in 2022 triggered a rapid surge in generative artificial intelligence mobile apps (i.e., Gen-AI apps). Despite widespread adoption, little is known about how end users perceive and evaluate these Gen-AI functionalities in practice. In this work, we conduct a user-centered analysis of 676,066 reviews from 173 Gen-AI apps on the Google Play Store. We introduce a four-phase methodology, SARA (Selection, Acquisition, Refinement, and Analysis), that enables the systematic extraction of user insights using prompt-based LLM techniques. First, we demonstrate the reliability of LLMs in topic extraction, achieving 91% accuracy through five-shot prompting and non-informative review filtering. Then, we apply this method to the informative reviews, identify the top 10 user-discussed topics (e.g., AI Performance, Content Quality, and Content Policy & Censorship) and analyze the key challenges and emerging opportunities. Finally, we examine how these topics evolve over time, offering insight into shifting user expectations and engagement patterns with Gen-AI apps. Based on our findings and observations, we present actionable implications for developers and researchers.
Generative AI, Large Language Models, User Review Analysis, Mobile Apps
ccs: Software and its engineering Software maintenance tools
1. Introduction
Mobile applications (apps) are a crucial aspect of people’s lives, and their global data usage continues to rise each year (Rathod and Agal, 2023). A major shift in the mobile app landscape occurred with the release of ChatGPT [^1] https://play.google.com/store/apps/details?id=com.openai.chatgpt&hl=en, which achieved 7.4 million downloads within its first ten days. This rapid adoption signalled strong consumer interest in generative artificial intelligence (Gen-AI) and sparked a surge of similar artificial intelligence (AI) powered apps across app marketplaces. In 2024, this trend contributed to a 112% year-over-year increase in downloads of AI chatbot apps (Ceci, 2025), such as Microsoft Copilot [^2] https://play.google.com/store/apps/details?id=com.microsoft.copilot and DeepSeek [^3] https://play.google.com/store/apps/details?id=com.deepseek.chat. Gen-AI refers to a class of AI that generates new and meaningful content, including text, images, audio and video (Feuerriegel et al., 2024; Zhang et al., 2024). Unlike traditional AI models, which rely on historical data to perform predictions and analytical tasks, Gen-AI models synthesize information from diverse sources to create original outputs (Gozalo-Brizuela and Garrido-Merchán, 2023).
Among the most prominent forms of Gen-AI are large language models (LLMs), i.e., deep learning models trained on massive text corpora to understand and generate human-like language (Hadi et al., 2023). LLMs power a wide range of natural language processing (NLP) tasks, including text generation, summarization, translation, and dialogue. Developers are increasingly integrating Gen-AI models into mobile apps to enable novel capabilities, such as conversational interfaces, voice interaction, and content generation (Hau et al., 2025). When embedded into apps, these models can deliver personalized and adaptive experiences by providing real-time, context-aware feedback that enhances user engagement and motivation (Yuen and Schlote, 2024). A key enabler of this integration is the growing availability of Gen-AI-powered application programming interfaces (APIs). APIs serve as software intermediaries that enable mobile apps to access language understanding and generation functionalities without having to host or fine-tune models locally (Chen et al., 2025b).
While the integration of Gen-AI into mobile apps offers clear benefits in functionality and user experience, it also introduces a range of technical and practical challenges. Implementing Gen-AI features requires consideration of API integration, prompt engineering, token limitation, inference cost, and the unpredictable model outputs (Chen et al., 2025a). These issues can degrade the quality of Gen-AI-powered apps, leading to functional failures (e.g., crashes, incorrect behaviour, or suboptimal user interfaces), reduced efficiency (e.g., latency or excessive token consumption), and potential security vulnerabilities (Shao et al., 2025). Additionally, beyond implementation, quality assurance for Gen-AI-powered apps remains a concern. Unlike traditional software systems, Gen-AI components often lack formal specifications, making it challenging to evaluate correctness or performance in a deterministic way (Nahar et al., 2024). As a result, ensuring the reliability and robustness of Gen-AI-integrated mobile apps continues to be a significant challenge for developers.
However, despite growing interest in integrating LLMs into mobile apps, little is known about the specific concerns and experiences of end-users. Existing research primarily focused on the challenges faced by developers. For example, Nahar et al. (Nahar et al., 2024) employed interviews and surveys to investigate how developers address Gen-AI integration issues, including latency, energy consumption, cost, and fairness. While informative, their work is limited to developer perspectives and does not consider how these technical challenges impact end-user experience. Chen et al. (Chen et al., 2025a) analyzed posts from an OpenAI developer forum to uncover challenges in prompting, API use, and plugin development. This study offers insight into implementation bottlenecks but also neglects user-facing concerns. Similarly, Shao et al. (Shao et al., 2025) examined open-source mobile apps with Gen-AI integration and found that 77% experienced integration challenges, which impacted functionality, performance, or security issues. However, their analysis is limited to code-level integration problems and does not investigate whether users are aware of or affected by these issues.
In contrast, our work examines the user-perceived impact of Gen-AI-enabled apps, using large-scale review data to understand how users perceive Gen-AI features in practice. Existing studies on user perspectives of Gen-AI apps tend to focus on specific domains, such as education (Golding et al., 2024; Kim et al., 2025b; Lee et al., 2024a; Shata and Hartley, 2025; Kim et al., 2025a; Yuen and Schlote, 2024), creative image generation (Tang et al., 2024; Haase et al., 2023; Oppenlaender et al., 2023), or mental health (Jin et al., 2025; Heinz et al., 2025). Our work takes a systematic view by analyzing user reviews across a wide range of Gen-AI tools and application areas. To our knowledge, this is the first study to systematically examine the full landscape of topics users discuss across diverse Gen-AI domains. We identify the technical pain points surfaced by users, in addition to emerging behavioural patterns, expectations, and sources of satisfaction or frustration. This user-centered perspective complements developer-focused studies, providing critical insights for improving the usability and effectiveness of Gen-AI mobile apps.
Figure 1. Examples of reviews of Gen-AI apps
To gain these insights, we analyze user reviews that offer valuable perspectives on the challenges, expectations, and satisfaction associated with Gen-AI apps. For example, Figure 1 illustrates representative reviews from three Gen-AI apps, highlighting both praise and frustration. Building on this foundation, we introduce a four-phase methodology, Selection, Acquisition, Refinement, and Analysis (SARA). Our proposed methodology enables the large-scale collection and processing of user reviews from Gen-AI apps, supporting effective analysis using LLMs. First, we identify relevant Gen-AI apps (i.e., Selection) and collect user reviews from the Google Play Store (i.e., Acquisition). We then apply a data cleaning pipeline to enhance the quality of input data for LLM-based analysis (i.e., Refinement). Finally, we employ prompt-based LLM techniques to analyze the reviews (i.e., Analysis). Using this methodology, we conduct an empirical study on 676,066 reviews from 173 Gen-AI apps to address the following research questions (RQs):
RQ1: How accurately can LLMs identify topics in user reviews of Gen-AI apps?
{adjustwidth}
0.8cm In this RQ, we investigate how prompt configuration (e.g., using 0-shot, 3-shot, and 5-shot) and input quality (e.g., the proportion of informative reviews) affect the performance of LLMs in extracting topics from user reviews. Our results show that increasing the number of few-shot examples and filtering non-informative reviews significantly improve accuracy. The best-performing setup achieves an overall accuracy of 91%, confirming the reliability of our LLM pipeline for downstream analysis.
RQ2: What are the most prominent topics discussed in user reviews of Gen-AI apps?
{adjustwidth}
0.8cm To answer this RQ, we apply the LLM-based topic labelling pipeline from RQ1 to informative reviews across 173 Gen-AI apps. We analyze the top 10 topics in Gen-AI app reviews to understand user behaviour, key challenges and opportunities, and derive actionable insights for developers and researchers. Our findings reveal that users tend to rate Gen-AI topics positively and expect high performance, often providing constructive feedback from a collaborative mindset. Notably, some users form emotional bonds with conversational agents and value personalized interactions. These insights have practical implications for developers aiming to improve the usability, engagement, and trustworthiness of Gen-AI features.
RQ3: What temporal trends emerge in user feedback on Gen-AI topics? {adjustwidth}0.8cm In this RQ, we explore how user feedback on Gen-AI topics evolves as the technology matures. We cluster topics to identify evolutionary trends and support the analysis by qualitative review interpretation. Our analysis reveals that while the average rating for Gen-AI topics remains high, topic trends diverge: Content Quality ratings decline due to rising user expectations, AI Performance remains stable due to balancing effects, and Content Policy & Censorship ratings improve with better filtering mechanisms with time. These shifts reflect an increasingly critical and experienced user base that raises expectations for Gen-AI apps over time.
Our study provides valuable insights for app developers and researchers by uncovering how users perceive and evaluate Gen-AI features in mobile apps. Specifically, our contributions are as follows:
  • We propose a four-phase methodology, SARA, for analyzing user reviews of Gen-AI apps. Applying this methodology, we conduct a large-scale empirical study of 676,066 reviews from 173 Gen-AI apps on the Google Play Store to 1) identify the most frequently discussed Gen-AI-related topics, 2) uncover user expectations and pain points, and 3) analyze temporal trends to reveal how user perceptions evolve.
  • We evaluate the effectiveness of LLMs for topic extraction and assignment of user reviews, demonstrating that increasing the number of labelled examples and filtering out non-informative reviews significantly improves accuracy.
  • We provide actionable feedback for developers to maintain competitiveness, for policymakers to explore the ethical concerns surrounding content policy and censorship, and for user requirement engineers to better understand user needs.
  • We provide a replication package that includes the dataset of Gen-AI apps, three manually labelled samples for evaluating the filtering prompt, and the complete set of code we used to run all prompt-based experiments presented in this paper.
We organize the remainder of this paper as follows. Section 2 illustrates existing work on analyzing user reviews. Section 3 presents our methodology. Sections 46 describe the approach and the results obtained in our RQs. Section 7 synthesizes our key findings and implications of our work. Section 8 discusses potential threats to the validity of our findings. Finally, Section 9 concludes the paper and discusses implications for future research.
2. Related Work
In our literature review, we examine three key areas: 1) the user perspectives of Gen-AI apps, 2) user review analysis techniques, and 3) studies on prompt design and evaluation.
2.1. User Perspective of Gen-AI Apps
The user’s perspective on utilizing Gen-AI tools in education has been well studied (Golding et al., 2024; Kim et al., 2025b; Lee et al., 2024a). Shata and Hartley (Shata and Hartley, 2025) used surveys grounded in the Technology Acceptance Model (TAM) and Social Cognitive Theory (SCT) to examine faculty attitudes, identifying trust as the key driver of Gen-AI adoption. Kim et al. (Kim et al., 2025a) compared student and faculty perceptions on the use of Gen-AI in learning. Their study revealed that students and faculty used Gen-AI at similar rates, while students found Gen-AI tools easier to learn. Both faculty and students believed that Gen-AI will have more negative than positive effects on learning. Yuen and Schlote (Yuen and Schlote, 2024) surveyed adult learners to explore their views on the integration of AI in language learning apps. Their participants expressed a desire for in-app social features to assist in learning a language, and the authors suggested AI integration in the form of AI tutors that act as conversation partners.
In the creative image domain, Tang et al. (Tang et al., 2024) surveyed users of Gen-AI image tools focusing on satisfaction, challenges, acceptance, perceptions, and applicability of Gen-AI image tools. They found that users most frequently rely on these tools during the early ideation stage rather than for producing final outputs. In support of that, Haase et al. (Haase et al., 2023) examined how Gen-AI images influence creative idea generation, and demonstrated that Gen-AI art effectively supports ideation. Oppenlaender et al. (Oppenlaender et al., 2023) focused more on the user’s understanding of technology, fears, and thoughts on the risks of Gen-AI image tools in their survey. Their participants voiced concerns over a societal risk, as people could use these tools to generate images that spread misinformation.
In mental health, applications of AI include detection, diagnosis, treatment, and ongoing condition management (Sharma and Patel, 2024; Madhuri et al., 2025; S et al., 2025; Kheterpal and Gill, 2024). Our focus is on Gen-AI-driven treatment through conversational agents. Jin et al. (Jin et al., 2025) reviewed barriers to sustained engagement in mental health apps and suggested integrating AI in the form of AI agents that can detect emotional changes in users, build strong relationships with users, and promote timely intervention. Heinz et al. (Heinz et al., 2025) studied the efficacy of a mental health AI chatbot. They found that it significantly reduced users’ symptoms, maintained high engagement, and achieved user acceptance, with a reported therapeutic alliance comparable to that of human therapists.
Unlike prior work that relied on expert synthesis or self-reported perceptions, our study analyzes large-scale, real-world reviews from users actively engaging with Gen-AI apps. We offer a complementary perspective on actual usage and feedback. We also take a broad perspective by examining Gen-AI apps across all domains, without prioritizing any specific application area.
2.2. User Review Analysis Techniques
In this section, we summarize prior work related to 1) filtering user reviews and 2) topic extraction and assignment to user reviews.
Filtering user reviews. Automated filtering has shown strong potential in enhancing user review analysis by effectively removing non-informative content. Chen et al. (Chen et al., 2014) introduced a framework called AR-miner that used Latent Dirichlet Allocation (LDA) to first filter informative reviews with a hit-rate of 70%, and second to group the informative reviews using topic modelling. Ghosh et al. (Ghosh et al., 2024) evaluated fine-tuned LLMs for filtering non-informative user reviews on the Google Play Store and found them highly effective, achieving accuracies of 92% for both BERT and DistilBERT, and 91% for Gemma. Al Wahshat et al. (Al Wahshat et al., 2023) highlighted the potential of using GPT-4 in filtering manipulated reviews due to its powerful language processing capabilities.
Topic extraction and assignment to user reviews. Before the advancement of LLMs, researchers used various methods to analyze user reviews. Genc Nayebi and Abran (Genc-Nayebi and Abran, 2017) conducted a literature review in 2016 that explored approaches for mining opinions from user reviews. They identified three studies that relied on manual analysis, while others used automated techniques, such as Classification and Regression Trees (CART), word frequency statistics, correlation analysis, LDA, and Hexbin plots. Researchers and developers still use non-LLM approaches, even in recent work. For example, Arambepola et al. (Arambepola et al., 2024) used the Valence Aware Dictionary and sEntiment Reasoner (VADER) for sentiment analysis to categorize reviews into positive, neutral, and negative categories. They then used an LDA model for topic modelling of user reviews of educational apps. Their positive topics achieved a coherence score of 0.56, and their negative topics achieved a score of 0.49. Assi et al. (Assi et al., 2021a) proposed the Global-Local sensitive Feature Extractor (GLFE) to identify high-level features in user reviews, achieving 79–82% precision and 74–77% recall on annotated review data.
LLMs have demonstrated promising performance in tasks related to topic extraction and assignment. Arambepola et al. (Arambepola et al., 2025) conducted a recent systematic review of mobile app review analysis techniques and found that integrating LLMs into app user review analysis provides a more holistic and context-rich approach to evaluating reviews and understanding users. Pham et al. (Pham et al., 2024) introduced TopicGPT, a framework to perform topic extraction by prompting Open AI’s GPT (OpenAI, 2023) to generate topics and assign them to documents. Their framework outperformed traditional topic models, such as LDA, SeededLDA, and BERTopic. Prakash et al. (Prakash et al., 2023) introduced PromptMTopics, a prompt-based ChatGPT model set to extract topics from textual meme captions that outperformed baseline models. Assi et al. (Assi et al., 2025a) introduced LLM-Cure, an approach that extracts and assigns features from user reviews to make suggestions for feature improvement. Their approach outperformed baseline models, achieving an F1-score of 85%. Sorathiya and Ginde (Sorathiya and Ginde, 2024) employed a hybrid approach that combines Natural Language Inference (NLI) and LLMs to identify privacy concerns in mental health app reviews, aiming to uncover ethically relevant user feedback that traditional keyword-based methods often overlook. Their approach achieved an F1-score of 81%.
Instead of extracting and assigning topics to user reviews, prior studies focused on a single task (i.e., either extracting topics from user reviews or assigning predefined topics to user reviews). For example, Wei et al. (Wei et al., 2023) focused on review assignment using predefined labels rather than extracting topics. They assigned reviews using fixed categories such as feature requests, problem reports, and irrelevant feedback. While useful, this approach limits the discovery of emergent or evolving user concerns. Dos Santos et al. (Dos Santos et al., 2023) extracted topics from user reviews without performing the assignment task. Their study examined the impact of app updates on accessibility by utilizing GPT-4 to identify key issues, barriers, improvements, update consequences, and user sentiments. While this approach is effective for uncovering topical insights, it does not account for the frequency or distribution of each topic, limiting its usefulness for trend or prevalence analysis. Ren et al. (Ren et al., 2024) used LDA for topic extraction and pretrained LLMs for requirement extraction. Rezaei Nasab et al. (Rezaei Nasab et al., 2025) conducted a study of AI-based mobile app reviews to identify fairness-related concerns. They constructed a labelled dataset of fairness vs. non-fairness reviews, developed machine learning and LLM-based classifiers to detect fairness mentions, applied K-Means clustering, and performed manual coding to identify six types of fairness concerns.
Our work builds on this line of research by utilizing LLMs to filter out non-informative reviews and identify topics in user reviews of Gen-AI apps. Based on the analysis of user reviews, we identify the challenges and potential opportunities to adopt Gen-AI functionalities in mobile apps.
2.3. Prompt Design and Evaluation
Prompt engineering is the process of designing and refining prompts to effectively guide LLMs in performing specific tasks (Koyuncu, 2025). Prior work discusses prompt design in three directions: 1) the process of prompt design, 2) the definition of the required tasks, and 3) identifying the number of included examples.
The process of prompt design. Marvin et al. (Marvin et al., 2024) emphasized the importance of prompt refinement and outlined several prompt engineering techniques, including dynamic prompting, which refers to iteratively adjusting the prompt based on the generated output. He et al. (He et al., 2024) investigated the impact of prompt format across various tasks and found that prompt structure significantly affects performance, with different formats performing best in different task types. In our work, we adopt a prompt format inspired by prior work (Assi et al., 2025a) and refine it to suit our study’s goals, employing dynamic prompting to enhance accuracy and relevance further.
The definition of the required tasks. To ensure accurate, ethical, and reliable AI outputs, effective prompt engineering requires understanding common pitfalls, such as ambiguity, bias, and lack of context (Giray, 2023). Prompts can be explicit, implicit, or creative, varying in the amount of guidance they offer to the AI. Explicit prompts provide clear instructions, while implicit and creative prompts allow more flexibility and imagination but may produce less predictable results (Data Science Horizons, 2023). Ho et al. (Ho et al., 2024) adopted a prompt classification framework consisting of the three types. The authors analyzed how different prompts affect AI-generated responses. Explicit prompts generate specific and fact based answers, implicit prompts produce broader and more flexible responses, while creative prompts encourage innovation and originality. In our work, we adopt an explicit prompting strategy to improve structure and reproducibility.
Identifying the number of included examples. Deshmukh et al. (Deshmukh et al., 2025) conducted a literature review of prompt engineering and discussed 16 LLM prompt patterns to enhance output accuracy. Their analysis highlighted the importance of clarity, specificity, defined output formats, and contextual examples to improve model accuracy. However, it did not address techniques for optimizing example selection (i.e., determining the optimal number of examples to include in the prompt). Complimenting this, Wang et al. (Wang et al., 2023) demonstrated that using similar examples and increasing the number to four enhances performance, while adding more than four examples yields only marginal gains. Their work focused on classification, multiple-choice, and generation tasks, rather than user review analysis tasks. Prakash et al. (Prakash et al., 2023) evaluated prompts with 2, 4, 6, and 8 examples, finding comparable topic coherence and diversity across configurations.
In this paper, we further extend this line of research by investigating how the number of in-context examples and the exclusion of noisy data affect topic extraction and assignment tasks for user review. Our work demonstrates that increasing the number of examples for topic extraction from 0-shot to 5-shot significantly improves the model’s accuracy. Our work also suggests that filtering out non-informative reviews enhances the accuracy.
3. Methodology
To study user feedback on Gen-AI apps, we design a four-phase methodology we call SARA, which stands for Selection, Acquisition, Refinement, and Analysis. SARA encompasses 1) the selection of relevant Gen-AI apps, 2) the acquisition of user reviews, 3) a refinement data cleaning pipeline to ensure data quality for LLM analysis, and finally, 4) the LLM-based analysis of user reviews through topic extraction and assignment. Our proposed methodology enables us to address our RQs concerning the impact of data quality and few-shot examples on LLM performance, the prevalent topics within Gen-AI user reviews, and the temporal evolution of user ratings toward these apps. We present an overview of our research approach in Figure 2.
Figure 2. Overview of our approach
3.1. Selection of Gen-AI apps
Our app selection process involves three steps: 1) generating keywords from prior work, 2) searching the Google Play Store and scraping relevant apps, and 3) manually screening and validating each app to confirm it is Gen-AI.
  1. 1.
    Keyword generation. We compile a list of keywords derived from a Gen-AI taxonomy by Gozalo-Brizuela and Garrido-Merchán (Gozalo-Brizuela and Garrido-Merchán, 2023). The authors organized Gen-AI apps into 14 categories, each with a taxonomy of the relevant technology. We exclude four categories, i.e.,“biotech,” “brain,” “drug discovery,” and “other,” because an Android mobile app user does not typically use them. We create a list of ten categories relevant to mobile apps and repurpose the taxonomy as keywords, as shown in Table 1.
  2. 2.
    App scraping. Using our list of keywords, we scrape the Google Play Store to identify all matching apps. Some keywords (e.g.,“Image Editing” and “Business”) are too broad, so we append “Generative AI” for all keywords when searching the Google Play Store. For each identified app, we collect descriptive metadata, including app ID, app name, Google Play Store category, star rating, number of downloads, number of users who rated the app, and the app description. In this paper, we refer to Google Play Store app categories as app categories.
  3. 3.
    Manual screening and validation. The initial search yields 303 apps. To ensure the inclusion of only genuine Gen-AI apps, the first and second authors independently review the apps by looking at the title, description, and images of the apps. We assess the inter-rater reliability, using Cohen’s Kappa (Cohen, 1960) ($\kappa=0.61$), which indicates substantial agreement. In case of disagreement, the two authors discuss and resolve the discrepancies for the ambiguous cases. This rigorous process results in a final list of 194 confirmed Gen-AI apps.
Table 1. Categories and keywords for app search in the Google Play Store
Category Keywords
Text Conversational AI, Text-to-Science, Text-to-Author Simulation, Text-to-Medical Advice, Text-to-Itinerary, Doc-to-Text
Image Image Editing, Artistic Images, Realistic Images
Video Text-to-Video, Video Production, Image-to-Video
3D Text-to-3D
Code and Software Text-to-Code, Text-to-Multilingual Code
Speech Text-to-Speech, AI Understanding, Speech-to-Text
Business Business
Marketing Marketing, Advertisement, Accounting
Gaming Videogame Creation
Music Music Generator
3.2. Acquisition of User Reviews
Following app selection, we collect user reviews using the Google Play Store Scraper (Mingyu, nd). For each retrieved review, the crawler extracts a comprehensive set of metadata, including reviewId, userName, userImage, content (i.e., review text), score (i.e., user rating from 1 to 5 stars), thumbsUpCount, reviewCreatedVersion, at (i.e., timestamp of the review), replyContent, repliedAt, and appVersion. Our subsequent analysis primarily focuses on the content, score, and at fields. We organize the collected reviews based on the Google Play Store categories (app categories) of their respective apps. At the end of this step, we collect 2,778,643 reviews of 186 apps, as eight apps from our list have no textual reviews.
3.3. Multi-Stage User Review Refinement
To ensure the robustness and reliability of our LLM-based analysis, we implement a multi-stage data cleaning pipeline designed to enhance the quality of the user review data progressively. This multi-stage refinement process comprises four distinct stages. We refine data in stages because we are interested in experimenting with different levels of data cleaning in RQ1. We summarize the number of reviews retained after each cleaning stage in Table 2. Each stage builds upon the previous one as follows:
Table 2. Number of reviews by app category and stage.
App category Stage 0 Stage 1 Stage 2 Stage 3
Productivity 362,210 354,829 233,311 169,119
Photography 1,035,946 430,490 188,382 126,277
Art & Design 332,881 238,291 157,320 125,362
Entertainment 205,467 205,467 156,636 118,669
Education 101,812 95,590 57,680 41,905
Video Players & Editors 364,124 109,610 51,182 36,507
Tools 231,865 104,636 48,807 35,046
Music & Audio 17,875 17,875 11,781 9,150
Health & Fitness 41,537 10,368 8,747 7,849
Games 29,922 6,254 5,555 5,070
Books & Reference 2,071 2,071 1,423 1,112
Total 2,725,710 1,575,481 920,824 676,066
Stage 0: basic cleaning. Text preprocessing, as demonstrated in prior work (Ghosh et al., 2024; Prakash et al., 2023; Roumeliotis et al., 2024; Lee et al., 2024b), is a standard practice for eliminating noise and ensuring more effective analysis. It involves the removal of punctuation, special characters, extra white spaces, emojis, and irrelevant data such as usernames and URLs. In the initial stage, we exclude empty reviews and apps with no reviews, as we are interested in analyzing the textual content of the reviews. Then, we perform fundamental text preprocessing, including replacing commas with white spaces and removing special characters and emojis from the review content.
We exclude five categories that contain fewer than 1,000 reviews, as small sample sizes tend to produce repetitive and low-quality topics. To validate this threshold, we experiment with categories of varying sizes and find that categories with fewer than 1,000 reviews yield topics that lack diversity and informativeness. For instance, the Personalization app category, which contains only 23 reviews, produces two nearly identical and non-informative topics labelled Positive Feedback and Praise. Additionally, some topics focus on non-informative content, such as Personal Comments or Signatures, which capture examples like “hello marlyn” and “Placau!!!!” that are not relevant to the app’s functionality or performance. Based on these observations, we adopt 1,000 reviews as the minimum threshold and exclude any app categories that fall below this cutoff.
Stage 1: temporal filtering. This stage filters reviews written prior to 2022 to ensure Gen-AI relevancy. Some apps included in our initial screening are not originally designed as Gen-AI tools but have integrated Gen-AI functionalities at a later stage. For instance, in the case of Microsoft Bing Search [^4] https://play.google.com/store/apps/details?id=com.microsoft.bing, Bing Chat is a recently introduced feature that extends the capabilities of the search engine by incorporating Gen-AI models (Iorliam and Ingio, 2024). Consequently, it is essential to exclude user reviews published before the integration of such features, as they reflect experiences with the app before the adoption of Gen-AI. While foundational work on AI models began as early as 2012, notable advancements in Gen-AI only started to gain momentum after 2018 (Sengar et al., 2024). The year 2018 also marked a rise in academic interest and publications related to Gen-AI (Dwivedi and Elluri, 2024). However, the field witnessed a major breakthrough in 2022, particularly with the public release and widespread adoption of ChatGPT, which significantly accelerated the prominence and deployment of Gen-AI technologies across various platforms (Akhtar, 2024). To maximize the relevance and accuracy of the user feedback analyzed in our study, we adopt 2022 as the cutoff year for including reviews.
Stage 2: short review exclusion. Building upon the cleaned data from Stage 1, this stage employs a simple heuristic to further reduce noise by excluding reviews containing three words or less. Short reviews of three words or less do not contain context for meaningful topics (Assi et al., 2021a; Arambepola et al., 2024). This step aims to eliminate overly brief and potentially non-informative feedback.
Stage 3: LLM-assisted informative review filtration. Non-informative reviews bring noise to user reviews analysis and influence the performance of topic extraction (Chen et al., 2014). To improve the quality of extracted topics, we exclude non-informative reviews. We define non-informative reviews as short, generic statements that express sentiment without referencing specific reasons or features. While Chen et al. (Chen et al., 2014) focused on informativeness from a developer’s perspective, we expand this definition to consider the informational needs of multiple stakeholders, including users, platform owners, developers and researchers. We carry out the LLM-assisted process in the following steps.
  1. 1.
    Prompt design. To filter non-informative reviews, we utilize Open AI’s (OpenAI, 2023) GPT-4o-mini model with zero temperature for reproducibility. We use gpt-4o-mini instead of the full gpt-4o model because it offers significantly faster response times while maintaining reasonable output quality for our task, which does not require complex reasoning (OpenAI, 2024b).
    We design a filtering prompt that instructs the LLM to classify reviews as either informative or non-informative. To guide the model effectively, we incorporate representative examples within the prompt.
  2. 2.
    Iterative prompt refinement. We use small statistically significant samples of 96 reviews (95% confidence interval, 10% margin of error) from three diverse app categories (Photography, Productivity, and Art & Design). Then, the first and second authors manually label the reviews in the samples as informative or non-informative. We calculate the inter-rater reliability using Kappa (Cohen, 1960) ($\kappa=0.91$), which indicates almost perfect agreement.
    Next, we design and test different versions of the filtering prompt against the manually labelled data, evaluating the model’s filtering accuracy and iteratively refining the prompt. Revisions include modifying the wording, structure, number of examples, and types of examples. We repeat this iterative process until we achieve a peak filtering accuracy of 91% on held-out validation samples. Figure 3 shows the final version of the filtering prompt, which includes ten examples (five informative and five non-informative) and yields optimal performance.
    Figure 3. Template of the filtering prompt used to filter out non-informative reviews.
  3. 3.
    Prompt validation. To assess the generalizability of the final filtering prompt and reduce the risk of overfitting to specific review instances, we manually evaluate its performance on a separate set of previously unseen reviews of the same categories and the same size. The model demonstrates strong generalization capabilities, achieving an average classification accuracy of 90%.
  4. 4.
    Large-scale review filtering. We apply the validated filtering prompt to the full dataset to filter all non-informative reviews. Reviews are processed in batches of 100 to minimize omissions, as larger batches sometimes result in the model returning only a subset of reviews. Any unreturned reviews are consolidated into new batches and resent. This process reduces 920,824 reviews to 676,066 informative reviews, as shown in Table 2.
Our data collection, spanning September and October 2024, resulted in an initial dataset of 2,778,643 user reviews across 186 unique Gen-AI apps. Following our multi-stage cleaning process, we obtain a dataset of 676,066 informative reviews across 173 Gen-AI apps, which forms the basis for our subsequent analysis.
3.4. LLM-powered user review analysis
We use the refined user review data to conduct topic extraction, topic assignment, trend analysis, and manual analysis. For RQ1, we design and evaluate our topic extraction and assignment prompts. We experiment with different data noise levels (i.e., different stages) and numbers of few-shot examples. For RQ2, we extract the top 10 topics per app category and assign them to user reviews. We conduct manual thematic analysis for each Gen-AI topic to derive user insights from reviews. For RQ3, we analyze how Gen-AI topics evolve over time using clustering and manual review. In the following sections, we describe the experiment design and the findings for each RQ.
4. RQ1: How accurately can LLMs identify topics in user reviews of Gen-AI apps?
Effective topic extraction and assignment of user reviews are essential for understanding user needs and improving app features for Gen-AI apps. Prior work has shown that prompt engineering techniques can help tailor LLMs to specific tasks (Marvin et al., 2024). Applying LLMs to topic extraction and assignment requires well-designed prompts to achieve high accuracy. In this RQ, we investigate two key aspects of prompt design: 1) the extent to which cleaning the input data prior to prompting improves accuracy, and 2) the effect of varying the number of few-shot examples in the prompt. The goal is to identify the optimal combination of preprocessing and few-shot-based prompting that enables more accurate and reliable topic extraction and assignment from user reviews.
4.1. Experiment Setup
To design an effective topic extraction prompt, we experiment with varying levels of data cleaning and different numbers of few-shot examples. As described in section 3, we design a four-stage data cleaning pipeline. Stage 0 includes basic textual cleaning. Stage 1 excludes reviews before the launch of ChatGPT in 2022 (Akhtar, 2024). Stage 2 excludes short reviews with three words or less. Stage 3 excludes non-informative reviews using an LLM-based filtering approach. We conduct experiments on Stage 1, Stage 2, and Stage 3, but not on Stage 0. We exclude Stage 0 because while it performs essential preprocessing, it does not remove irrelevant data as done in Stage 1. We approach this problem in four steps: 1) we generate random samples of user reviews for topic extraction and assignment drawing from each stage of data cleaning, 2) we apply three prompting strategies with varying numbers of few-shot examples (i.e., 0-shot, 3-shot, and 5-shot) to extract topics from these samples, 3) assign the extracted topics to a new set of sample reviews using the same prompts, and (4) we evaluate the accuracy of the topic assignment to determine the optimal combination of data cleaning and few-shot examples. We summarize this process in Figure 4 and detail it in this section.
Figure 4. An overview of our RQ1 experiment design, where $x\in\{Stage1,Stage2,Stage3\}$ and $y\in\{Photography,Productivity,Art\&Design,Entertainment,VideoPlayers\&Editors\}$.
Sampling. To evaluate the effectiveness of our topic extraction prompts across different stages of data cleaning, we define $x\in\{Stage1,Stage2,Stage3\}$ to represent the three data cleaning stages. For each stage $x$, we select five representative app categories from the Google Play Store, i.e., Photography, Productivity, Art & Design, Entertainment, and Video Players & Editors, denoted as $y$. For each combination of cleaning stage $x$ and app category $y$, we generate two separate statistically representative random samples of user reviews from our review dataset. First, we extract a large sample denoted $S_{\text{large}}(x,y)$, computed using a 95% confidence level and a 2% margin of error with an average size of 2,337 reviews. This large sample is used for topic extraction to ensure the coverage of the user feedback space. Second, we draw a small sample denoted $S_{\text{small}}(x,y)$ from our review dataset, computed using a 95% confidence level and a 10% margin of error with an average size of 96 reviews. This sample is used for topic assignment, where manual evaluation is feasible and interpretable.
Topic extraction. To evaluate how prompt design affects topic extraction performance, we design three experiments, i.e., E1: 0 shot, E2: 3 shot and E3: 5 shot, with three prompt configurations differing in the number of few-shot examples, i.e., 0-shot, 3-shot and 5-shot. These values are chosen based on prior research and practical considerations in prompt-based learning  (Dos Santos et al., 2023; Assi et al., 2025a; Pham et al., 2024). In E1: 0 shot, the prompt contains only the instructions. In E2: 3 shot, the prompt includes the instruction with three shots, i.e., three examples of user reviews with their topics. In E3: 5 shot, the prompt includes the instructions with five examples of user reviews with their topics. For each cleaning stage $x\in\{\text{Stage 1},\text{Stage 2},\text{Stage 3}\}$ and app category $y$, we apply each of the three prompt configurations to the large sample $S_{\text{large}}(x,y)$. This results in a set of extracted top topics, denoted as $T(x,y,n)$, where $n\in\{0,3,5\}$ indicates the number of few-shot examples used in the prompt. Figure 5 shows the three prompt configuration templates.
Figure 5. Template of the topic extraction prompts for (E1: 0-shot, E2: 3-shot, and E3: 5-shot) used to extract the top topics from a large sample of stage $x$ and app category $y$. Insert category name is where we insert the name of the app category $y$. Insert sample of reviews is where we list our large sample of reviews separated by a new line escape.
Topic assignment. To evaluate the effectiveness of the extracted topic sets $T(x,y,n)$ produced by each topic extraction experiment with $n\in\{0,3,5\}$, we perform a follow-up topic assignment task using LLM. Specifically, for each experiment $n$, we design a dedicated assignment prompt, illustrated in Figure 6, that takes as input: 1) the set of top topics $T(x,y,n)$ extracted in that experiment, 2) five few-shot example reviews manually selected and labelled from the corresponding large sample $S_{\text{large}}(x,y)$, and 3) the target reviews from the small sample $S_{\text{small}}(x,y)$ to which topics are assigned. We manually select and assign topics to five few-shot examples to improve assignment accuracy as done in prior work (Assi et al., 2025a; Deshmukh et al., 2025).
Figure 6. Template of the P5: topic assignment prompt used to assign topics to a list of reviews. We insert the set of top topics $T(x,y,n)$, five few-shot examples, and the small sample $S_{\text{small}}(x,y)$.
Accuracy evaluation. To assess the accuracy of the LLM-assigned topics, we conduct a manual annotation of the assignments. The first and second authors manually and independently label the correctness of LLM assignment. An assignment is considered correct if the LLM assigned topic accurately reflects the content of the review. The Cohen’s Kappa agreement score (Cohen, 1960) is computed, yielding a score of 0.56, indicative of moderate agreement. The two authors discuss and resolve the initial disagreements to align on labelling criteria. In total, 4,353 LLM topic assignments are manually evaluated for correctness for five categories, three stages, and three experiments. Table 3 shows the accuracy for each cleaning stage, app category and shot examples. To determine whether the observed accuracy differences are statistically significant, we perform statistical tests. Specifically, we apply the non-parametric test, Kruskal–Wallis H test (Chan and Walmsley, 1997), to evaluate the effect of few-shot prompt configurations on overall accuracy. Additionally, post-hoc pairwise comparisons are conducted using the Mann–Whitney U test (Mann and Whitney, 1947) with Bonferroni correction to adjust for multiple testing (Cabin and Mitchell, 2000). We also use one-way ANOVA (Judd et al., 2017) to test for significant differences across cleaning stages and few-shot settings.
4.2. RQ1 Findings
Observation 1.1: Providing more examples significantly improves classification and assignment accuracy. The average accuracy increases across experiments for each cleaning stage, with E3: 5-shot consistently achieving the highest accuracy in all three stages, as shown in Table 3. Notably, E3: 5-shot is the only setting to achieve average accuracies above 80% across all samples. The Kruskal–Wallis test indicates that the number of few-shot examples has a statistically significant effect on overall accuracy ($H=7.778$, $p=0.0205$). Post hoc Mann–Whitney U tests with Bonferroni correction show that the accuracy improvement from E1: 0 shot to E3: 5 shot is statistically significant ($p=0.0181$), whereas differences between E1: 0 shot and E2: 3 shot ($p=1.000$) and E2: 3 shot and E3: 5 shot ($p=0.2519$) are not. These results suggest that increasing the number of few-shot examples, specifically 5-shot, significantly enhances overall model accuracy.
Table 3. Accuracy across three stages and three experiments of five app categories: 1) Photography (Photo), 2) Productivity (Pro), 3) Art & Design (Art), 4) Entertainment (Ent), and 5) Video Players & Editors (Video).
Experiment Stage 1: Temporal reviews Stage 2: Short reviews Stage 3: Non-informative reviews
Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg.
E1: 0 shot 86% 77% 91% 69% 95% 83% 84% 85% 85% 80% 90% 85% 72% 86% 77% 86% 89% 82%
E2: 3 shot 91% 86% 80% 66% 95% 83% 91% 80% 86% 82% 95% 87% 76% 83% 88% 92% 90% 86%
E3: 5 shot 93% 81% 94% 90% 92% 90% 90% 84% 92% 88% 93% 89% 95% 86% 86% 95% 91% 91%
Observation 1.2: The proportion of informative reviews varies by cleaning stage, which impacts overall accuracy. As shown in Table 3, accuracy does not follow a consistent trend across the three cleaning stages (i.e., Stage 1, Stage 2, and Stage 3). For instance, the average accuracy for E1: 0 shot and E2: 3 shot increase from Stage 1 to Stage 2, but decrease from Stage 2 to Stage 3. In contrast, E3: 5 shot shows a slight decrease from Stage 1 to Stage 2, followed by a slight increase in Stage 3. This inconsistency may be explained by differences in the proportion of informative reviews across stages: 43% in Stage 1, 66% in Stage 2, and 88% in Stage 3. Since assigning topics to non-informative reviews is generally easier and less dependent on nuanced understanding, variations in their prevalence can skew overall accuracy. As a result, overall accuracy may not accurately reflect the model’s effectiveness in handling more meaningful and content-rich reviews across different cleaning stages.
Observation 1.3: Stage 3 consistently yields the highest accuracy on informative reviews across all app categories and experiments. Given the inconsistencies observed in overall accuracy explained in observation 1.2, we isolate informative reviews to better evaluate model performance. Table 4 shows that, across all experiments and app categories, average accuracy on informative reviews is highest in Stage 3. This suggests that data cleaning improves the model’s ability to assign topics to reviews with informative content correctly. We conduct a one-way ANOVA test to evaluate whether the cleaning stage has a significant effect on classification accuracy. The ANOVA for few-shot experiments does not reveal significant differences ($F(2,42)=2.25$, $p=0.118$).
Table 4. Accuracy of informative reviews across stages and app categories, Photography (Photo), Productivity (Pro), Art & Design (Art), Entertainment (Ent), Video Players & Editors (Video).
Experiment Stage 1: Temporal reviews Stage 2: Short reviews Stage 3: Non-informative reviews
Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg.
E1: 0 shot 83% 76% 90% 88% 76% 83% 84% 83% 89% 74% 82% 82% 81% 91% 81% 86% 89% 86%
E2: 3 shot 87% 91% 87% 86% 76% 85% 93% 78% 89% 89% 94% 88% 81% 88% 89% 91% 90% 88%
E3: 5 shot 87% 85% 94% 82% 76% 85% 80% 79% 90% 82% 89% 84% 94% 92% 90% 95% 97% 94%
Observation 1.4: The best performing setup combines filtering non-informative reviews, i.e., Stage 3 and more few shot examples, i.e., E3: 5-shot. Our findings suggest that using LLMs to filter the non-informative reviews and increasing the number of few-shot examples to five improves the accuracy of the model. Based on these findings, we adopt the Stage 3 – E3: 5-shot configuration for topic classification and assignment in RQ2, as it yielded the highest overall accuracy of 91% and the highest informative accuracy of 94%.
Summary of RQ 1 Our findings indicate that increasing the number of few-shot examples in the prompt significantly improves output accuracy. Additionally, filtering out non-informative reviews enhances the accuracy of extraction and assignment for informative reviews. The model using the 5-shot prompt combined with LLM-based filtering of non-informative reviews achieves the highest overall accuracy of 91%. Accordingly, we adopt this configuration for topic extraction in RQ2.
5. RQ2: What are the most prominent topics discussed in user reviews of Gen-AI apps?
Understanding the issues that users care about is essential for improving the design and evaluation of Gen-AI apps. In RQ2, we uncover the top 10 topics that users discuss in reviews of Gen-AI apps across different app categories. We then analyze topic frequency and associated review ratings to interpret common user behaviours and needs, supporting the identification of actionable insights that inform the design of more adaptive and user-centered Gen-AI experiences.
5.1. Experiment Setup
In RQ1, we experiment with varying levels of data cleaning and numbers of few-shot examples. We find that filtering out non-informative reviews using an LLM and providing five-shot examples improves topic extraction and assignment accuracy. In RQ2, we apply the optimal topic extraction prompt identified in RQ1 to the remaining app categories and assign topics to all reviews in Stage 3. In this section, we provide a detailed description of the implementation of RQ2.
Topic extraction and assignment. The results of RQ1 indicate that the optimal dataset for analysis is Stage 3 data, which underwent four stages of cleaning: basic cleaning, temporal filtering, short review exclusion, and LLM-assisted informative review filtration. Our dataset spans 11 app categories, totalling 676,066 user reviews as shown in Table 2. As in RQ1, we sample a statistically significant set of informative reviews (large sample) per app category, using a 95% confidence level and 2% margin of error. The resulting sample sizes range from 761 to 2,368 reviews, with an average of 2,031 per category. For each app category $y$, we extract the top 10 topics from the large sample using the validated 5-shot topic extraction prompt shown in Figure 5. Next, we use the topic assignment prompt shown in Figure 6 to assign one of the extracted topics to each review in the dataset. For each app category $y$, the corresponding top 10 topics are used as input to label reviews with a single topic.
Classification and categorization. As our primary interest lies in understanding user feedback related to Gen-AI functionality, we identify and analyze reviews that discuss such features. To do so, we manually examine all extracted topics and label them as either Gen-AI topics, which reference Gen-AI capabilities, or non-Gen-AI topics, which pertain to general app features outside the scope of Gen-AI. We further group semantically related topics into high-level topic categories based on shared themes or functionality, facilitating higher-level interpretation of user concerns. In line with prior studies (Assi et al., 2025a; Pagano and Maalej, 2013; Chen et al., 2014), we also classify reviews with 1 or 2-star ratings as negative, 3-star rating as neutral, and 4 or 5-star rating as positive.
Analysis. We perform quantitative analysis on both app categories and topic categories. We compute the following metrics:
  1. (1)
    Average Gen-AI Rating (AGR): the average star rating of reviews labelled with Gen-AI topics.
  2. (2)
    Average Non-Gen-AI Rating (ANR): the average rating of reviews labelled with non-Gen-AI topics.
  3. (3)
    Gen-AI Review Count (GRC): the number of reviews assigned to Gen-AI
  4. (4)
    Non-Gen-AI Review Count (NRC): the number of reviews assigned to non-Gen-AI topics
  5. (5)
    Gen-AI Review Percentage (GRP): the proportion of each type relative to the total number of reviews in each app category
  6. (6)
    Non-Gen-AI Review Percentage (NRP): the proportion of each type relative to the total number of reviews in each topic category
We conduct the Mann–Whitney U test to examine whether reviews labelled as Gen-AI-related receive significantly higher ratings than those that are not.
To gain deeper insights into user concerns and feedback within each topic category, we randomly sample reviews using a 95% confidence level and a 10% margin of error, resulting in a dataset of 901 reviews. We manually code and analyze all sampled reviews to provide richer, contextualized insights that may surface actionable implications for both researchers and developers.
5.2. RQ2 Findings
Observation 2.1: Gen-AI reviews consistently receive higher user ratings (AGR) than non-Gen-AI reviews (ANR) across all app categories. As shown in Table 5, the average rating of reviews discussing Gen-AI functionality exceeds the average rating of non-Gen-AI reviews in every app category, indicating higher user satisfaction with the Gen-AI features.
To assess the statistical significance of this difference, we apply the Mann–Whitney U test. The results reveal a significant difference in the rating distributions between the two groups ($U=65{,}113{,}513{,}850$, $p<0.0001$). Gen-AI reviews tend to receive higher ratings, with a median rating of 5.0 and a mean of 4.10, compared to a median of 4.0 and a mean of 3.32 for non-Gen-AI reviews. These findings suggest that Gen-AI functionality is a major contributor to positive user experiences across a wide range of app categories.
Table 5. Summary of app categories, including average Gen-AI rating (AGR), average non-Gen-AI rating (ANR), Gen-AI Reviews Count (GRC), Gen-AI reviews percentage (GRP), and number of distinct Gen-AI topics
App Category AGR ANR GRC GRP # Gen-AI Topics
Productivity 4.2 3.5 84,122 50% 6
Entertainment 4.0 3.2 56,672 48% 6
Education 4.6 3.3 31,465 75% 6
Art & Design 3.4 3.2 26,585 21% 3
Photography 3.9 3.3 15,567 12% 2
Tools 4.0 3.4 7,782 22% 3
Health & Fitness 4.8 4.0 6,896 88% 7
Video Players & Editors 4.6 3.9 5,179 14% 2
Games 3.6 2.7 1,788 35% 5
Music & Audio 3.0 2.6 1,679 18% 4
Books & Reference 4.1 2.3 428 39% 4
Average 4.0 3.2 21,651 38% 4
Observation 2.2: Gen-AI functionality is a prominent focus in user feedback, comprising an average of 38% of reviews across app categories and reaching as high as 88% for the Health & Fitness category. As reported in Table 5, the prevalence of Gen-AI–focused reviews varies by app category, reflecting the differing roles Gen-AI plays across app domains. For example, categories such as Health & Fitness, Education, and Productivity contain the highest proportions of Gen-AI reviews, with 88%, 75%, and 50%, respectively. Conversely, categories such as Photography and Video Players & Editors exhibit lower percentages of Gen-AI reviews. In these categories, 58% of the reviews focus on non-Gen-AI topics, such as overall user experience, technical issues, and subscription or pricing models, suggesting that these factors outweighed interest in Gen-AI features.
Figure 7 illustrates the distribution of Gen-AI review percentages across individual apps within various categories, highlighting key differences in engagement. Several categories, such as Education and Productivity, exhibit a moderate spread, indicating consistent Gen-AI engagement across apps. Other categories, such as Art & Design and Entertainment present a wider spread, suggesting more variability in Gen-AI adoption among apps within these groups.
Figure 7. Box plot showing the distribution of the percentage of reviews classified as Gen-AI at the app level, grouped by app category
$\Rightarrow$ Implication 1.
Gen-AI reviews consistently yield higher user ratings across app categories. Developers must recognize Gen-AI as a crucial tool for boosting overall user satisfaction. They must strategically incorporate Gen-AI features to remain competitive.
Observation 2.3: The Gen-AI topics with the highest numbers of reviews are AI Performance, Utility & Use Cases, and Content Quality. Table 6 and Table 7 present the top Gen-AI and non-Gen-AI topics, respectively. In terms of user satisfaction, the Gen-AI topics with the highest average ratings are Human AI Interaction & Emotional Connection, Utility & Use Cases, and Customer Support & Community. The topics most commonly mentioned across different apps are AI Performance, Content Quality, and Content Policy & Censorship. We further examine each of these categories in the subsections below.
Table 6. Summary of Gen-AI topic categories, including the Gen-AI Reviews Count (GRC), the average Gen-AI rating (AGR), the number of apps, and the number of app categories in which each topic category appears
Gen-AI Topic Category GRC AGR # Apps # App Categories
AI Performance 61,071 4.0 127 10
Utility & Use Cases 56,505 4.8 45 3
Content Quality 44,594 3.6 131 8
Content Policy & Censorship 27,521 2.9 76 5
Human AI Interaction & Emotional Connection 25,142 4.8 13 4
Features & Functionality 16,275 4.0 58 4
Customer Support & Community 3,361 4.7 12 3
Comparison to Other Apps 2,315 3.9 13 3
Personalization & Customization 1,283 4.0 30 3
Accessibility & Inclusivity 96 4.1 4 1
Table 7. Summary of non-Gen-AI topic categories, including the non-Gen-AI Review Count (NRC), the average non-Gen-AI rating (ANR), the number of apps, and the number of app categories in which each topic category appears
Non-Gen-AI Topic Category NRC ANR # Apps # App Categories
Technical Difficulties 99,621 2.2 157 10
Monetization Methods & Structure 85,756 2.3 159 11
User Interface & Experience 80,827 4.5 156 11
Utility & Use Cases 21,464 4.8 34 1
Updates & Evolution 21,100 2.8 92 7
Personalization & Customization 6,860 4.8 18 1
Customer Support & Community 6,848 1.8 85 6
Limitations of Usage 5,971 2.5 45 4
Features & Functionality 1,810 3.0 11 1
Comparison to Other Apps 1,493 3.4 17 1
Accessibility & Inclusivity 401 2.6 5 2
Other (unassigned) 105,752 4.2 165 11
1) AI Performance. AI Performance refers to how well the AI model performs in terms of 1) accuracy (27%), 2) speed (22%), 3) understanding (15%), 4) memory (5%), 5) reliability (5%), and 6) creativity & randomness (3%). AI accuracy describes the precision of the AI response to the user’s prompt. AI speed represents how quickly the AI responds to a user’s prompt. AI understanding indicates the AI’s ability to comprehend the meaning, intent, and context behind the user’s prompt. AI memory signifies how well the AI remembers prior conversations and links them to the current one. AI can have both short-term memory, which spans a single conversation, and long-term memory, which spans multiple conversations. Reliability refers to the consistency and dependability of the AI system in performing its intended tasks without crashing, freezing, timing out, or failing to generate output. Creativity captures the degree of originality, uniqueness, or unpredictability in the AI’s responses.
Figure 8. Examples of reviews discussing AI Performance
Insights. We find that 79% of the users who discussed AI Performance provided a positive review with a rating of 4 or 5. Reviews 1 and 2 in Figure 8 show examples of positive reviews. Despite this overall satisfaction, 16% of reviews mentioned limitations in AI Performance, with the most frequently cited issue being AI understanding (5%), followed by inconsistency in accuracy (4%), reliability (3%), speed (2%), and AI memory (1%).
A recurring theme in the less favourable reviews is poor AI understanding, which leads to irrelevant or partially correct responses. Some users show awareness of prompt engineering and attempt to guide the AI with examples, yet still encounter issues. For instance, review 3 in Figure 8 illustrates a user struggling to communicate effectively with the AI, despite providing clear examples. Similarly, review 4 in Figure 8 highlights inconsistent performance for complex requests. While the AI excels at handling simple tasks, it often struggles with more complex ones.
Although memory-related concerns (e.g., forgetting prior context or repeating content) are occasionally mentioned as shown in review 5 in Figure 8, they tend to be rated more leniently, indicating that users may be more forgiving of these limitations.
$\Rightarrow$ Implication 2.
While users are generally satisfied with the overall performance of Gen-AI apps, a portion of their concerns relate to limitations in human-AI communication, particularly when the AI struggles to interpret nuanced or complex prompts. To better meet user expectations and improve usability, developers should prioritize enhancing the Gen-AI’s ability to understand and respond appropriately to more complex interactions.
2) Content Quality. Content Quality refers to users’ assessment of AI-generated content in terms of usefulness, coherence, creativity, engagement, structure and presentation.
Figure 9. Examples of reviews discussing Content Quality
Insights. In our sample, 67% of reviews express satisfaction, praising the creative and polished outputs such as producing “radio-ready” songs and appending useful resources to textual answers as illustrated in reviews 1 and 2 in Figure 9 respectively.
However, 33% of reviews reveal concerns in three domains. First, users report that image-generation features frequently fail to produce accurate or faithful visual outputs, particularly for complex or creative prompts (e.g., failed attempts to depict two people and an eagle wearing denim shorts in reviews 3 and 4 in Figure 9, respectively.) This finding aligns with broader evidence that AI image generators often collapse under high-complexity tasks (Petsiuk et al., 2022). Second, 8% of reviews highlight concerns regarding insufficient diversity and inclusion. For instance, review 5 in Figure 9 highlights a disparity in image quality between indigenous women and people of colour. Third, a minority of users (3% of reviews) highlight formatting and user interface issues in text outputs, such as grammatical errors, excessive length, inconsistent structure, indicating a need for improved polish as illustrated in review 6 in Figure 9.
$\Rightarrow$ Implication 3.
Although users generally regard Gen-AI outputs as high quality, enhancing image accuracy, diversity, inclusion, formatting, and user interface presentation remains essential for delivering truly reliable and user-centered content. Developers should prioritize these areas to refine trust, inclusivity, and satisfaction across varied use cases.
3) Content Policy & Censorship. Content policy refers to rules and guidelines that define what type of AI-generated content is permissible. Censorship refers to automated or manual mechanisms used to block, modify, or suppress inappropriate, offensive, harmful, or illegal content based on those rules.
Figure 10. Examples of reviews discussing Content Policy & Censorship
Insights. In our review sample, 73% of users comment on censorship experience. Of these 23% report frustration with the filtering that often block benign content, adversely affecting usability. For example, the author of review 1 in Figure 10 reports getting censored for saying “hi.” On the other hand, 10% of the raised concerns relate to lack of effective filtering such as unethical imagery involving children bypassing filters as expressed in review 2 in Figure 10. Regarding adult content, 67% of concerned users express dissatisfaction with the excessive restriction of access. While 19% explicitly request an optional censorship toggle for adult users as illustrated in review 3 in Figure 10. Additionally, users express concern over unclear content ownership. Some users report that their creations were reused or uploaded by the platform without consent, as shown in review 4 in Figure 10.
$\Rightarrow$ Implication 4.
Developers need to strike the right balance in filtering content to avoid blocking genuinely harmful material while preventing harmless content. Ideally, developers could offer optional and adjustable censorship settings for mature or context-specific use. At the same time, platforms must clearly state who owns AI-generated outputs and how they can be reused, to maintain transparency and protect user trust.
4) Features & Functionality. Features & Functionality refer to the Gen-AI-driven capabilities offered by the app, such as voice interaction and prompt autocompletion, which shape the user experience and determine which features users find valuable or lacking.
Insights. In our sample, 75% of user reviews praise specific Gen-AI features, while 25% criticize them. For example, some users react negatively when a “stop-generation” button is removed and replaced with a voice mode feature, as illustrated in review 1 in Figure 11. While 11% of users enjoy the voice features, some describe them as useless. This indicates that even well‑received Gen-AI additions (such as voice interaction) can alienate users if they disrupt existing valued controls.
$\Rightarrow$ Implication 5.
Developers should design Gen-AI features, such as voice mode, in a way that complements existing controls rather than replacing them. For instance, it is crucial to ensure new AI-driven tools coexist with familiar functionalities, so that developers can both innovate and maintain intuitive user experiences.
Figure 11. Examples of reviews discussing Features & Functionality
5) Utility & Use Cases. Utility & Use Cases refer to the practical roles Gen-AI plays in users’ daily lives, including enhancing accessibility, supporting professional tasks, and enabling creative expression across diverse contexts.
Figure 12. Examples of reviews discussing Utility & Use Cases
Insights. In our sample, 64% of reviews emphasize AI’s educational utility, such as helping with homework, essay writing, tutoring, and breaking down complex topics as shown in review 1 in Figure 12. Additionally, some users specifically praise how Gen-AI can act as an accessible on-demand learning assistant for dyslexic users, as commented in review 2 in Figure 12. Beyond education, 11% of reviews describe how the Gen-AI tools facilitate routine work tasks, such as inspiring new ideas, structuring project workflows, and even acting as fitness or guitar instructors, as illustrated in review 3 in Figure 12.
$\Rightarrow$ Implication 6.
Despite the launch of tools like ChatGPT Edu (OpenAI, 2024a), which is tailored for academic contexts, Gen-AI platforms must go further by prioritizing accessibility-first design, for example, enabling assistive note-taking, voice transcription, and dyslexia support features. Developers should also integrate customizable, role-based modes or templates, such as “legal advisor,” “fitness coach,” or “creative brainstorming partner,” to make these tools more adaptable and effective across diverse professional and personal applications.
6) Personalization & Customization. Personalization & Customization describe how Gen-AI systems adapt to user preferences and allow users to tailor outputs. This includes customizing output formats, editing generated content, and adopting preferred AI personas, e.g., voices or AI character roles.
Figure 13. Examples of reviews discussing Personalization & Customization
Insights. In our sample, 38% of reviews emphasize the AI’s inability to create a personalized experience. For example, review 1 in Figure 13, discusses the app’s inability to maintain shopping and to-do lists. Additionally, 10% of users request inline editing capabilities to support customization by refining generated content directly, rather than having to regenerate entirely. For instance, review 2 in Figure 13 suggests the option to edit lyrics in an already generated song rather than regenerate a new one entirely. Meanwhile, 9% of reviews request more diversity in AI characters and custom voices. This reflects a growing interest in personalizing AI identities to match preferences such as female voice options. For example, review 3 in Figure 13 describes the absence of customized voice features as a “deal breaker.”
$\Rightarrow$ Implication 7.
Developers should implement customizable output options, enabling users to format, export, and save AI-generated content (e.g., lists, summaries) in a manner that suits their workflows. They should also add inline editing tools, so users can refine content directly without needing to regenerate responses. Additionally, offering persona customization, such as selectable character roles or voice styles, will enhance user engagement and make interactions feel more personalized.
7) Comparison to Other Apps. Comparison to Other Apps refers to comparing the AI app to other tools users have previously tried.
Figure 14. Examples of reviews discussing Comparisons to Other Apps
Insights. We find that 66% of the reviews draw comparisons to other AI-powered apps (e.g., comparing Microsoft Bing Search with Google). We also observe that 9% of the reviews compare the app with non-Gen-AI alternatives. For instance, review 1 in Figure 14 criticizes the forced integration of the AI chat feature, expressing a strong preference for traditional search functionality. Finally, 18% of reviews draw comparisons between the apps they review and general-purpose AI tools (e.g., ChatGPT), highlighting expectations shaped by user experiences with AI tools, as shown in review 2 in Figure 14. 77% of the reviews that compare the app with ChatGPT believe that ChatGPT outperforms the app in question, while 15% prefer the performance of the domain-specific app. These comparisons highlight a key challenge for domain-specific Gen-AI app developers. Their apps must offer distinct value that goes beyond the general-purpose functionality of ChatGPT.
$\Rightarrow$ Implication 8.
ChatGPT is a baseline comparator to users, in order to succeed developers must offer value beyond what ChatGPT offers. Developers should also consider creating a toggle button to allow users to opt in or out of AI-enabled features, such as AI search. The regular features may be sufficient for users (e.g., regular search can satisfy users’ searching needs, and AI search is not always necessary).
8) Customer Support & Community. Customer Support & Community refers to reviews that address developers directly to express their experience, such as reporting user enjoyment with specific features or requesting to fix app shortcomings.
Figure 15. Examples of reviews discussing Customer Support & Community
Insights. In our sample, we find that 51% of users express enjoyment with specific features (e.g., review 1 in Figure 15), 19% suggest new features to enhance their experience (e.g., review 2 in Figure 15), 9% request to fix app shortcomings (e.g., review 3 in Figure 15), and 21% discuss generic enjoyment of the app without relating it to a specific feature.
$\Rightarrow$ Implication 9.
User reviews highlight key features that developers should preserve when updating their apps. Prior work provided approaches that facilitate competitor analysis (Assi et al., 2025a; Hassan et al., 2022; Assi et al., 2021b). For competing developers, analyzing user feedback from similar apps can also uncover desirable features worth adopting to remain competitive in the market.
9) Human AI Interaction & Emotional Connection. Human AI Interaction & Emotional Connection reviews highlight the emotional and relational dimensions of human–AI interaction, particularly with conversational Gen-AI apps.
Figure 16. Examples of reviews discussing Human AI Interaction & Emotional Connection
Insights. We observe that users enjoy their interactions with conversational Gen-AI apps for multiple reasons. First, users enjoy a realistic interaction (i.e., closely similar to human dialogue). This perceived realism fosters a sense of companionship and emotional connection (e.g., review 1 in Figure 16). Second, reviews highlight the comfort in the emotional support provided by Gen-AI apps, particularly because the interaction lacks real human judgment, as shown in review 2 in Figure 16. This illustrates the importance of AI’s perceived neutrality and non-judgmental nature, especially when discussing sensitive or emotionally charged topics. Finally, users point to the therapeutic potential of conversational AI in providing mental and emotional support, such as managing mental health challenges (e.g., review 3 in Figure 16) and developing strategies to achieve personal goals.
$\Rightarrow$ Implication 10.
The perception of conversational Gen-AI apps as emotionally supportive and non-judgmental suggests their growing role as trusted companions, particularly in contexts involving emotional vulnerability. Developers should account for these evolving user expectations by ensuring that Gen-AI interfaces maintain a tone of empathy and neutrality, while also incorporating safeguards to avoid over-reliance or the illusion of professional psychological care. Furthermore, this trend highlights the need for interdisciplinary collaboration with mental health professionals to responsibly shape the affective design of Gen-AI systems used for emotionally sensitive interactions.
10) Accessibility & Inclusivity. Accessibility and inclusivity refer to the degree to which the Gen-AI apps accommodate users for affordability, multilingual support, age appropriateness, and usability for individuals with cognitive or physical impairments.
Figure 17. Examples of reviews discussing Accessibility & Inclusivity
Insights. Users discuss different accessibility and inclusivity topics including: 1) financial barriers (raised in 31% of the reviews, such as review 1 in Figure 17), 2) greater language support (raised in 22% of the reviews, such as review 2 in Figure 17), 3) disability-related needs (raised in 8% of the reviews, such as reviews 3 and 4 in Figure 17), 4) age-related concerns (raised in 10% of the reviews, such as Reviews 5 and 6 in Figure 17), and 5) general accessibility concerns (raised in 31% of the reviews).
$\Rightarrow$ Implication 11.
These findings highlight the importance of designing Gen-AI apps with diverse accessibility needs in mind. Developers should prioritize inclusive design by integrating features such as read-aloud support, customizable text presentation, and adaptive interfaces tailored to neurodiverse users (e.g., Attention Deficit Hyperactivity Disorder, Dyslexia and Dyspraxia).
Summary of RQ 2 From our analysis of user reviews across diverse Gen-AI app categories, we find that users are increasingly experimenting with prompts, offering feedback, and forming emotional connections with non-human agents. They expect fast, creative, and context-aware responses, along with personalization and transparency. For developers, this means moving beyond baseline functionality to support user control (e.g., ability to disable Gen-AI features) and differentiated value compared with AI tools (e.g., ChatGPT).
6. RQ3: What temporal trends emerge in user feedback on Gen-AI topics?
To understand how users perceive Gen-AI features over time, we explore the temporal dynamics of user feedback and ratings in app reviews. Our goal is to uncover how user opinions about Gen-AI topics shift as Gen-AI capabilities mature and become more integrated into mobile apps.
6.1. Experiment Setup
We track the evolution of user feedback over time by conducting a two-level trend analysis. First, we conduct a high-level comparison of overall trends between Gen-AI and non-Gen-AI topics by examining how 1) average ratings and 2) the percentage of reviews evolve over time. This perspective helps highlight the increasing user attention to Gen-AI features. Second, we perform a more fine-grained, category-level analysis of Gen-AI topics to uncover detailed temporal patterns. We track the evolution of individual Gen-AI topic categories using clustering techniques and qualitative review analysis. Based on the observed rating dynamics, we select three representative topic categories for deeper investigation: Content Quality (decreasing trend), AI Performance (stable trend), and Content Policy & Censorship (increasing trend). The following steps outline how we conduct the analysis:
Step 1: data preparation. We group reviews according to their topic category $(T)$ and time period $(P)$. Specifically, for every topic category $T$, we organize reviews into half-year intervals (i.e., every 6 months) based on their submission dates. Then, for each topic category $T$, we construct two time series across all time periods $P$: 1) $Avg_{(T,P)}$ as the average rating of reviews in category $T$ posted during period $P$, and 2) $Pr_{(T,P)}$ as the percentage of reviews in period $P$ that belong to category $T$. We exclude the topic category Accessibility & Inclusivity because it only contains 96 reviews in total, which are not enough to conduct a time series analysis.
Step 2: clustering methodology. To identify patterns in how Gen-AI topics evolve over time, we cluster the topics based on their trends. For each time series $Avg_{(T,P)}$ and $Pr_{(T,P)}$, we extract slope and standard deviation as features following the three-step process listed below:
  1. 1.
    Features extraction. Feature selection in clustering involves identifying the most relevant features to uncover patterns in unlabeled data (Dash and Koot, 2009). Feature selection enhances interpretability, aligns with research goals  (Kreienkamp et al., 2025), and enables grouping of topics with similar temporal trends (e.g., increasing, fluctuating, or stable). For each topic category $(T)$, we compute the slope $m$ and the standard deviation $\sigma$ for both $Avg_{(T,P)}$ and $Pr_{(T,P)}$. Selecting these features supports more meaningful clustering outcomes by capturing both the directional trend $m$ and variability $\sigma$ of each topic’s trend. It further enables the grouping of topics that exhibit similar temporal behaviours, such as consistently increasing or decreasing and fluctuating or stable patterns.
  2. 2.
    Clustering algorithm. K-Means clustering is a partitioning method that groups unlabeled data into ($k$) clusters by minimizing within-cluster variance (Warren Liao, 2005). We apply K-Means clustering to group topics based on their slope $m$ and standard deviation $\sigma$ values. We select K-Means because it is efficient, flexible, and widely used for clustering in data mining tasks (Ikotun et al., 2023; Assi et al., 2025b). K-Means is suitable for low-dimensional feature spaces (Ikotun et al., 2023) and hence enables us to identify generalizable trend patterns across different topics and apps.
  3. 3.
    Optimal number of clusters. To determine the optimal number of clusters ($k$), we compute Silhouette Scores, one of the most widely adopted methods for evaluating clustering quality (Ashari et al., 2023), across multiple values of $k$. Based on the silhouette results, we select the $k$ value closest to 1. The optimal silhouette scores correspond to $k$=3 for both the average rating trend series $Avg_{(T,P)}$ and the review percentage trend series $Pr_{(T,P)}$. We complement this with qualitative validation through manual inspection, confirming that $k$=3 yields the most interpretable and distinct groupings.
Step 3: manual trend investigation. To interpret the identified clusters, we perform qualitative analysis on random review samples from key time periods. For each selected topic category, we identify three representative intervals, typically from the beginning, a period of fluctuation, and the end of the timeline. From each interval, we extract a statistically representative random sample of reviews using a 95% confidence level and a 10% margin of error. We manually code the sampled reviews to identify recurring user concerns, behavioural patterns, and shifts in expectations. This process provides qualitative context that enhances our understanding of the observed quantitative trends.
6.2. Findings
Observation 3.1: Gen-AI topics consistently receive higher user ratings than non-Gen-AI topics, with an average difference of 0.8 rating stars over time. Specifically, the average rating for Gen-AI topics is 4.0, compared to 3.2 for non-Gen-AI topics. As illustrated in Figure 18, the percentage of reviews discussing Gen-AI topics increases over time. The percentage of Gen-AI reviews increases by 25%, from 18% in 2022-H1 to 43% in 2024-H2, reflecting a growing user focus on Gen-AI features. Despite this surge in volume, the average rating for Gen-AI topics remains stable over time, as shown in Figure 19 with only a slight change of ± 0.2 stars across the observed period. Upon further investigation, we find that this stability is the result of diverging trends across individual topic categories. As illustrated in Figure 20, some topic categories (e.g., Content Policy & Censorship) exhibit increasing trends, others (e.g., Content Quality) decline, while several topics (e.g., AI Performance) remain stable. These opposing directional trends counterbalance each other when aggregated, leading to the observed overall stability in Gen-AI topic ratings.
Figure 18. Evolution of the Percentage of Gen-AI Reviews
Figure 19. Evolution of the Average Ratings of Gen-AI vs Non-Gen-AI Topics
Gen-AI topics analysis. As established in our experimental setup, $k=3$ yields the most interpretable and well-separated clusters. These clusters reflect three general trends: increasing, decreasing, and stable as shown in Figure 20 and 21 for the average ratings and the percentage of reviews, respectively.
To deepen our understanding of the observed rating trends, we perform a manual trend investigation, i.e., a qualitative analysis, on three representative topic categories, each corresponding to a distinct trend direction. For the Decreasing trend, we select Content Quality due to the unexpected nature of its decline, particularly in light of ongoing advancements in AI capabilities as shown in Figure 20. For the Stable trend, we examine AI Performance, where we had anticipated an upward trend reflecting improvements in system performance. For the Increasing trend, we analyze Content Policy & Censorship to investigate how user perceptions of policy changes have evolved and to uncover the factors contributing to rising satisfaction in this area.
Figure 20. Trend Clusters for Average Ratings of Gen-AI Topics
Figure 21. Trend Clusters for Percentage of Gen-AI Topics
Observation 3.2: The Content Quality topic category shows a decline in rating over time, despite improvements in Gen-AI technology. As shown in Figure 20, the average rating for Content Quality decreases from 4.5 in 2022-H1 to 3.3 in 2024-H2. To better understand the unexpected downward trend in Content Quality ratings, we conduct a manual analysis of user reviews across three strategically selected time periods: 2022-H1 (corresponding to the highest average rating), 2022-H2 (marking a period of sharp decline), and 2024-H2 (representing the lowest observed average rating). This sampling allows us to investigate user behavior before, during, and after the major shifts in rating averages. Based on the manual inspection, we identify two key user-related factors driving the decrease as follows.
Figure 22. Examples of reviews discussing Content Quality at different time periods.
Observation 3.3: Over time, the Content Quality reviews become less tolerant of flaws and more likely to penalize them. Excitement for emerging Gen-AI technologies appears to influence early adopters’ evaluation criteria, making them less likely to penalize imperfections. In our analysis of the 2022-H1 sample, 96% of the reviews are positive, with only 5% of them including constructive feedback despite assigning a perfect 5-star rating. For example, the user writing review 1 as shown in Figure 22 awards the app the highest rating while reporting that the AI generated low-quality facial images. Such reviews reflect a pattern in which users prioritize innovation and novelty over technical quality, demonstrating a higher tolerance for flaws during the early adoption phase.
Over time, users become less tolerant of imperfections. The percentage of positive reviews declines from 96% in 2022-H1 to 59% in 2024-H2, indicating a significant shift in user ratings. Moreover, positive reviews become increasingly critical: in 2022-H2, 12% of all reviews offer constructive feedback while still assigning a high rating (4 or 5 stars), rising to 18% in 2024-H2. This trend suggests that users’ expectations have heightened.
More notably, an increasing number of users begin deducting rating star points for flaws over time. In 2022-H2, only 2% of constructive reviews assign a 5-star rating, while 10% assign a 4-star rating. By 2024-H2, this trend intensifies: 15% of constructive reviewers assign a 4-star rating, whereas only 4% assign a 5-star rating. For instance, review 2 in Figure 22 expresses the need for improvement in graphics and gives a 4-star rating. This shift presents a growing criticality among users. While early reviews tend to overlook shortcomings in favour of innovation, later reviews are more discerning, penalizing points even when expressing generally positive sentiment.
Observation 3.4: Over time, the Content Quality reviews are more experimental and demanding. Initial reviews frequently convey surprise and satisfaction with relatively modest capabilities, indicating limited expectations during the early adoption phase. For example, review 3 in Figure 22 from 2022-H1 expresses surprise at the high quality of the content, despite initially having low expectations. Such statements suggest that early positive ratings are often driven by low expectations of the novel technology rather than its actual performance.
As Gen-AI technologies become more widespread, user expectations rise accordingly. Reviews begin to reflect more ambitious and exploratory usage patterns, with users increasingly testing the system’s capabilities through complex prompts. When the app fails to meet these heightened demands, users express dissatisfaction. For example, review 4 in Figure 22 in 2022-H2, reports success in simple requests but horrific results with complex requests, ultimately giving the app a poor 1-star rating. This evolution in review content presents a shift from initial awe to critical evaluation, as users become more experimental and less tolerant of limitations.
We refer to users who review the app in 2022 as early adopters because 2022 is the year ChatGPT was launched (Akhtar, 2024), and we refer to users in 2023 and 2024 as mainstream users as they started using the app as it rose in popularity. Early adopters are typically motivated by innovation and are more tolerant of uncertainties and imperfections in emerging technologies than mainstream users (Paulino and Gudmundsson, 2024). Hence, the decline in the average rating of Content Quality can be explained by the change in observed characteristics between early adopters and mainstream users.
$\Rightarrow$ Implication 12.
Developers entering a competitive Gen-AI market must ensure that their apps are robust, possess refined capabilities and are capable of meeting the higher expectations of mainstream users. Over-promising performance can lead to user dissatisfaction and reputational harm, whereas transparent communication about capabilities and limitations can foster trust.
Observation 3.5: The AI Performance topic exhibits a stable trend in average user ratings over time, characterized by a slight increase in 2023-H1, and followed by a modest decline in 2024-H1. To account for this overall stability, we analyze fluctuations at the app category level. We identify a compensatory pattern revealing that improvements in user-perceived AI Performance in some categories are offset by declines in others. This inter-category balancing effect contributes to the relatively flat trend observed in the aggregate ratings.
Figure 23. Examples of reviews discussing AI Performance at different time periods.
Observation 3.6: Advances in image generation capabilities contribute to the slight increase in the average rating of the AI Performance topic in 2023-H1. In 2022-H1, negative reviews of image generation apps constitute 9% of reviews in our sample and account for 47% of all negative reviews during that period. These users express frustration with the visual output. For example, review 1 in Figure 23 expresses frustration over the performance of the image generator.
By 2023-H1, these issues appear to be addressed, as the percentage of negative reviews of image generation apps drops to 5%, and their share of total negative reviews decreases to 29%. This reduction in dissatisfaction likely suggests improvements in the AI Performance of image generation, which helps lift the overall AI Performance rating during this period.
Observation 3.7: Underperformance in Productivity apps leads to a slight decline in the average rating of the AI Performance topic in 2024. By 2024-H1, Productivity apps emerge as the leading source of performance-related complaints, accounting for 8% of all reviews in our sample and 44% of all negative reviews. Unlike creative tools, which are often evaluated in terms of novelty or expressiveness, Productivity apps are frequently judged against well-established, non-AI alternatives that offer straightforward functionality. Consequently, users express notable frustration with AI-driven features that are perceived as unnecessarily complex or unreliable. For example, review 2 in Figure 23 prefers a regular assistant over an AI assistant because the AI assistant complicates simple tasks, and review 3 in Figure 23 prefers a calculator over an AI due to accuracy issues.
These comments highlight that Gen-AI integration can complicate tasks that would otherwise be simple. Rather than enhancing the user experience, AI features may undermine usability if they are poorly aligned with user expectations or fail to deliver practical value. Some reviews explicitly question the necessity of AI features. Review 2 in Figure 23 describes abandoning their AI assistant after it offered a detailed classification of fan technologies instead of simply turning on a fan. This indicates that forced integration may backfire, especially when it disrupts core functionality.
Observation 3.8: The stability in the average rating of the AI Performance topic is driven by offsetting trends across app types. Some app categories show improvement (e.g., image generation), while others show decline (e.g., Productivity). These shifts tend to cancel each other out at the global level. In 2023, rising satisfaction with creative apps contributes to a mild increase in performance ratings. In 2024, dissatisfaction in Productivity apps is balanced by relatively stable perceptions in creative and educational apps. This compensatory effect explains the flat trend in average performance ratings over time.
Although the global average for AI Performance ratings remains stable over time, this apparent stability masks meaningful variation across app categories. Our analysis reveals divergent trends: categories such as Art & Design, Entertainment, and Books & Reference show moderate declines in average AI Performance ratings (–0.84 to –0.93), while others like Education and Games experience notable increases (+1.39 and +0.96, respectively). These opposing shifts suggest that stability at the aggregate level may result from the balancing of improvements and declines across distinct domains. This trend may also reflect differences in user expectations. Creative tools are often evaluated for imagination and output quality, while Productivity apps are judged more on efficiency and reliability.
$\Rightarrow$ Implication 13.
AI integration should be purpose-driven and context-aware. Raji et al. (Raji et al., 2022) discussed how AI systems are often deployed without sufficient consideration of their actual functionality and utility. The authors argue that the assumption of AI’s inherent value can lead to ineffective or even harmful implementations, stressing the need for critical assessment before integration. Our findings further compel developers to critically assess whether Gen-AI meaningfully enhances a given task, rather than assuming its value by default. Inappropriate or unnecessary use of AI can degrade usability and frustrate users.
Observation 3.9: The average user ratings for the Content Policy & Censorship topic follow an increasing trend, with fluctuations in 2022-H2 and 2024-H1. We observe that user ratings decrease in 2022-H2, rise steadily through 2023-H2, and decrease slightly in 2024-H1. To understand this trend, we examine the raised user concerns in three random samples of user reviews in 2022-H2, 2023-H2, and 2024-H1. We find that content policy complaints discuss three main concerns: 1) ethical AI training practices, 2) the effectiveness of content filters, and 3) false-positive censorship.
Figure 24. Examples of reviews discussing Content Policy at different time periods.
Observation 3.10: Ethical concerns around AI-generated art are more prominent in early samples but disappear in later ones. In 2022-H2, 10% of sampled reviews express concerns about the ethics of training AI on copyrighted human art without consent or compensation. For example, review 1 in Figure 24 expresses how Gen-AI harms artists whose art is stolen. This concern disappears from the random samples of 2023-H2 and 2024-H1.
This highlights the importance of transparent and ethical data use policies. However, the absence of similar comments in later samples suggests either that the user-base shifted or that these concerns were deprioritized in users’ feedback. The drop may also indicate user desensitization or limited platform changes aimed at addressing the issue. Nevertheless, the ethics of AI content policy should still be addressed by researchers and developers.
Observation 3.11: Content filters improve over time in blocking unwanted adult content (i.e., the raised concerns drop from 12% to 3%) and illegal content (i.e., the raised concerns drop from 4% to 0%). In our early sample, 15% of the studied reviews flag harmful content slipping past moderation, especially involving sexual or predatory material. This drops to 5% in 2023-H2 and to 3% by 2024-H1, indicating increased user satisfaction with content safety mechanisms. Examples of such reviews can be found in reviews 2, 3, and 4 in Figure 24.
The proportion of users reporting the unintentional generation of adult content drops from 12% in 2022-H2 to 3% in 2024-H1. Reports of illegal or predatory content involving children drops from 4% to 0%. These declines reflect measurable improvement in censorship mechanisms for harmful content.
Observation 3.12: Complaints about false-positive censorship grow over time (from 35% to 59%), especially among users engaging in regular and appropriate use. As filters improve at blocking harmful content with time, users increasingly report unnecessary censorship during benign interactions. In 2022-H2, 35% of users in our sample request a less strict filter. This jumps to 71% in 2023-H2 and settles at 59% in 2024-H1.
While the proportion of users requesting more freedom to generate explicit or violent content decreases slightly (from 21% in 2022-H2 to 14% in 2024-H1), more users begin to complain about filters interfering with regular usage. For example, 6% of reviews in 2022-H2 flag such issues, rising to 30% in 2024-H1. Review 5 in Figure 24 shows an example of reviews discussing false-positive censorship, where the user reports receiving a guidelines warning for writing a non-explicit role-play scene. This suggests that while filters are more successful in removing harmful content, they are over-correcting and obstructing benign, creative, or functional uses. Over-correction can hinder the Gen-AI user experience, particularly in role-play and storytelling apps.
Observation 3.13: Despite an overall positive trend, over-correction in censorship filters causes a minor ratings setback in 2024. The average rating for Content Policy and Censorship follows an increasing trend over time, reflecting growing user trust in content moderation systems. As previously discussed, this trend is supported by a substantial decline in complaints about harmful content. Correspondingly, the share of positive reviews that discuss censorship rises from 4% in 2022-H2 to 51% in 2023-H2. This rise indicates that while many users enjoy the app enough to leave a high rating, they still express frustration with the limitations imposed by the censorship filters.
However, this progress introduces a new challenge of false-positive filtering in 2024-H1. As previously discussed, while moderation systems are successful in blocking inappropriate content, they begin flagging benign prompts as violations. This over-correction results in a moderate decline in the share of positive censorship-related reviews (20% in 2024-H1) and contributes to a slight drop in the average rating. Nevertheless, this dip does not reverse the overall gains. Ratings in 2024-H1 remain higher than in 2022-H2, suggesting that while usability concerns emerge, users still view the improvements in content safety as a net positive. This also highlights an important reality for developers; progress in content moderation is unlikely to be perfectly linear. Minor setbacks and fluctuations are to be expected as systems evolve and adapt to user needs.
$\Rightarrow$ Implication 14.
Developers must strike a careful balance between safety and creative freedom. While stricter filters reduce the occurrence of disturbing or illegal content, they increasingly disrupt legitimate use. To avoid over-correction, developers should implement modular moderation systems with customizable filters or toggles that can be adjusted by personal preference. Clear content policy guidelines and transparent communication about AI behaviour are also essential. Additionally, ethical considerations around AI training data should not be ignored by developers or researchers, even if fewer users express concerns. These considerations remain essential for maintaining long-term trust and ensuring platform responsibility.
Summary of RQ 3 Our analysis reveals that Gen-AI topics evolve differently over time with increasing, stable, or decreasing trends. Content Quality ratings steadily decline, not due to technological regression, but due to increased user expectations. In contrast, AI Performance shows a stable rating trend, driven by a balancing effect between improving and declining domains. Lastly, Content Policy & Censorship follows an overall positive trend, with improved content filters increasing user ratings. These findings reflect a maturing user base that is more experienced, discerning, and vocal. Our user base of Gen-AI apps raise the bar for what constitutes a high-quality Gen-AI experience.
7. Discussion
In this section, we reflect on the broader implications of our findings and provide targeted recommendations for six key stakeholder groups: 1) users, 2) platform owners, 3) developers, 4) testing tool designers, 5) requirements engineers, and 6) policymakers and regulators.
7.1. Implications for Users
Our analysis reveals that users are not passive recipients of Gen-AI technology but are increasingly engaged, expressive, and collaborative (Observation 3.3). Reviewers provide constructive feedback to co-create better experiences with developers. As Gen-AI becomes more embedded in users’ lives, particularly in the emotional or educational contexts, users come to expect empathy, personalization, and reliability (Implications 2, 7, and 10). This shift highlights the importance of treating users not merely as consumers, but as stakeholders who help shape and refine Gen-AI apps.
Furthermore, our findings highlight the need for further research into the evolving relationship between users and Gen-AI conversational agents. Users increasingly describe these systems as trusted companions or emotional supports. Understanding the psychological and social bonds users form will be crucial for guiding the ethical design, deployment, and regulation of AI in emotionally sensitive contexts (Implication 21).
7.2. Implications for Platform Owners
Platform owners must bear responsibility for ensuring that Gen-AI apps comply with ethical standards (Observation 3.10). As AI capabilities evolve rapidly, app store operators should implement both technical tools and regulatory frameworks that evaluate whether submitted Gen-AI apps meet baseline criteria for safety, transparency, and responsible data usage. This may include requiring apps to disclose whether AI is used, clarifying how outputs are generated, and verifying that training data sources are ethically obtained. By introducing certification processes or audit mechanisms for Gen-AI features (particularly in apps that claim educational, therapeutic, or professional utility), platforms can promote accountability and mitigate harm, fostering greater trust among users and developers alike.
App platforms play a critical role in shaping the quality of user feedback. Our results show that a large proportion of user reviews are non-informative (Observation 1.2), which limits their value for developers and researchers. Future research could explore how redesigning review interfaces might encourage more meaningful feedback. One possible approach is to prompt users to answer specific questions or rate distinct topics, such as usability, design, or content accuracy, separately from the overall rating.
7.3. Implications for Developers
Developers face growing pressure to meet user expectations that are increasingly shaped by leading models, such as ChatGPT (Implication 8). Users demand fast, accurate, and context-aware outputs across a wide range of use cases, including education, professional assistance, and emotional support (Implications 6 and 10). Our findings highlight several critical areas for improvement: fine-tuning censorship filters to avoid false positives (Observation 3.9; Implication 4), supporting accessibility and inclusivity concerns (Implication 11), and enabling more personalized and flexible user experiences (Implication 7). Developers should also consider offering modular features and customization options, such as adjustable censorship settings, adjustable AI functionalities, and adaptive content formats. Importantly, developers must think critically about when and how AI is used, ensuring that integration adds meaningful value, as not every problem or task requires an AI solution (Observation 3.7; Implication 13).
7.4. Implications for Testing Tool Designers
Our LLM-based framework achieves 91% accuracy in topic extraction and assignment (Observation 1.4), demonstrating the effectiveness of prompt design and data pre-filtering (Observations 1.1 and 1.3). These results point to new opportunities in automated review analysis, particularly in testing and quality assurance workflows. Designers of testing tools should consider incorporating LLMs for top topic extraction and assignment. Our methods can help uncover nuanced user feedback across languages, accessibility contexts, and time, significantly enhancing product refinement cycles.
7.5. Implications for Requirements Engineers
Requirements engineers can benefit from the structured insights our topic analysis approach provides. The ability to track how user expectations evolve by topic over time can inform roadmap planning and feature prioritization (Observation 3.3). Our analysis also reveals that users clearly articulate feature-specific frustrations and preferences, providing actionable guidance for interface design, feature retention, and refinement. Engineers should leverage such review data to validate or revise assumptions about user needs, especially in light of shifting demographics and rising expectations for Gen-AI-powered functionality.
7.6. Implications for Policy Makers and Regulators
The emergence of Gen-AI apps raises important questions about ethics, safety, and user protection. While some early users voiced concerns about data ethics and copyright in training data, these concerns faded in later samples (Observation 3.10), which may indicate normalization rather than resolution. Policy makers must ensure that ethical standards around content ownership, training transparency, and accountability are formalized and enforced (Implication 4). Furthermore, content moderation must strike a balance between safety and usability. Regulators should develop protocols to mitigate the harm caused by unregulated AI generations (Observation 3.11; Implication 4), particularly as Gen-AI becomes increasingly pervasive in sensitive domains such as education and emotional support.
8. Threats to Validity
Construct threats to validity refer to the risk that the measures we use in our study do not accurately reflect the concepts we aim to study. First, we use the average app review rating as a proxy for user satisfaction. However, this metric may not fully capture the sentiment expressed in the review text. For example, users sometimes assign high ratings while simultaneously reporting critical issues or limitations. To mitigate this threat, we supplement the quantitative rating data with manual qualitative analysis of review content, ensuring that nuanced sentiments, both positive and negative, are appropriately captured. Second, our methodology relies on LLMs to automatically filter and classify reviews. To control for potential biases introduced by the LLM-based automation, we manually annotate samples and our results demonstrate high agreement, i.e., 91% accuracy, between the LLM-generated classifications and human annotations, supporting the reliability of using LLMs for these tasks. Additionally, we measure inter-rater reliability to confirm the consistency and trustworthiness of the automated classifications.
Internal threats to validity concern the soundness of the procedures used to collect, analyze, and interpret the data. A key internal threat relates to the uncertainty surrounding the timeline for introducing the Gen-AI features app. While we manually verify that all selected apps currently support Gen-AI functionality, we cannot determine the exact date of integration. To reduce the risk of analyzing reviews unrelated to Gen-AI, we exclude all reviews prior to 2022, i.e., the year ChatGPT was released. While this filtering improves temporal alignment, it cannot fully guarantee that all remaining reviews pertain to post-Gen-AI functionality.
External threats to validity concern the ability to generalize the results. Our analysis focuses on publicly available user reviews from the Google Play Store, which may not fully reflect user experiences on other platforms, such as the Apple App Store. We take several steps to enhance the generalizability of our results. First, we include a diverse set of apps across multiple categories to ensure a wide range of topics and app functionalities. Second, although the selected apps may vary in implementation details from one platform to another, many of the Gen-AI topics discussed (i.e., Content Quality, Content Policy & Censorship, and Accessibility & Inclusivity) are prevalent across platforms and commonly adopted by competing apps within and across categories. Consequently, the insights derived from our analysis of user perceptions and feedback are likely to be transferable to other apps that implement similar Gen-AI capabilities. Developers of comparable apps, including those not represented in our dataset or operating on different platforms, may benefit from these findings. In particular, the results can help them better understand user concerns, inform feature design decisions, and mitigate risks associated with Gen-AI integration.
9. Conclusion
The rapid integration of Gen-AI features into mobile apps has transformed how users interact with technology. In this study, we present a user-centered analysis of Gen-AI mobile apps by examining over 676,000 reviews from 173 apps on the Google Play Store. Using our proposed four-phase SARA methodology and prompt-based LLM techniques, we uncover actionable insights into how users experience and evaluate Gen-AI functionalities.
Our results show that LLMs can reliably extract user-discussed topics with high accuracy, i.e., 91%, when supported by five-shot prompting and data cleaning. We identify the 10 most frequently discussed topics revealing user behavior and challenges that inform actionable insights for developers and researchers. Furthermore, our findings suggest an evolution in user expectations, with a shift in user behavior, such as shifting from early adopters who were more forgiving of limitations to a mainstream user base that demands higher expectations. Future work will explore cross-platform comparisons and further investigate how expectations vary across different user profiles (e.g., casual vs. expert users) as Gen-AI tools continue and mature.
Data Availability. We provide a replication package, including manually labelled data, and Python scripts to replicate the analyses at https://github.com/Safwat-UofT/Gen-AI_User_Reviews.
Acknowledgements.
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [RGPIN-2021-03969].

### v3

@@
 Despite widespread adoption, little is known about how end users perceive and evaluate these Gen-AI functionalities in practice.
-In this work, we conduct a user-centered analysis of 676,066 reviews from 173 Gen-AI apps on the Google Play Store.
-We introduce a four-phase methodology, SARA (Selection, Acquisition, Refinement, and Analysis), that enables the systematic extraction of user insights using prompt-based LLM techniques.
-First, we demonstrate the reliability of LLMs in topic extraction, achieving 91% accuracy through five-shot prompting and non-informative review filtering.
-Then, we apply this method to the informative reviews, identify the top 10 user-discussed topics (e.g., AI Performance, Content Quality, and Content Policy & Censorship) and analyze the key challenges and emerging opportunities.
+In this work, we conduct a user-centered analysis of 676,066 automatically labeled reviews from 173 Gen-AI apps on the Google Play Store.
+We propose a structured four-phase framework, SARA (Selection, Acquisition, Refinement, and Analysis), which integrates and extends state-of-the-art techniques for large-scale review collection, filtering, and analysis using prompt-based LLMs.
+First, we empirically validate the reliability of LLMs for topic extraction and assignment, achieving 91% accuracy through five-shot prompting and LLM-based filtering of non-informative reviews.
+We then apply the framework to informative reviews to identify the ten most discussed topics (e.g., AI Performance, Content Quality, and Content Policy & Censorship) and analyze the key challenges and emerging opportunities.
 Finally, we examine how these topics evolve over time, offering insight into shifting user expectations and engagement patterns with Gen-AI apps.
@@
 Generative AI, Large Language Models, User Review Analysis, Mobile Apps
-ccs: Software and its engineering Software maintenance tools
+$c$s: Software and its engineering Software maintenance tools
 1. Introduction
-Mobile applications (apps) are a crucial aspect of people’s lives, and their global data usage continues to rise each year (Rathod and Agal, 2023).
+Mobile applications (apps) are a crucial aspect of people’s lives, and their global data usage continues to rise each year (Rathod_Agal_2023).
 A major shift in the mobile app landscape occurred with the release of ChatGPT [^1] https://play.google.com/store/apps/details?id=com.openai.chatgpt&hl=en, which achieved 7.4 million downloads within its first ten days.
-This rapid adoption signalled strong consumer interest in generative artificial intelligence (Gen-AI) and sparked a surge of similar artificial intelligence (AI) powered apps across app marketplaces.
-In 2024, this trend contributed to a 112% year-over-year increase in downloads of AI chatbot apps (Ceci, 2025), such as Microsoft Copilot [^2] https://play.google.com/store/apps/details?id=com.microsoft.copilot and DeepSeek [^3] https://play.google.com/store/apps/details?id=com.deepseek.chat.
-Gen-AI refers to a class of AI that generates new and meaningful content, including text, images, audio and video (Feuerriegel et al., 2024; Zhang et al., 2024).
-Unlike traditional AI models, which rely on historical data to perform predictions and analytical tasks, Gen-AI models synthesize information from diverse sources to create original outputs (Gozalo-Brizuela and Garrido-Merchán, 2023).
-Among the most prominent forms of Gen-AI are large language models (LLMs), i.e., deep learning models trained on massive text corpora to understand and generate human-like language (Hadi et al., 2023).
+This rapid adoption signaled strong consumer interest in generative artificial intelligence (Gen-AI) and sparked a surge of similar artificial intelligence (AI) powered apps across app marketplaces.
+In 2024, this trend contributed to a 112% year-over-year increase in downloads of AI chatbot apps (Ceci2025), such as Microsoft Copilot [^2] https://play.google.com/store/apps/details?id=com.microsoft.copilot and DeepSeek [^3] https://play.google.com/store/apps/details?id=com.deepseek.chat.
+Gen-AI refers to a class of AI that generates new and meaningful content, including text, images, audio and video (Feuerriegel_Hartmann_Janiesch_Zschech_2024; 10628027).
+Unlike traditional AI models, which rely on historical data to perform predictions and analytical tasks, Gen-AI models synthesize information from diverse sources to create original outputs (gozalobrizuela2023surveygenerativeaiapplications).
+Among the most prominent forms of Gen-AI are large language models (LLMs), i.e., deep learning models trained on massive text corpora to understand and generate human-like language (Hadi_Tashi_Qureshi_Shah_Muneer_Irfan_Zafar_Shaikh_Akhtar_Wu_et_al_2023).
 LLMs power a wide range of natural language processing (NLP) tasks, including text generation, summarization, translation, and dialogue.
-Developers are increasingly integrating Gen-AI models into mobile apps to enable novel capabilities, such as conversational interfaces, voice interaction, and content generation (Hau et al., 2025).
-When embedded into apps, these models can deliver personalized and adaptive experiences by providing real-time, context-aware feedback that enhances user engagement and motivation (Yuen and Schlote, 2024).
+Developers are increasingly integrating Gen-AI models into mobile apps to enable novel capabilities, such as conversational interfaces, voice interaction, and content generation (Hau_Hassan_Zhou_2025).
+When embedded into apps, these models can deliver personalized and adaptive experiences by providing real-time, context-aware feedback that enhances user engagement and motivation (Yuen_Schlote_2024).
 A key enabler of this integration is the growing availability of Gen-AI-powered application programming interfaces (APIs).
-APIs serve as software intermediaries that enable mobile apps to access language understanding and generation functionalities without having to host or fine-tune models locally (Chen et al., 2025b).
+APIs serve as software intermediaries that enable mobile apps to access language understanding and generation functionalities without having to host or fine-tune models locally (Chen_Liu_Zhou_Zhao_Wang_Wang_Chen_Bissyandé_Klein_Li_2024).
 While the integration of Gen-AI into mobile apps offers clear benefits in functionality and user experience, it also introduces a range of technical and practical challenges.
-Implementing Gen-AI features requires consideration of API integration, prompt engineering, token limitation, inference cost, and the unpredictable model outputs (Chen et al., 2025a).
-These issues can degrade the quality of Gen-AI-powered apps, leading to functional failures (e.g., crashes, incorrect behaviour, or suboptimal user interfaces), reduced efficiency (e.g., latency or excessive token consumption), and potential security vulnerabilities (Shao et al., 2025).
+Implementing Gen-AI features requires consideration of API integration, prompt engineering, token limitation, inference cost, and the unpredictable model outputs (Chen_Gao_Chen_Zhang_Liu_2025).
+These issues can degrade the quality of Gen-AI-powered apps, leading to functional failures (e.g., crashes, incorrect behaviour, or suboptimal user interfaces), reduced efficiency (e.g., latency or excessive token consumption), and potential security vulnerabilities (Shao_Huang_Shen_Ma_Su_Wan_2024).
 Additionally, beyond implementation, quality assurance for Gen-AI-powered apps remains a concern.
-Unlike traditional software systems, Gen-AI components often lack formal specifications, making it challenging to evaluate correctness or performance in a deterministic way (Nahar et al., 2024).
+Unlike traditional software systems, Gen-AI components often lack formal specifications, making it challenging to evaluate correctness or performance in a deterministic way (Nahar_Kästner_Butler_Parnin_Zimmermann_Bird_2024).
 As a result, ensuring the reliability and robustness of Gen-AI-integrated mobile apps continues to be a significant challenge for developers.
@@
 For example, Nahar et al.
-(Nahar et al., 2024) employed interviews and surveys to investigate how developers address Gen-AI integration issues, including latency, energy consumption, cost, and fairness.
+(Nahar_Kästner_Butler_Parnin_Zimmermann_Bird_2024) employed interviews and surveys to investigate how developers address Gen-AI integration issues, including latency, energy consumption, cost, and fairness.
 While informative, their work is limited to developer perspectives and does not consider how these technical challenges impact end-user experience.
 Chen et al.
-(Chen et al., 2025a) analyzed posts from an OpenAI developer forum to uncover challenges in prompting, API use, and plugin development.
+(Chen_Gao_Chen_Zhang_Liu_2025) analyzed posts from an OpenAI developer forum to uncover challenges in prompting, API use, and plugin development.
 This study offers insight into implementation bottlenecks but also neglects user-facing concerns.
 Similarly, Shao et al.
-(Shao et al., 2025) examined open-source mobile apps with Gen-AI integration and found that 77% experienced integration challenges, which impacted functionality, performance, or security issues.
+(Shao_Huang_Shen_Ma_Su_Wan_2024) examined open-source mobile apps with Gen-AI integration and found that 77% experienced integration challenges, which impacted functionality, performance, or security issues.
 However, their analysis is limited to code-level integration problems and does not investigate whether users are aware of or affected by these issues.
 In contrast, our work examines the user-perceived impact of Gen-AI-enabled apps, using large-scale review data to understand how users perceive Gen-AI features in practice.
-Existing studies on user perspectives of Gen-AI apps tend to focus on specific domains, such as education (Golding et al., 2024; Kim et al., 2025b; Lee et al., 2024a; Shata and Hartley, 2025; Kim et al., 2025a; Yuen and Schlote, 2024), creative image generation (Tang et al., 2024; Haase et al., 2023; Oppenlaender et al., 2023), or mental health (Jin et al., 2025; Heinz et al., 2025).
+Existing studies on user perspectives of Gen-AI apps tend to focus on specific domains, such as education (Golding_Lippert_Neuschatz_Salomon_Burke_2024; Kim_Yu_Detrick_Li_2025; Lee_Arnold_Srivastava_Plastow_Strelan_Ploeckl_Lekkas_Palmer_2024; Shata_Hartley_2025; Kim_Klopfer_Grohs_Eldardiry_Weichert_Cox_Pike_2025; Yuen_Schlote_2024), creative image generation (Tang_Zhang_Ciancia_Wang_2024; haase2023art; Oppenlaender_Silvennoinen_Paananen_Visuri_2023), or mental health (Jin_Kim_Han_2025; Heinz_Mackin_Trudeau_Bhattacharya_Wang_Banta_Jewett_Salzhauer_Griffin_Jacobson_2025).
 Our work takes a systematic view by analyzing user reviews across a wide range of Gen-AI tools and application areas.
@@
 For example, Figure 1 illustrates representative reviews from three Gen-AI apps, highlighting both praise and frustration.
-Building on this foundation, we introduce a four-phase methodology, Selection, Acquisition, Refinement, and Analysis (SARA).
-Our proposed methodology enables the large-scale collection and processing of user reviews from Gen-AI apps, supporting effective analysis using LLMs.
+Building on prior work in user review mining and LLM-based topic analysis (Maram_LLM_Cure; Chen_Lin_Hoi_Xiao_Zhang_2014), we propose a structured four-phase framework called Selection, Acquisition, Refinement, and Analysis (SARA).
+This framework enables the systematic collection, cleaning, and analysis of user reviews from Gen-AI apps, demonstrating the feasibility of large-scale, reproducible studies using LLMs.
 First, we identify relevant Gen-AI apps (i.e., Selection) and collect user reviews from the Google Play Store (i.e., Acquisition).
-We then apply a data cleaning pipeline to enhance the quality of input data for LLM-based analysis (i.e., Refinement).
+We then apply a data cleaning framework to enhance the quality of input data for LLM-based analysis (i.e., Refinement).
 Finally, we employ prompt-based LLM techniques to analyze the reviews (i.e., Analysis).
-Using this methodology, we conduct an empirical study on 676,066 reviews from 173 Gen-AI apps to address the following research questions (RQs):
+Using this integrative framework, we automatically label and quantitatively analyze 676,066 user reviews from 173 Gen-AI apps using LLMs, complemented by qualitative analyses on statistically representative samples.
+Before analyzing Gen-AI user reviews (RQ2–RQ3), we first evaluate the accuracy and reliability of our LLM-based framework (RQ1) to establish its validity for large-scale analysis.
+This ensures that the topic extraction and assignment process used in the subsequent analyses is accurate and trustworthy.
 RQ1: How accurately can LLMs identify topics in user reviews of Gen-AI apps?
-{adjustwidth}
-0.8cm In this RQ, we investigate how prompt configuration (e.g., using 0-shot, 3-shot, and 5-shot) and input quality (e.g., the proportion of informative reviews) affect the performance of LLMs in extracting topics from user reviews.
+In this RQ, we investigate how prompt configuration (e.g., using 0-shot, 3-shot, and 5-shot) and input quality (e.g., the proportion of informative reviews) affect the performance of LLMs in extracting topics from user reviews.
 Our results show that increasing the number of few-shot examples and filtering non-informative reviews significantly improve accuracy.
-The best-performing setup achieves an overall accuracy of 91%, confirming the reliability of our LLM pipeline for downstream analysis.
-RQ2: What are the most prominent topics discussed in user reviews of Gen-AI apps?
-{adjustwidth}
-0.8cm To answer this RQ, we apply the LLM-based topic labelling pipeline from RQ1 to informative reviews across 173 Gen-AI apps.
+The best-performing setup achieves an overall accuracy of 91%, confirming the reliability of our LLM framework for downstream analysis.
+RQ2: What are the most prominent topics discussed in user reviews and in developer responses of Gen-AI apps?
+To answer this RQ, we apply the LLM-based topic labeling framework from RQ1 to informative reviews across 173 Gen-AI apps.
 We analyze the top 10 topics in Gen-AI app reviews to understand user behaviour, key challenges and opportunities, and derive actionable insights for developers and researchers.
+While our primary focus is on user reviews, we also include a small supplementary examination of developer responses to better contextualize how developers engage with Gen-AI feedback.
 Our findings reveal that users tend to rate Gen-AI topics positively and expect high performance, often providing constructive feedback from a collaborative mindset.
@@
 RQ3: What temporal trends emerge in user feedback on Gen-AI topics?
-{adjustwidth}0.8cm In this RQ, we explore how user feedback on Gen-AI topics evolves as the technology matures.
+In this RQ, we explore how user feedback on Gen-AI topics evolves as the technology matures.
 We cluster topics to identify evolutionary trends and support the analysis by qualitative review interpretation.
@@
 Our study provides valuable insights for app developers and researchers by uncovering how users perceive and evaluate Gen-AI features in mobile apps. Specifically, our contributions are as follows:
-  • We propose a four-phase methodology, SARA, for analyzing user reviews of Gen-AI apps.
-Applying this methodology, we conduct a large-scale empirical study of 676,066 reviews from 173 Gen-AI apps on the Google Play Store to 1) identify the most frequently discussed Gen-AI-related topics, 2) uncover user expectations and pain points, and 3) analyze temporal trends to reveal how user perceptions evolve.
-  • We evaluate the effectiveness of LLMs for topic extraction and assignment of user reviews, demonstrating that increasing the number of labelled examples and filtering out non-informative reviews significantly improves accuracy.
+  • We propose a structured four-phase framework, SARA, that builds upon established techniques in user review analysis and LLM-based topic modeling.
+Using SARA, we conduct a large-scale empirical study of 676,066 reviews from 173 Gen-AI apps on the Google Play Store to 1) identify the most frequently discussed Gen-AI-related topics, 2) uncover user expectations and pain points, and 3) analyze temporal trends to reveal how user perceptions evolve.
+While validated in the context of Gen-AI apps, SARA’s modular design makes it generalizable to analyzing user feedback in other mobile app domains and AI-integrated systems.
+  • We evaluate the effectiveness of LLMs for topic extraction and assignment of user reviews, demonstrating that increasing the number of labeled examples and filtering out non-informative reviews significantly improves accuracy.
   • We provide actionable feedback for developers to maintain competitiveness, for policymakers to explore the ethical concerns surrounding content policy and censorship, and for user requirement engineers to better understand user needs.
-  • We provide a replication package that includes the dataset of Gen-AI apps, three manually labelled samples for evaluating the filtering prompt, and the complete set of code we used to run all prompt-based experiments presented in this paper.
+  • We provide a replication package that includes the dataset of Gen-AI apps, three manually labeled samples for evaluating the filtering prompt, and the complete set of code we used to run all prompt-based experiments presented in this paper.
 We organize the remainder of this paper as follows.
@@
 Section 3 presents our methodology.
-Sections 46 describe the approach and the results obtained in our RQs.
+Sections 46 describe the experiment design and the findings obtained for each RQ.
 Section 7 synthesizes our key findings and implications of our work.
@@
 2. Related Work
-In our literature review, we examine three key areas: 1) the user perspectives of Gen-AI apps, 2) user review analysis techniques, and 3) studies on prompt design and evaluation.
+In our literature review, we examine two key areas: 1) the user perspectives of Gen-AI apps, and 2) user review analysis techniques.
 2.1. User Perspective of Gen-AI Apps
-The user’s perspective on utilizing Gen-AI tools in education has been well studied (Golding et al., 2024; Kim et al., 2025b; Lee et al., 2024a).
-Shata and Hartley (Shata and Hartley, 2025) used surveys grounded in the Technology Acceptance Model (TAM) and Social Cognitive Theory (SCT) to examine faculty attitudes, identifying trust as the key driver of Gen-AI adoption.
+The user’s perspective on utilizing Gen-AI tools in education has been well studied (Golding_Lippert_Neuschatz_Salomon_Burke_2024; Kim_Yu_Detrick_Li_2025; Lee_Arnold_Srivastava_Plastow_Strelan_Ploeckl_Lekkas_Palmer_2024).
+Shata and Hartley (Shata_Hartley_2025) used surveys grounded in the Technology Acceptance Model (TAM) and Social Cognitive Theory (SCT) to examine faculty attitudes, identifying trust as the key driver of Gen-AI adoption.
 Kim et al.
-(Kim et al., 2025a) compared student and faculty perceptions on the use of Gen-AI in learning.
+(Kim_Klopfer_Grohs_Eldardiry_Weichert_Cox_Pike_2025) compared student and faculty perceptions on the use of Gen-AI in learning.
 Their study revealed that students and faculty used Gen-AI at similar rates, while students found Gen-AI tools easier to learn.
 Both faculty and students believed that Gen-AI will have more negative than positive effects on learning.
-Yuen and Schlote (Yuen and Schlote, 2024) surveyed adult learners to explore their views on the integration of AI in language learning apps.
+Yuen and Schlote (Yuen_Schlote_2024) surveyed adult learners to explore their views on the integration of AI in language learning apps.
 Their participants expressed a desire for in-app social features to assist in learning a language, and the authors suggested AI integration in the form of AI tutors that act as conversation partners.
 In the creative image domain, Tang et al.
-(Tang et al., 2024) surveyed users of Gen-AI image tools focusing on satisfaction, challenges, acceptance, perceptions, and applicability of Gen-AI image tools.
+(Tang_Zhang_Ciancia_Wang_2024) surveyed users of Gen-AI image tools focusing on satisfaction, challenges, acceptance, perceptions, and applicability of Gen-AI image tools.
 They found that users most frequently rely on these tools during the early ideation stage rather than for producing final outputs.
 In support of that, Haase et al.
-(Haase et al., 2023) examined how Gen-AI images influence creative idea generation, and demonstrated that Gen-AI art effectively supports ideation.
+(haase2023art) examined how Gen-AI images influence creative idea generation, and demonstrated that Gen-AI art effectively supports ideation.
 Oppenlaender et al.
-(Oppenlaender et al., 2023) focused more on the user’s understanding of technology, fears, and thoughts on the risks of Gen-AI image tools in their survey.
+(Oppenlaender_Silvennoinen_Paananen_Visuri_2023) focused more on the user’s understanding of technology, fears, and thoughts on the risks of Gen-AI image tools in their survey.
 Their participants voiced concerns over a societal risk, as people could use these tools to generate images that spread misinformation.
-In mental health, applications of AI include detection, diagnosis, treatment, and ongoing condition management (Sharma and Patel, 2024; Madhuri et al., 2025; S et al., 2025; Kheterpal and Gill, 2024).
+In mental health, applications of AI include detection, diagnosis, treatment, and ongoing condition management (10842666; 10963316; 10985281; 10911418).
 Our focus is on Gen-AI-driven treatment through conversational agents.
 Jin et al.
-(Jin et al., 2025) reviewed barriers to sustained engagement in mental health apps and suggested integrating AI in the form of AI agents that can detect emotional changes in users, build strong relationships with users, and promote timely intervention.
+(Jin_Kim_Han_2025) reviewed barriers to sustained engagement in mental health apps and suggested integrating AI in the form of AI agents that can detect emotional changes in users, build strong relationships with users, and promote timely intervention.
 Heinz et al.
-(Heinz et al., 2025) studied the efficacy of a mental health AI chatbot.
+(Heinz_Mackin_Trudeau_Bhattacharya_Wang_Banta_Jewett_Salzhauer_Griffin_Jacobson_2025) studied the efficacy of a mental health AI chatbot.
 They found that it significantly reduced users’ symptoms, maintained high engagement, and achieved user acceptance, with a reported therapeutic alliance comparable to that of human therapists.
-Unlike prior work that relied on expert synthesis or self-reported perceptions, our study analyzes large-scale, real-world reviews from users actively engaging with Gen-AI apps.
-We offer a complementary perspective on actual usage and feedback.
-We also take a broad perspective by examining Gen-AI apps across all domains, without prioritizing any specific application area.
+Prior studies on user perceptions of Gen-AI apps have primarily relied on interviews and surveys to capture attitudes toward adoption, trust, and usefulness within specific domains such as education (Shata_Hartley_2025; Kim_Klopfer_Grohs_Eldardiry_Weichert_Cox_Pike_2025) and image generation (Tang_Zhang_Ciancia_Wang_2024; Oppenlaender_Silvennoinen_Paananen_Visuri_2023).
+These works offer valuable insight into how users conceptualize Gen-AI, but they often reflect processed opinions rather than immediate experiences.
+In contrast, our analysis of large-scale user reviews captures lived interactions, representing users’ spontaneous reactions written in context and often within moments of use.
+This perspective complements survey-based research by revealing practical concerns and responses as they naturally occur.
+For example, while prior studies in mental health (Jin_Kim_Han_2025; Heinz_Mackin_Trudeau_Bhattacharya_Wang_Banta_Jewett_Salzhauer_Griffin_Jacobson_2025) demonstrated that conversational AI can foster engagement and therapeutic rapport, our findings extend this understanding by showing that users explicitly value Gen-AI’s perceived neutrality, emotional support, and human-like realism.
+Furthermore, while earlier work examined single domains in isolation, our study provides a holistic view across multiple app categories, revealing shared topics and concerns that transcend context.
+By organizing insights by functional features rather than by domain, our findings highlight common usability, quality, and ethical challenges that characterize Gen-AI technologies more broadly.
 2.2. User Review Analysis Techniques
-In this section, we summarize prior work related to 1) filtering user reviews and 2) topic extraction and assignment to user reviews.
+In this section, we summarize prior work related to 1) filtering user reviews, 2) pre-LLM approaches to topic extraction, 3) LLM-based topic extraction and assignment methods, 4) single-task topic analysis, and 5) domain-specific applications of LLMs in user review analysis.
+These studies collectively form the foundation for our own framework SARA, which builds upon their methods but advances them by integrating filtering, topic extraction, and topic assignment into a unified, scalable framework for analyzing Gen-AI app reviews.
 Filtering user reviews.
@@
 Chen et al.
-(Chen et al., 2014) introduced a framework called AR-miner that used Latent Dirichlet Allocation (LDA) to first filter informative reviews with a hit-rate of 70%, and second to group the informative reviews using topic modelling.
+(Chen_Lin_Hoi_Xiao_Zhang_2014) introduced a framework called AR-miner that used Latent Dirichlet Allocation (LDA) to first filter informative reviews with a hit-rate of 70%, and second to group the informative reviews using topic modeling.
 Ghosh et al.
-(Ghosh et al., 2024) evaluated fine-tuned LLMs for filtering non-informative user reviews on the Google Play Store and found them highly effective, achieving accuracies of 92% for both BERT and DistilBERT, and 91% for Gemma.
+(Ghosh_Pargaonkar_Eisty_2024) evaluated fine-tuned LLMs for filtering non-informative user reviews on the Google Play Store and found them highly effective, achieving accuracies of 92% for both BERT and DistilBERT, and 91% for Gemma.
 Al Wahshat et al.
-(Al Wahshat et al., 2023) highlighted the potential of using GPT-4 in filtering manipulated reviews due to its powerful language processing capabilities.
-Topic extraction and assignment to user reviews.
+(Al_Wahshat_Abu_ulbeh_Yusoff_Zakaria_Amir_Fazamin_Wan_Hamzah_P_2023) highlighted the potential of using GPT-4 in filtering manipulated reviews due to its powerful language processing capabilities.
+Building on this line of work, our study integrates automated filtering as the first step of a unified, LLM-assisted framework that combines filtering with topic extraction and assignment.
+While prior work primarily evaluated filtering accuracy in isolation, we demonstrate that incorporating this process within a broader analysis framework improves data quality and supports large-scale review interpretation.
+Our filtering component achieves comparable accuracy (90%) while operating at a much larger scale across diverse Gen-AI app categories.
+Pre-LLM approaches to topic extraction.
 Before the advancement of LLMs, researchers used various methods to analyze user reviews.
-Genc Nayebi and Abran (Genc-Nayebi and Abran, 2017) conducted a literature review in 2016 that explored approaches for mining opinions from user reviews.
+Genc Nayebi and Abran (Genc-Nayebi_Abran_2017) conducted a literature review in 2016 that explored approaches for mining opinions from user reviews.
 They identified three studies that relied on manual analysis, while others used automated techniques, such as Classification and Regression Trees (CART), word frequency statistics, correlation analysis, LDA, and Hexbin plots.
@@
 For example, Arambepola et al.
-(Arambepola et al., 2024) used the Valence Aware Dictionary and sEntiment Reasoner (VADER) for sentiment analysis to categorize reviews into positive, neutral, and negative categories.
-They then used an LDA model for topic modelling of user reviews of educational apps.
+(10499727) used the Valence Aware Dictionary and sEntiment Reasoner (VADER) for sentiment analysis to categorize reviews into positive, neutral, and negative categories.
+They then used an LDA model for topic modeling of user reviews of educational apps.
 Their positive topics achieved a coherence score of 0.56, and their negative topics achieved a score of 0.49.
 Assi et al.
-(Assi et al., 2021a) proposed the Global-Local sensitive Feature Extractor (GLFE) to identify high-level features in user reviews, achieving 79–82% precision and 74–77% recall on annotated review data.
+(Assi_Hassan_Tian_Zou_2021) proposed the Global-Local sensitive Feature Extractor (GLFE) to identify high-level features in user reviews, achieving 79–82% precision and 74–77% recall on annotated review data.
+These early techniques established the foundation for automated review analysis but relied heavily on statistical or lexicon-based models.
+With the advent of LLMs, topic extraction has shifted toward context-aware and semantically richer methods, as discussed next.
+LLM-based topic extraction and assignment methods.
 LLMs have demonstrated promising performance in tasks related to topic extraction and assignment.
 Arambepola et al.
-(Arambepola et al., 2025) conducted a recent systematic review of mobile app review analysis techniques and found that integrating LLMs into app user review analysis provides a more holistic and context-rich approach to evaluating reviews and understanding users.
+(Arambepola_Lalendra_Wimalasena_Munasinghe_2025) conducted a recent systematic review of mobile app review analysis techniques and found that integrating LLMs into app user review analysis provides a more holistic and context-rich approach to evaluating reviews and understanding users.
 Pham et al.
-(Pham et al., 2024) introduced TopicGPT, a framework to perform topic extraction by prompting Open AI’s GPT (OpenAI, 2023) to generate topics and assign them to documents.
+(Pham_Hoyle_Sun_Resnik_Iyyer_2024) introduced TopicGPT, a framework to perform topic extraction by prompting Open AI’s GPT (openai2023gpt) to generate topics and assign them to documents.
 Their framework outperformed traditional topic models, such as LDA, SeededLDA, and BERTopic.
 Prakash et al.
-(Prakash et al., 2023) introduced PromptMTopics, a prompt-based ChatGPT model set to extract topics from textual meme captions that outperformed baseline models.
+(Prakash_Wang_Hoang_Hee_Lee_2023) introduced PromptMTopics, a prompt-based ChatGPT model set to extract topics from textual meme captions that outperformed baseline models.
+While these studies evaluated LLM-based extraction on textual datasets such as documents or memes, our work applies similar techniques directly to large-scale user reviews.
 Assi et al.
-(Assi et al., 2025a) introduced LLM-Cure, an approach that extracts and assigns features from user reviews to make suggestions for feature improvement.
+(Maram_LLM_Cure) introduced LLM-Cure, an approach that extracts and assigns features from user reviews to make suggestions for feature improvement.
 Their approach outperformed baseline models, achieving an F1-score of 85%.
-Sorathiya and Ginde (Sorathiya and Ginde, 2024) employed a hybrid approach that combines Natural Language Inference (NLI) and LLMs to identify privacy concerns in mental health app reviews, aiming to uncover ethically relevant user feedback that traditional keyword-based methods often overlook.
-Their approach achieved an F1-score of 81%.
-Instead of extracting and assigning topics to user reviews, prior studies focused on a single task (i.e., either extracting topics from user reviews or assigning predefined topics to user reviews).
+Among these, LLM-Cure most closely resembles our framework in combining both extraction and assignment tasks, though its focus was on generating feature suggestions rather than large-scale trend analysis.
+Single-task topic analysis.
+Instead of extracting and assigning topics to user reviews, some studies focused on a single task (i.e., either extracting topics from user reviews or assigning predefined topics to user reviews).
 For example, Wei et al.
-(Wei et al., 2023) focused on review assignment using predefined labels rather than extracting topics.
+(Wei_Courbis_Lambolais_Xu_Bernard_Dray_2023) focused on review assignment using predefined labels rather than extracting topics.
 They assigned reviews using fixed categories such as feature requests, problem reports, and irrelevant feedback.
@@
 Dos Santos et al.
-(Dos Santos et al., 2023) extracted topics from user reviews without performing the assignment task.
-Their study examined the impact of app updates on accessibility by utilizing GPT-4 to identify key issues, barriers, improvements, update consequences, and user sentiments.
+(Dos_Santos_Oliveira_De_Jesus_Aljedaani_Eler_2023) and Ren et al.
+(Ren_Nakagawa_Tsuchiya_2024) extracted topics from user reviews without performing the assignment task.
 While this approach is effective for uncovering topical insights, it does not account for the frequency or distribution of each topic, limiting its usefulness for trend or prevalence analysis.
-Ren et al.
-(Ren et al., 2024) used LDA for topic extraction and pretrained LLMs for requirement extraction.
+Domain-specific applications of LLMs in user review analysis.
+Beyond general topic extraction, some studies applied LLMs to narrow domains of app analysis.
+Sorathiya and Ginde (Sorathiya_Ginde_2024) employed a hybrid approach combining Natural Language Inference (NLI) and LLMs to identify privacy concerns in mental health app reviews, uncovering ethically relevant feedback that keyword-based methods often miss.
 Rezaei Nasab et al.
-(Rezaei Nasab et al., 2025) conducted a study of AI-based mobile app reviews to identify fairness-related concerns.
-They constructed a labelled dataset of fairness vs.
-non-fairness reviews, developed machine learning and LLM-based classifiers to detect fairness mentions, applied K-Means clustering, and performed manual coding to identify six types of fairness concerns.
-Our work builds on this line of research by utilizing LLMs to filter out non-informative reviews and identify topics in user reviews of Gen-AI apps.
-Based on the analysis of user reviews, we identify the challenges and potential opportunities to adopt Gen-AI functionalities in mobile apps.
-2.3. Prompt Design and Evaluation
-Prompt engineering is the process of designing and refining prompts to effectively guide LLMs in performing specific tasks (Koyuncu, 2025).
-Prior work discusses prompt design in three directions: 1) the process of prompt design, 2) the definition of the required tasks, and 3) identifying the number of included examples.
-The process of prompt design.
-Marvin et al.
-(Marvin et al., 2024) emphasized the importance of prompt refinement and outlined several prompt engineering techniques, including dynamic prompting, which refers to iteratively adjusting the prompt based on the generated output.
-He et al.
-(He et al., 2024) investigated the impact of prompt format across various tasks and found that prompt structure significantly affects performance, with different formats performing best in different task types.
-In our work, we adopt a prompt format inspired by prior work (Assi et al., 2025a) and refine it to suit our study’s goals, employing dynamic prompting to enhance accuracy and relevance further.
-The definition of the required tasks.
-To ensure accurate, ethical, and reliable AI outputs, effective prompt engineering requires understanding common pitfalls, such as ambiguity, bias, and lack of context (Giray, 2023).
-Prompts can be explicit, implicit, or creative, varying in the amount of guidance they offer to the AI.
-Explicit prompts provide clear instructions, while implicit and creative prompts allow more flexibility and imagination but may produce less predictable results (Data Science Horizons, 2023).
-Ho et al.
-(Ho et al., 2024) adopted a prompt classification framework consisting of the three types.
-The authors analyzed how different prompts affect AI-generated responses.
-Explicit prompts generate specific and fact based answers, implicit prompts produce broader and more flexible responses, while creative prompts encourage innovation and originality.
-In our work, we adopt an explicit prompting strategy to improve structure and reproducibility.
-Identifying the number of included examples.
-Deshmukh et al.
-(Deshmukh et al., 2025) conducted a literature review of prompt engineering and discussed 16 LLM prompt patterns to enhance output accuracy.
-Their analysis highlighted the importance of clarity, specificity, defined output formats, and contextual examples to improve model accuracy.
-However, it did not address techniques for optimizing example selection (i.e., determining the optimal number of examples to include in the prompt).
-Complimenting this, Wang et al.
-(Wang et al., 2023) demonstrated that using similar examples and increasing the number to four enhances performance, while adding more than four examples yields only marginal gains.
-Their work focused on classification, multiple-choice, and generation tasks, rather than user review analysis tasks.
-Prakash et al.
-(Prakash et al., 2023) evaluated prompts with 2, 4, 6, and 8 examples, finding comparable topic coherence and diversity across configurations.
-In this paper, we further extend this line of research by investigating how the number of in-context examples and the exclusion of noisy data affect topic extraction and assignment tasks for user review.
-Our work demonstrates that increasing the number of examples for topic extraction from 0-shot to 5-shot significantly improves the model’s accuracy.
-Our work also suggests that filtering out non-informative reviews enhances the accuracy.
+(RezaeiNasab_Dashti_Shahin_Zahedi_Khalajzadeh_Arora_Liang_2025) examined fairness concerns in AI-based mobile apps, using LLM and machine learning based classifiers to detect fairness mentions and cluster them into six categories.
+While these studies reveal how LLMs can surface ethics-related insights within specific domains, our dataset and focus differ substantially.
+We analyze user reviews from a large-scale collection of Gen-AI apps spanning multiple functional categories, including but not limited to health & fitness, education, productivity, and entertainment.
+Unlike prior domain-specific datasets centered on single themes such as privacy or fairness, our corpus captures a broad range of emergent Gen-AI topics (i.e.
+AI performance, content quality, and content policy) that are unique to this new class of applications.
+This wider scope allows us to identify both category-specific differences and cross-domain similarities, showing how users across varied contexts often express comparable concerns and expectations about Gen-AI functionality.
+Collectively, these studies demonstrate the growing potential of LLMs for review analysis yet reveal a gap in unified, scalable approaches that integrate filtering, topic extraction, and topic assignment across diverse app domains.
+To address this, we design the SARA framework, which operationalizes these steps into a cohesive process for large-scale analysis of Gen-AI user feedback.
 3. Methodology
-To study user feedback on Gen-AI apps, we design a four-phase methodology we call SARA, which stands for Selection, Acquisition, Refinement, and Analysis.
-SARA encompasses 1) the selection of relevant Gen-AI apps, 2) the acquisition of user reviews, 3) a refinement data cleaning pipeline to ensure data quality for LLM analysis, and finally, 4) the LLM-based analysis of user reviews through topic extraction and assignment.
-Our proposed methodology enables us to address our RQs concerning the impact of data quality and few-shot examples on LLM performance, the prevalent topics within Gen-AI user reviews, and the temporal evolution of user ratings toward these apps.
-We present an overview of our research approach in Figure 2.
-Figure 2. Overview of our approach
+To study user feedback on Gen-AI apps, we propose and validate a structured four-phase framework we call SARA, which stands for Selection, Acquisition, Refinement, and Analysis.
+SARA encompasses 1) the selection of relevant Gen-AI apps, 2) the acquisition of user reviews, 3) a refinement data cleaning framework to ensure data quality for LLM analysis, and finally, 4) the LLM-based analysis of user reviews through topic extraction and assignment.
+SARA integrates and extends state-of-the-art methods in user review analysis (Maram_LLM_Cure; Chen_Lin_Hoi_Xiao_Zhang_2014), combining automated filtering, topic extraction, and topic assignment into a unified, empirically validated workflow that supports large-scale, LLM-driven analysis of user feedback.
+Although evaluated within the Gen-AI domain, SARA’s modular design enables its application to other mobile or AI-enabled systems for analyzing large-scale user feedback.
+We first evaluate the reliability of SARA in accurately labeling review topics (RQ1) and then apply it in subsequent analyses to explore user concerns and temporal trends (RQ2–RQ3).
+We present an overview of our research framework in Figure 2.
 3.1. Selection of Gen-AI apps
@@
     Keyword generation.
-We compile a list of keywords derived from a Gen-AI taxonomy by Gozalo-Brizuela and Garrido-Merchán (Gozalo-Brizuela and Garrido-Merchán, 2023).
+We compile a list of keywords derived from a Gen-AI taxonomy by Gozalo-Brizuela and Garrido-Merchán (gozalobrizuela2023surveygenerativeaiapplications).
 The authors organized Gen-AI apps into 14 categories, each with a taxonomy of the relevant technology.
@@
 We create a list of ten categories relevant to mobile apps and repurpose the taxonomy as keywords, as shown in Table 1.
+    Figure 2. Overview of our framework
   2. 2.
@@
 To ensure the inclusion of only genuine Gen-AI apps, the first and second authors independently review the apps by looking at the title, description, and images of the apps.
-We assess the inter-rater reliability, using Cohen’s Kappa (Cohen, 1960) ($\kappa=0.61$), which indicates substantial agreement.
+We assess the inter-rater reliability, using Cohen’s Kappa (cohen1960kappa) ($\kappa=0.61$), which indicates substantial agreement.
 In case of disagreement, the two authors discuss and resolve the discrepancies for the ambiguous cases.
@@
 3.2. Acquisition of User Reviews
-Following app selection, we collect user reviews using the Google Play Store Scraper (Mingyu, nd).
+Following app selection, we collect user reviews using the Google Play Store Scraper (googleplayscraper).
 For each retrieved review, the crawler extracts a comprehensive set of metadata, including reviewId, userName, userImage, content (i.e., review text), score (i.e., user rating from 1 to 5 stars), thumbsUpCount, reviewCreatedVersion, at (i.e., timestamp of the review), replyContent, repliedAt, and appVersion.
@@
 3.3. Multi-Stage User Review Refinement
-To ensure the robustness and reliability of our LLM-based analysis, we implement a multi-stage data cleaning pipeline designed to enhance the quality of the user review data progressively.
+To ensure the robustness and reliability of our LLM-based analysis, we implement a multi-stage data cleaning framework designed to enhance the quality of the user review data progressively.
 This multi-stage refinement process comprises four distinct stages.
@@
 Each stage builds upon the previous one as follows:
-Table 2. Number of reviews by app category and stage.
+Table 2. Number of reviews by app category and stage (Full dataset considered).
 App category Stage 0 Stage 1 Stage 2 Stage 3
@@
 Stage 0: basic cleaning.
-Text preprocessing, as demonstrated in prior work (Ghosh et al., 2024; Prakash et al., 2023; Roumeliotis et al., 2024; Lee et al., 2024b), is a standard practice for eliminating noise and ensuring more effective analysis.
+Text preprocessing, as demonstrated in prior work (Ghosh_Pargaonkar_Eisty_2024; Prakash_Wang_Hoang_Hee_Lee_2023; Roumeliotis_Tselikas_Nasiopoulos_2024; Lee_Lee_Seo_2024), is a standard practice for eliminating noise and ensuring more effective analysis.
 It involves the removal of punctuation, special characters, extra white spaces, emojis, and irrelevant data such as usernames and URLs.
@@
 To validate this threshold, we experiment with categories of varying sizes and find that categories with fewer than 1,000 reviews yield topics that lack diversity and informativeness.
-For instance, the Personalization app category, which contains only 23 reviews, produces two nearly identical and non-informative topics labelled Positive Feedback and Praise.
+For instance, the Personalization app category, which contains only 23 reviews, produces two nearly identical and non-informative topics labeled Positive Feedback and Praise.
 Additionally, some topics focus on non-informative content, such as Personal Comments or Signatures, which capture examples like “hello marlyn” and “Placau!!!!” that are not relevant to the app’s functionality or performance.
@@
 Some apps included in our initial screening are not originally designed as Gen-AI tools but have integrated Gen-AI functionalities at a later stage.
-For instance, in the case of Microsoft Bing Search [^4] https://play.google.com/store/apps/details?id=com.microsoft.bing, Bing Chat is a recently introduced feature that extends the capabilities of the search engine by incorporating Gen-AI models (Iorliam and Ingio, 2024).
+For instance, in the case of Microsoft Bing Search [^4] https://play.google.com/store/apps/details?id=com.microsoft.bing, Bing Chat is a recently introduced feature that extends the capabilities of the search engine by incorporating Gen-AI models (Iorliam_Ingio_2024).
 Consequently, it is essential to exclude user reviews published before the integration of such features, as they reflect experiences with the app before the adoption of Gen-AI.
-While foundational work on AI models began as early as 2012, notable advancements in Gen-AI only started to gain momentum after 2018 (Sengar et al., 2024).
-The year 2018 also marked a rise in academic interest and publications related to Gen-AI (Dwivedi and Elluri, 2024).
-However, the field witnessed a major breakthrough in 2022, particularly with the public release and widespread adoption of ChatGPT, which significantly accelerated the prominence and deployment of Gen-AI technologies across various platforms (Akhtar, 2024).
+While foundational work on AI models began as early as 2012, notable advancements in Gen-AI only started to gain momentum after 2018 (Sengar_Hasan_Kumar_Carroll_2024).
+The year 2018 also marked a rise in academic interest and publications related to Gen-AI (Dwivedi_Elluri_2024).
+However, the field witnessed a major breakthrough in 2022, particularly with the public release and widespread adoption of ChatGPT, which significantly accelerated the prominence and deployment of Gen-AI technologies across various platforms (Akhtar_2024).
 To maximize the relevance and accuracy of the user feedback analyzed in our study, we adopt 2022 as the cutoff year for including reviews.
@@
 Building upon the cleaned data from Stage 1, this stage employs a simple heuristic to further reduce noise by excluding reviews containing three words or less.
-Short reviews of three words or less do not contain context for meaningful topics (Assi et al., 2021a; Arambepola et al., 2024).
+Short reviews of three words or less do not contain context for meaningful topics (Assi_Hassan_Tian_Zou_2021; 10499727).
 This step aims to eliminate overly brief and potentially non-informative feedback.
 Stage 3: LLM-assisted informative review filtration.
-Non-informative reviews bring noise to user reviews analysis and influence the performance of topic extraction (Chen et al., 2014).
+Non-informative reviews bring noise to user reviews analysis and influence the performance of topic extraction (Chen_Lin_Hoi_Xiao_Zhang_2014).
 To improve the quality of extracted topics, we exclude non-informative reviews.
@@
 While Chen et al.
-(Chen et al., 2014) focused on informativeness from a developer’s perspective, we expand this definition to consider the informational needs of multiple stakeholders, including users, platform owners, developers and researchers.
+(Chen_Lin_Hoi_Xiao_Zhang_2014) focused on informativeness from a developer’s perspective, we expand this definition to consider the informational needs of multiple stakeholders, including users, platform owners, developers and researchers.
 We carry out the LLM-assisted process in the following steps.
@@
     Prompt design.
-To filter non-informative reviews, we utilize Open AI’s (OpenAI, 2023) GPT-4o-mini model with zero temperature for reproducibility.
-We use gpt-4o-mini instead of the full gpt-4o model because it offers significantly faster response times while maintaining reasonable output quality for our task, which does not require complex reasoning (OpenAI, 2024b).
-    We design a filtering prompt that instructs the LLM to classify reviews as either informative or non-informative.
-To guide the model effectively, we incorporate representative examples within the prompt.
+To filter non-informative reviews, we utilize Open AI’s (openai2023gpt) GPT-4o-mini model with zero temperature for reproducibility.
+We use gpt-4o-mini instead of the full gpt-4o model because it offers significantly faster response times while maintaining reasonable output quality for our task, which does not require complex reasoning (OpenAI_O4Mini_2024).
+    Prior research comparing explicit, implicit, and creative prompting (Giray_2023; Ho_Mayberry_Nguyen_Dhulipala_Pallipuram_2024; Deshmukh_Raut_Bhavsar_Gurav_Patil_2025) shows that explicit, well-structured prompts produce more factual and task-aligned outputs, while underspecified ones invite unpredictable reasoning.
+Consistent with these findings and with approaches such as LLM-Cure (Maram_LLM_Cure), we design an explicit, structured prompt that classifies each review as either informative or non-informative.
+The prompt includes representative examples, a fixed binary label set, and a CSV-style output format to constrain the model’s responses.
+These measures ensure deterministic labeling and minimize hallucination by preventing free-form or extraneous text generation.
   2. 2.
     Iterative prompt refinement.
+Prior studies highlight that prompt structure and iterative refinement (“dynamic prompting”) substantially influence LLM performance  (Marvin_Hellen_Jjingo_Nakatumba-Nabende_2024; He_Rungta_Koleczek_Sekhon_Wang_Hasan_2024).
+In our setup, we adopt this principle by refining prompts through multiple pilot runs to improve label consistency and clarity.
 We use small statistically significant samples of 96 reviews (95% confidence interval, 10% margin of error) from three diverse app categories (Photography, Productivity, and Art & Design).
 Then, the first and second authors manually label the reviews in the samples as informative or non-informative.
-We calculate the inter-rater reliability using Kappa (Cohen, 1960) ($\kappa=0.91$), which indicates almost perfect agreement.
-    Next, we design and test different versions of the filtering prompt against the manually labelled data, evaluating the model’s filtering accuracy and iteratively refining the prompt.
+We calculate the inter-rater reliability using Kappa (cohen1960kappa) ($\kappa=0.91$), which indicates almost perfect agreement.
+    Next, we design and test different versions of the filtering prompt against the manually labeled data, evaluating the model’s filtering accuracy and iteratively refining the prompt.
 Revisions include modifying the wording, structure, number of examples, and types of examples.
@@
 Effective topic extraction and assignment of user reviews are essential for understanding user needs and improving app features for Gen-AI apps.
-Prior work has shown that prompt engineering techniques can help tailor LLMs to specific tasks (Marvin et al., 2024).
+Prior work has shown that prompt engineering techniques can help tailor LLMs to specific tasks (Marvin_Hellen_Jjingo_Nakatumba-Nabende_2024; 10.1145/3736408).
 Applying LLMs to topic extraction and assignment requires well-designed prompts to achieve high accuracy.
+In our setup, we adopt this principle by refining prompts through multiple pilot runs on statistically significant samples to improve label consistency and clarity.
 In this RQ, we investigate two key aspects of prompt design: 1) the extent to which cleaning the input data prior to prompting improves accuracy, and 2) the effect of varying the number of few-shot examples in the prompt.
@@
 To design an effective topic extraction prompt, we experiment with varying levels of data cleaning and different numbers of few-shot examples.
-As described in section 3, we design a four-stage data cleaning pipeline.
+As described in section 3, we design a four-stage data cleaning framework.
 Stage 0 includes basic textual cleaning.
-Stage 1 excludes reviews before the launch of ChatGPT in 2022 (Akhtar, 2024).
+Stage 1 excludes reviews before the launch of ChatGPT in 2022 (Akhtar_2024).
 Stage 2 excludes short reviews with three words or less.
@@
 This sample is used for topic assignment, where manual evaluation is feasible and interpretable.
+Figure 5.
+Template of the topic extraction prompts for (E1: 0-shot, E2: 3-shot, and E3: 5-shot) used to extract the top topics from a large sample of stage $x$ and app category $y$.
+Insert category name is where we insert the name of the app category $y$.
+Insert sample of reviews is where we list our large sample of reviews separated by a new line escape.
 Topic extraction.
-To evaluate how prompt design affects topic extraction performance, we design three experiments, i.e., E1: 0 shot, E2: 3 shot and E3: 5 shot, with three prompt configurations differing in the number of few-shot examples, i.e., 0-shot, 3-shot and 5-shot.
-These values are chosen based on prior research and practical considerations in prompt-based learning  (Dos Santos et al., 2023; Assi et al., 2025a; Pham et al., 2024).
+Because prior work reported mixed effects of few-shot examples (Wang_Yang_Wei_2023; Prakash_Wang_Hoang_Hee_Lee_2023), we design three experiments, i.e., E1: 0 shot, E2: 3 shot and E3: 5 shot, with three prompt configurations differing in the number of few-shot examples, i.e., 0-shot, 3-shot and 5-shot.
+These values are chosen based on prior research and practical considerations in prompt-based learning  (Dos_Santos_Oliveira_De_Jesus_Aljedaani_Eler_2023; Maram_LLM_Cure; Pham_Hoyle_Sun_Resnik_Iyyer_2024).
 In E1: 0 shot, the prompt contains only the instructions.
-In E2: 3 shot, the prompt includes the instruction with three shots, i.e., three examples of user reviews with their topics.
+In E2: 3 shot, the prompt includes the instructions with three shots, i.e., three examples of user reviews with their topics.
 In E3: 5 shot, the prompt includes the instructions with five examples of user reviews with their topics.
@@
 Figure 5 shows the three prompt configuration templates.
-Figure 5.
-Template of the topic extraction prompts for (E1: 0-shot, E2: 3-shot, and E3: 5-shot) used to extract the top topics from a large sample of stage $x$ and app category $y$.
-Insert category name is where we insert the name of the app category $y$.
-Insert sample of reviews is where we list our large sample of reviews separated by a new line escape.
-Topic assignment.
-To evaluate the effectiveness of the extracted topic sets $T(x,y,n)$ produced by each topic extraction experiment with $n\in\{0,3,5\}$, we perform a follow-up topic assignment task using LLM.
-Specifically, for each experiment $n$, we design a dedicated assignment prompt, illustrated in Figure 6, that takes as input: 1) the set of top topics $T(x,y,n)$ extracted in that experiment, 2) five few-shot example reviews manually selected and labelled from the corresponding large sample $S_{\text{large}}(x,y)$, and 3) the target reviews from the small sample $S_{\text{small}}(x,y)$ to which topics are assigned.
-We manually select and assign topics to five few-shot examples to improve assignment accuracy as done in prior work (Assi et al., 2025a; Deshmukh et al., 2025).
+To mitigate hallucination during topic extraction, we anchor the prompt in actual user reviews, instruct the model to output only ten distinct high-level topics that describe app features, functionalities, and utility, and provide few-shot examples that illustrate the expected level of abstraction and discourage free-form or overly creative topic invention.
+To further adapt the LLM to the Gen-AI domain, the few-shot examples are drawn from real reviews of Gen-AI apps, exposing the model to characteristic terminology and contexts (e.g., discussions of AI Performance, Technical Difficulties, and Content Quality).
 Figure 6.
@@
 We insert the set of top topics $T(x,y,n)$, five few-shot examples, and the small sample $S_{\text{small}}(x,y)$.
+Topic assignment.
+To evaluate the effectiveness of the extracted topic sets $T(x,y,n)$ produced by each topic extraction experiment with $n\in\{0,3,5\}$, we perform a follow-up topic assignment task using LLM.
+Specifically, for each experiment $n$, we design a dedicated assignment prompt, illustrated in Figure 6, that takes as input: 1) the category-specific set of top topics $T(x,y,n)$ extracted for that app category in that experiment, 2) five few-shot example reviews manually selected and labeled from the corresponding large sample $S_{\text{large}}(x,y)$, and 3) the target reviews from the small sample $S_{\text{small}}(x,y)$ to which topics are assigned.
+We manually select and assign topics to five few-shot examples to improve assignment accuracy as done in prior work (Maram_LLM_Cure; Deshmukh_Raut_Bhavsar_Gurav_Patil_2025).
+To mitigate hallucination during topic assignment, we constrain the model’s output to a predefined set of deterministic labels, i.e., the extracted topics $T(x,y,n)$.
+The model must choose from these labels or select “Other” if none accurately represent the review.
+This closed-set format enables immediate detection of hallucination, as any label outside the predefined list signals a deviation from valid outputs.
+Additionally, structured CSV-formatted output further enforces adherence to the expected schema and prevents free-form or unconstrained generation.
+To ensure domain relevance, the five few-shot examples are handpicked from real reviews of Gen-AI apps, representing common contexts and linguistic patterns found in this domain (e.g., user reviews of AI Performance, Utility & Use Cases, or Content Quality).
 Accuracy evaluation.
@@
 An assignment is considered correct if the LLM assigned topic accurately reflects the content of the review.
-The Cohen’s Kappa agreement score (Cohen, 1960) is computed, yielding a score of 0.56, indicative of moderate agreement.
-The two authors discuss and resolve the initial disagreements to align on labelling criteria.
+The Cohen’s Kappa agreement score (cohen1960kappa) is computed, yielding a score of 0.56, indicative of moderate agreement.
+The two authors discuss and resolve the initial disagreements to align on labeling criteria.
 In total, 4,353 LLM topic assignments are manually evaluated for correctness for five categories, three stages, and three experiments.
@@
 To determine whether the observed accuracy differences are statistically significant, we perform statistical tests.
-Specifically, we apply the non-parametric test, Kruskal–Wallis H test (Chan and Walmsley, 1997), to evaluate the effect of few-shot prompt configurations on overall accuracy.
-Additionally, post-hoc pairwise comparisons are conducted using the Mann–Whitney U test (Mann and Whitney, 1947) with Bonferroni correction to adjust for multiple testing (Cabin and Mitchell, 2000).
-We also use one-way ANOVA (Judd et al., 2017) to test for significant differences across cleaning stages and few-shot settings.
+Specifically, we apply the non-parametric test, Kruskal–Wallis H test (Chan_Walmsley_1997), to evaluate the effect of few-shot prompt configurations on overall accuracy.
+Additionally, post-hoc pairwise comparisons are conducted using the Mann–Whitney U test (Mann_Whitney_1947) with Bonferroni correction to adjust for multiple testing (3b6386c9-eeae-33a7-8c59-c0604aa87bef).
+We also use one-way ANOVA (Judd_McClelland_Ryan_2017) to test for significant differences across cleaning stages and few-shot settings.
+Table 3.
+Accuracy across three stages and three experiments of five app categories: 1) Photography (Photo), 2) Productivity (Pro), 3) Art & Design (Art), 4) Entertainment (Ent), and 5) Video Players & Editors (Video).
+(Representative subset considered).
+Experiment Stage 1: Temporal reviews Stage 2: Short reviews Stage 3: Non-informative reviews
+Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg.
+E1: 0 shot 86% 77% 91% 69% 95% 83% 84% 85% 85% 80% 90% 85% 72% 86% 77% 86% 89% 82%
+E2: 3 shot 91% 86% 80% 66% 95% 83% 91% 80% 86% 82% 95% 87% 76% 83% 88% 92% 90% 86%
+E3: 5 shot 93% 81% 94% 90% 92% 90% 90% 84% 92% 88% 93% 89% 95% 86% 86% 95% 91% 91%
 4.2. RQ1 Findings
@@
 These results suggest that increasing the number of few-shot examples, specifically 5-shot, significantly enhances overall model accuracy.
-Table 3.
-Accuracy across three stages and three experiments of five app categories: 1) Photography (Photo), 2) Productivity (Pro), 3) Art & Design (Art), 4) Entertainment (Ent), and 5) Video Players & Editors (Video).
-Experiment Stage 1: Temporal reviews Stage 2: Short reviews Stage 3: Non-informative reviews
-Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg.
-E1: 0 shot 86% 77% 91% 69% 95% 83% 84% 85% 85% 80% 90% 85% 72% 86% 77% 86% 89% 82%
-E2: 3 shot 91% 86% 80% 66% 95% 83% 91% 80% 86% 82% 95% 87% 76% 83% 88% 92% 90% 86%
-E3: 5 shot 93% 81% 94% 90% 92% 90% 90% 84% 92% 88% 93% 89% 95% 86% 86% 95% 91% 91%
 Observation 1.2: The proportion of informative reviews varies by cleaning stage, which impacts overall accuracy.
@@
 As a result, overall accuracy may not accurately reflect the model’s effectiveness in handling more meaningful and content-rich reviews across different cleaning stages.
+Table 4.
+Accuracy of informative reviews across stages and app categories, Photography (Photo), Productivity (Pro), Art & Design (Art), Entertainment (Ent), Video Players & Editors (Video).
+(Representative subset considered).
+Experiment Stage 1: Temporal reviews Stage 2: Short reviews Stage 3: Non-informative reviews
+Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg.
+E1: 0 shot 83% 76% 90% 88% 76% 83% 84% 83% 89% 74% 82% 82% 81% 91% 81% 86% 89% 86%
+E2: 3 shot 87% 91% 87% 86% 76% 85% 93% 78% 89% 89% 94% 88% 81% 88% 89% 91% 90% 88%
+E3: 5 shot 87% 85% 94% 82% 76% 85% 80% 79% 90% 82% 89% 84% 94% 92% 90% 95% 97% 94%
 Observation 1.3: Stage 3 consistently yields the highest accuracy on informative reviews across all app categories and experiments.
@@
 The ANOVA for few-shot experiments does not reveal significant differences ($F(2,42)=2.25$, $p=0.118$).
-Table 4. Accuracy of informative reviews across stages and app categories, Photography (Photo), Productivity (Pro), Art & Design (Art), Entertainment (Ent), Video Players & Editors (Video).
-Experiment Stage 1: Temporal reviews Stage 2: Short reviews Stage 3: Non-informative reviews
-Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg. Photo Pro Art Ent Video Avg.
-E1: 0 shot 83% 76% 90% 88% 76% 83% 84% 83% 89% 74% 82% 82% 81% 91% 81% 86% 89% 86%
-E2: 3 shot 87% 91% 87% 86% 76% 85% 93% 78% 89% 89% 94% 88% 81% 88% 89% 91% 90% 88%
-E3: 5 shot 87% 85% 94% 82% 76% 85% 80% 79% 90% 82% 89% 84% 94% 92% 90% 95% 97% 94%
 Observation 1.4: The best performing setup combines filtering non-informative reviews, i.e., Stage 3 and more few shot examples, i.e., E3: 5-shot.
@@
 Based on these findings, we adopt the Stage 3 – E3: 5-shot configuration for topic classification and assignment in RQ2, as it yielded the highest overall accuracy of 91% and the highest informative accuracy of 94%.
+Observation 1.5: Under the best-performing configuration (E3: 5-shot), only 9% of reviews were misclassified.
+Among these, 2% of all reviews involved incorrect topic assignments, 3% were missed topics (valid reviews marked as “Other”), and 4% were over-assignments (non-topical reviews given a label).
+Most errors occurred in short or ambiguous reviews with limited context, where topic boundaries were less clear.
+These findings indicate that misclassifications are infrequent and stem from understandable edge cases rather than uncontrolled generation or hallucination, reinforcing the reliability of the LLM-based labeling process.
 Summary of RQ 1 Our findings indicate that increasing the number of few-shot examples in the prompt significantly improves output accuracy.
@@
 Accordingly, we adopt this configuration for topic extraction in RQ2.
-5. RQ2: What are the most prominent topics discussed in user reviews of Gen-AI apps?
+5. RQ2: What are the most prominent topics discussed in user reviews and in developer responses of Gen-AI apps?
+Having validated the accuracy of our LLM-based topic extraction and assignment framework in RQ1, we now apply it to the full dataset to examine what users discuss in their reviews of Gen-AI apps.
 Understanding the issues that users care about is essential for improving the design and evaluation of Gen-AI apps.
@@
 The results of RQ1 indicate that the optimal dataset for analysis is Stage 3 data, which underwent four stages of cleaning: basic cleaning, temporal filtering, short review exclusion, and LLM-assisted informative review filtration.
-Our dataset spans 11 app categories, totalling 676,066 user reviews as shown in Table 2.
+Our dataset spans 11 app categories, totaling 676,066 user reviews as shown in Table 2.
 As in RQ1, we sample a statistically significant set of informative reviews (large sample) per app category, using a 95% confidence level and 2% margin of error.
@@
 To do so, we manually examine all extracted topics and label them as either Gen-AI topics, which reference Gen-AI capabilities, or non-Gen-AI topics, which pertain to general app features outside the scope of Gen-AI.
+All reviews originate from Gen-AI apps; the distinction here is between review types, not between Gen-AI and non-Gen-AI apps.
 We further group semantically related topics into high-level topic categories based on shared themes or functionality, facilitating higher-level interpretation of user concerns.
-In line with prior studies (Assi et al., 2025a; Pagano and Maalej, 2013; Chen et al., 2014), we also classify reviews with 1 or 2-star ratings as negative, 3-star rating as neutral, and 4 or 5-star rating as positive.
-Analysis. We perform quantitative analysis on both app categories and topic categories. We compute the following metrics:
+In line with prior studies (Maram_LLM_Cure; Pagano_Maalej_2013; Chen_Lin_Hoi_Xiao_Zhang_2014), we also classify reviews with 1 or 2-star ratings as negative, 3-star rating as neutral, and 4 or 5-star rating as positive.
+Analysis.
+We perform large-scale quantitative analysis on all automatically labeled reviews across both app categories and topic categories, followed by qualitative analysis on a representative sample.
+Specifically, for the quantitative component, we compute the following metrics:
   1. (1)
-    Average Gen-AI Rating (AGR): the average star rating of reviews labelled with Gen-AI topics.
+    Average Gen-AI Rating (AGR): the average star rating of reviews labeled with Gen-AI topics.
   2. (2)
-    Average Non-Gen-AI Rating (ANR): the average rating of reviews labelled with non-Gen-AI topics.
+    Average Non-Gen-AI Rating (ANR): the average rating of reviews labeled with non-Gen-AI topics.
   3. (3)
-    Gen-AI Review Count (GRC): the number of reviews assigned to Gen-AI
+    Gen-AI Review Count (GRC): the number of reviews assigned to Gen-AI topics.
   4. (4)
-    Non-Gen-AI Review Count (NRC): the number of reviews assigned to non-Gen-AI topics
+    Non-Gen-AI Review Count (NRC): the number of reviews assigned to non-Gen-AI topics.
   5. (5)
-    Gen-AI Review Percentage (GRP): the proportion of each type relative to the total number of reviews in each app category
+    Gen-AI Review Percentage (GRP): the proportion of Gen-AI reviews relative to the total number of reviews.
   6. (6)
-    Non-Gen-AI Review Percentage (NRP): the proportion of each type relative to the total number of reviews in each topic category
-We conduct the Mann–Whitney U test to examine whether reviews labelled as Gen-AI-related receive significantly higher ratings than those that are not.
-To gain deeper insights into user concerns and feedback within each topic category, we randomly sample reviews using a 95% confidence level and a 10% margin of error, resulting in a dataset of 901 reviews.
+    Non-Gen-AI Review Percentage (NRP): the proportion of non-Gen-AI reviews relative to the total number of reviews.
+We conduct the Mann–Whitney U test to examine whether reviews labeled as Gen-AI-related receive significantly higher ratings than those that are not.
+For developer-response analysis, we compute reply ratios, response delays, and response-type distributions for both Gen-AI and non–Gen-AI reviews, and we compare per-app reply behavior using a Wilcoxon signed-rank test (wilcoxon1992).
+We complement our quantitative analyses with qualitative examination by manually coding statistically significant samples (95% confidence level, 10% margin of error) of both user reviews and developer responses.
+For user reviews we draw samples from every Gen-AI topic category resulting in a total of 901 reviews.
 We manually code and analyze all sampled reviews to provide richer, contextualized insights that may surface actionable implications for both researchers and developers.
+For developer responses, we perform two qualitative analyses.
+First, we code a representative sample of 96 replies drawn across all app categories to characterize the overall distribution of developer responses.
+Second, we manually code an additional 281 replies from three focal categories (Games, Education, and Entertainment) to capture deeper, category-specific response practices.
+Together, these qualitative analyses provide contextual grounding for the quantitative findings and offer a more complete understanding of both user concerns and developer engagement within Gen-AI apps.
 5.2. RQ2 Findings
-Observation 2.1: Gen-AI reviews consistently receive higher user ratings (AGR) than non-Gen-AI reviews (ANR) across all app categories.
-As shown in Table 5, the average rating of reviews discussing Gen-AI functionality exceeds the average rating of non-Gen-AI reviews in every app category, indicating higher user satisfaction with the Gen-AI features.
+Observation 2.1: Gen-AI reviews consistently receive higher user ratings (AGR) than non-Gen-AI reviews (ANR) across all app categories of Gen-AI apps.
+As shown in Table 5, the average rating of reviews discussing Gen-AI functionality (i.e., AI Performance and Utility & Use Cases) exceeds the average rating of non-Gen-AI reviews (i.e., Technical Difficulties and Monetization Methods & Structure) in every app category of Gen-AI apps.
+This indicates that, when users explicitly comment on Gen-AI features, they tend to rate their experiences more positively than when they comment on other aspects of the same apps.
 To assess the statistical significance of this difference, we apply the Mann–Whitney U test.
-The results reveal a significant difference in the rating distributions between the two groups ($U=65{,}113{,}513{,}850$, $p<0.0001$).
+The results reveal a significant difference in the rating distributions between the two groups (U=65,113,513,850U=65{,}113{,}513{,}850, $p<0.0001$).
 Gen-AI reviews tend to receive higher ratings, with a median rating of 5.0 and a mean of 4.10, compared to a median of 4.0 and a mean of 3.32 for non-Gen-AI reviews.
 These findings suggest that Gen-AI functionality is a major contributor to positive user experiences across a wide range of app categories.
+This pattern is especially pronounced in categories such as Education and Books & Reference, which show some of the largest gaps between AGR and ANR.
+In Books & Reference, Gen-AI topics such as AI Performance and Human AI Interaction & Emotional Connection receive high ratings, while non-Gen-AI topics like Technical Difficulties and Customer Support are rated much lower due to crashes, generation failures, and unresolved support issues.
+Similarly, in Education, users highly rate Gen-AI topics such as Content Quality and Utility & Use Cases, but express dissatisfaction with Monetization Methods & Structure and Technical Difficulties.
+Together, these results indicate that strong Gen-AI features can substantially boost user satisfaction, but poor reliability, monetization, or support can still pull overall ratings down.
+Developers should therefore treat Gen-AI capabilities as amplifiers of perceived value, not as substitutes for robust app quality, stability, and fair pricing.
 Table 5.
-Summary of app categories, including average Gen-AI rating (AGR), average non-Gen-AI rating (ANR), Gen-AI Reviews Count (GRC), Gen-AI reviews percentage (GRP), and number of distinct Gen-AI topics
+Summary of app categories, including average Gen-AI rating (AGR), average non-Gen-AI rating (ANR), Gen-AI Reviews Count (GRC), Gen-AI reviews percentage (GRP), and number of distinct Gen-AI topics (Full dataset considered).
 App Category AGR ANR GRC GRP # Gen-AI Topics
@@
 Observation 2.2: Gen-AI functionality is a prominent focus in user feedback, comprising an average of 38% of reviews across app categories and reaching as high as 88% for the Health & Fitness category.
-As reported in Table 5, the prevalence of Gen-AI–focused reviews varies by app category, reflecting the differing roles Gen-AI plays across app domains.
-For example, categories such as Health & Fitness, Education, and Productivity contain the highest proportions of Gen-AI reviews, with 88%, 75%, and 50%, respectively.
-Conversely, categories such as Photography and Video Players & Editors exhibit lower percentages of Gen-AI reviews.
-In these categories, 58% of the reviews focus on non-Gen-AI topics, such as overall user experience, technical issues, and subscription or pricing models, suggesting that these factors outweighed interest in Gen-AI features.
-Figure 7 illustrates the distribution of Gen-AI review percentages across individual apps within various categories, highlighting key differences in engagement.
-Several categories, such as Education and Productivity, exhibit a moderate spread, indicating consistent Gen-AI engagement across apps.
-Other categories, such as Art & Design and Entertainment present a wider spread, suggesting more variability in Gen-AI adoption among apps within these groups.
-Figure 7. Box plot showing the distribution of the percentage of reviews classified as Gen-AI at the app level, grouped by app category
-$\Rightarrow$ Implication 1.
-Gen-AI reviews consistently yield higher user ratings across app categories.
-Developers must recognize Gen-AI as a crucial tool for boosting overall user satisfaction.
-They must strategically incorporate Gen-AI features to remain competitive.
-Observation 2.3: The Gen-AI topics with the highest numbers of reviews are AI Performance, Utility & Use Cases, and Content Quality.
+As shown in Table 5, the prevalence of Gen-AI–focused reviews varies by app category, reflecting the differing roles Gen-AI plays across app domains.
+Categories such as Health & Fitness, Education, and Productivity contain the highest proportions of Gen-AI reviews, with 88%, 75%, and 50%, respectively, indicating that users in these domains directly engage with AI features to achieve personal or task-oriented goals such as self-improvement, learning, or workflow assistance.
+Figure 7. Box plot showing the distribution of the percentage of reviews classified as Gen-AI at the app level, grouped by app category (Full dataset considered).
+In contrast, categories such as Photography and Video Players & Editors exhibit lower percentages of Gen-AI reviews, where 58% of feedback still focuses on non-Gen-AI topics such as usability, technical issues, and subscription or pricing models.
+This suggests that in creative or media-focused apps, users remain more concerned with traditional performance and value-for-money factors than with AI-driven functionality.
+Figure 7 illustrates the distribution of Gen-AI review percentages across individual apps within various categories, highlighting how consistently users discuss Gen-AI features within each domain.
+Categories such as Entertainment and Health & Fitness exhibit the widest spreads, suggesting substantial variation among apps; some are highly AI-centric, while others receive minimal user attention to Gen-AI features.
+In contrast, categories like Video Players & Editors and Photography show narrow spreads concentrated at low percentages, indicating that across these domains, Gen-AI features consistently play an assistive or complementary role rather than being the central focus of user experience.
+Together, these patterns suggest that while certain app categories demonstrate uneven user emphasis on Gen-AI, others achieve relatively uniform user engagement of Gen-AI features.
+Observation 2.3: Users discuss a wide range of Gen-AI topics, revealing where developers excel and where improvement is needed.
 Table 6 and Table 7 present the top Gen-AI and non-Gen-AI topics, respectively.
-In terms of user satisfaction, the Gen-AI topics with the highest average ratings are Human AI Interaction & Emotional Connection, Utility & Use Cases, and Customer Support & Community.
+Developers should pay particular attention to topics that appear most frequently across apps because they represent the aspects users discuss and evaluate most.
+Excelling in these areas is critical for maintaining competitiveness and user satisfaction.
+The Gen-AI topics with the highest numbers of reviews are AI Performance, Utility & Use Cases, and Content Quality.
 The topics most commonly mentioned across different apps are AI Performance, Content Quality, and Content Policy & Censorship.
-We further examine each of these categories in the subsections below.
+Similarly, topics with high average ratings (e.g., Human AI Interaction & Emotional Connection and Utility & use cases) serve as benchmarks for best user experiences that other apps can emulate.
+Conversely, lower-rated or frequently criticized topics (e.g., Content Policy & Censorship and Content Quality) highlight recurring weaknesses across Gen-AI apps.
+Developers should treat these as priority areas for refinement to stay ahead of the competition and sustain user trust.
+We further examine each of the Gen-AI categories in the subsections below.
 Table 6.
-Summary of Gen-AI topic categories, including the Gen-AI Reviews Count (GRC), the average Gen-AI rating (AGR), the number of apps, and the number of app categories in which each topic category appears
+Summary of Gen-AI topic categories, including the Gen-AI Reviews Count (GRC), the average Gen-AI rating (AGR), the number of apps, and the number of app categories in which each topic category appears (Full dataset considered).
 Gen-AI Topic Category GRC AGR # Apps # App Categories
@@
 Table 7.
-Summary of non-Gen-AI topic categories, including the non-Gen-AI Review Count (NRC), the average non-Gen-AI rating (ANR), the number of apps, and the number of app categories in which each topic category appears
+Summary of non-Gen-AI topic categories, including the non-Gen-AI Review Count (NRC), the average non-Gen-AI rating (ANR), the number of apps, and the number of app categories in which each topic category appears (Full dataset considered).
 Non-Gen-AI Topic Category NRC ANR # Apps # App Categories
@@
 Although memory-related concerns (e.g., forgetting prior context or repeating content) are occasionally mentioned as shown in review 5 in Figure 8, they tend to be rated more leniently, indicating that users may be more forgiving of these limitations.
-$\Rightarrow$ Implication 2.
-While users are generally satisfied with the overall performance of Gen-AI apps, a portion of their concerns relate to limitations in human-AI communication, particularly when the AI struggles to interpret nuanced or complex prompts.
-To better meet user expectations and improve usability, developers should prioritize enhancing the Gen-AI’s ability to understand and respond appropriately to more complex interactions.
 2) Content Quality. Content Quality refers to users’ assessment of AI-generated content in terms of usefulness, coherence, creativity, engagement, structure and presentation.
@@
 However, 33% of reviews reveal concerns in three domains.
-First, users report that image-generation features frequently fail to produce accurate or faithful visual outputs, particularly for complex or creative prompts (e.g., failed attempts to depict two people and an eagle wearing denim shorts in reviews 3 and 4 in Figure 9, respectively.) This finding aligns with broader evidence that AI image generators often collapse under high-complexity tasks (Petsiuk et al., 2022).
+First, users report that image-generation features frequently fail to produce accurate or faithful visual outputs, particularly for complex or creative prompts (e.g., failed attempts to depict two people and an eagle wearing denim shorts in reviews 3 and 4 in Figure 9, respectively.) This finding aligns with broader evidence that AI image generators often collapse under high-complexity tasks (Petsiuk_Siemenn_Surbehera_Chin_Tyser_Hunter_Raghavan_Hicke_Plummer_Kerret_et_al_2022).
 Second, 8% of reviews highlight concerns regarding insufficient diversity and inclusion.
@@
 Third, a minority of users (3% of reviews) highlight formatting and user interface issues in text outputs, such as grammatical errors, excessive length, inconsistent structure, indicating a need for improved polish as illustrated in review 6 in Figure 9.
-$\Rightarrow$ Implication 3.
+$\Rightarrow$ Implication 1.
+Gen-AI reviews consistently yield higher user ratings across app categories.
+Developers must recognize Gen-AI as a crucial driver of overall user satisfaction and strategically incorporate these features to remain competitive.
+However, they should treat Gen-AI capabilities as amplifiers of perceived value, not as substitutes for robust app quality, stability, and fair pricing.
+Even the most advanced AI functionality cannot compensate for poor technical reliability, weak customer support, or frustrating monetization structures.
+While users are generally satisfied with the overall performance of Gen-AI apps, a portion of their concerns relate to limitations in human-AI communication, particularly when the AI struggles to interpret nuanced or complex prompts.
+To better meet user expectations and improve usability, developers should prioritize enhancing the Gen-AI’s ability to understand and respond appropriately to more complex interactions.
 Although users generally regard Gen-AI outputs as high quality, enhancing image accuracy, diversity, inclusion, formatting, and user interface presentation remains essential for delivering truly reliable and user-centered content.
@@
 Some users report that their creations were reused or uploaded by the platform without consent, as shown in review 4 in Figure 10.
-$\Rightarrow$ Implication 4.
+$\Rightarrow$ Implication 2.
 Developers need to strike the right balance in filtering content to avoid blocking genuinely harmful material while preventing harmless content.
@@
 This indicates that even well‑received Gen-AI additions (such as voice interaction) can alienate users if they disrupt existing valued controls.
-$\Rightarrow$ Implication 5.
-Developers should design Gen-AI features, such as voice mode, in a way that complements existing controls rather than replacing them.
-For instance, it is crucial to ensure new AI-driven tools coexist with familiar functionalities, so that developers can both innovate and maintain intuitive user experiences.
 Figure 11. Examples of reviews discussing Features & Functionality
@@
 Beyond education, 11% of reviews describe how the Gen-AI tools facilitate routine work tasks, such as inspiring new ideas, structuring project workflows, and even acting as fitness or guitar instructors, as illustrated in review 3 in Figure 12.
-$\Rightarrow$ Implication 6.
-Despite the launch of tools like ChatGPT Edu (OpenAI, 2024a), which is tailored for academic contexts, Gen-AI platforms must go further by prioritizing accessibility-first design, for example, enabling assistive note-taking, voice transcription, and dyslexia support features.
-Developers should also integrate customizable, role-based modes or templates, such as “legal advisor,” “fitness coach,” or “creative brainstorming partner,” to make these tools more adaptable and effective across diverse professional and personal applications.
 6) Personalization & Customization.
@@
 For example, review 3 in Figure 13 describes the absence of customized voice features as a “deal breaker.”
-$\Rightarrow$ Implication 7.
+$\Rightarrow$ Implication 3.
+Developers should design Gen-AI features, such as voice mode, in a way that complements existing controls rather than replacing them.
+For instance, it is crucial to ensure new AI-driven tools coexist with familiar functionalities, so that developers can both innovate and maintain intuitive user experiences.
 Developers should implement customizable output options, enabling users to format, export, and save AI-generated content (e.g., lists, summaries) in a manner that suits their workflows.
@@
 Additionally, offering persona customization, such as selectable character roles or voice styles, will enhance user engagement and make interactions feel more personalized.
+Despite the launch of tools like ChatGPT Edu (openaiChatGPTedu), which is tailored for academic contexts, Gen-AI platforms must go further by prioritizing accessibility-first design, for example, enabling assistive note-taking, voice transcription, and dyslexia support features.
+Developers should also integrate customizable, role-based modes or templates, such as “legal advisor,” “fitness coach,” or “creative brainstorming partner,” to make these tools more adaptable and effective across diverse professional and personal applications.
 7) Comparison to Other Apps. Comparison to Other Apps refers to comparing the AI app to other tools users have previously tried.
@@
 Their apps must offer distinct value that goes beyond the general-purpose functionality of ChatGPT.
-$\Rightarrow$ Implication 8.
-ChatGPT is a baseline comparator to users, in order to succeed developers must offer value beyond what ChatGPT offers.
-Developers should also consider creating a toggle button to allow users to opt in or out of AI-enabled features, such as AI search.
-The regular features may be sufficient for users (e.g., regular search can satisfy users’ searching needs, and AI search is not always necessary).
 8) Customer Support & Community.
@@
 In our sample, we find that 51% of users express enjoyment with specific features (e.g., review 1 in Figure 15), 19% suggest new features to enhance their experience (e.g., review 2 in Figure 15), 9% request to fix app shortcomings (e.g., review 3 in Figure 15), and 21% discuss generic enjoyment of the app without relating it to a specific feature.
-$\Rightarrow$ Implication 9.
+$\Rightarrow$ Implication 4.
+ChatGPT is a baseline comparator to users, in order to succeed developers must offer value beyond what ChatGPT offers.
+Developers should also consider creating a toggle button to allow users to opt in or out of AI-enabled features, such as AI search.
+The regular features may be sufficient for users (e.g., regular search can satisfy users’ searching needs, and AI search is not always necessary).
 User reviews highlight key features that developers should preserve when updating their apps.
-Prior work provided approaches that facilitate competitor analysis (Assi et al., 2025a; Hassan et al., 2022; Assi et al., 2021b).
+Prior work provided approaches that facilitate competitor analysis (Maram_LLM_Cure; PeerApps_SANER; Assi_Hassan_Tian_Zou_2021).
 For competing developers, analyzing user feedback from similar apps can also uncover desirable features worth adopting to remain competitive in the market.
@@
 Finally, users point to the therapeutic potential of conversational AI in providing mental and emotional support, such as managing mental health challenges (e.g., review 3 in Figure 16) and developing strategies to achieve personal goals.
-$\Rightarrow$ Implication 10.
-The perception of conversational Gen-AI apps as emotionally supportive and non-judgmental suggests their growing role as trusted companions, particularly in contexts involving emotional vulnerability.
-Developers should account for these evolving user expectations by ensuring that Gen-AI interfaces maintain a tone of empathy and neutrality, while also incorporating safeguards to avoid over-reliance or the illusion of professional psychological care.
-Furthermore, this trend highlights the need for interdisciplinary collaboration with mental health professionals to responsibly shape the affective design of Gen-AI systems used for emotionally sensitive interactions.
 10) Accessibility & Inclusivity.
@@
 Users discuss different accessibility and inclusivity topics including: 1) financial barriers (raised in 31% of the reviews, such as review 1 in Figure 17), 2) greater language support (raised in 22% of the reviews, such as review 2 in Figure 17), 3) disability-related needs (raised in 8% of the reviews, such as reviews 3 and 4 in Figure 17), 4) age-related concerns (raised in 10% of the reviews, such as Reviews 5 and 6 in Figure 17), and 5) general accessibility concerns (raised in 31% of the reviews).
-$\Rightarrow$ Implication 11.
-These findings highlight the importance of designing Gen-AI apps with diverse accessibility needs in mind.
+$\Rightarrow$ Implication 5.
+The perception of conversational Gen-AI apps as emotionally supportive and non-judgmental suggests their growing role as trusted companions, particularly in contexts involving emotional vulnerability.
+Developers should account for these evolving user expectations by ensuring that Gen-AI interfaces maintain a tone of empathy and neutrality, while also incorporating safeguards to avoid over-reliance or the illusion of professional psychological care.
+Furthermore, this trend highlights the need for interdisciplinary collaboration with mental health professionals to responsibly shape the affective design of Gen-AI systems used for emotionally sensitive interactions.
+The findings further highlight the importance of designing Gen-AI apps with diverse accessibility needs in mind.
 Developers should prioritize inclusive design by integrating features such as read-aloud support, customizable text presentation, and adaptive interfaces tailored to neurodiverse users (e.g., Attention Deficit Hyperactivity Disorder, Dyslexia and Dyspraxia).
+Observation 2.4: Developer Responses to Gen-AI Reviews Are Primarily Courteous or Diagnostic, With Few Immediate Resolutions.
+Table 8 shows that for Gen-AI apps, the most common developer response type is Acknowledgment and Appreciation (31%), which consists largely of brief thank-you messages.
+These replies most often accompany positive reviews and serve as courteous engagement rather than substantive technical follow-up.
+The second most frequent category, Request for Additional Information (30%), reflects the practical limits of troubleshooting Gen-AI issues through review threads.
+Even when space is not a constraint, users rarely provide the detailed context needed to diagnose problems that relate to model behavior, device configuration, or account-specific issues.
+Developers therefore redirect users to email or in-app support channels where logs, device information, or credentials can be shared securely.
+By contrast, Issue Resolved and Update Advised appears rarely (2%), whereas Acknowledged and Under Investigation is more common (10%).
+This pattern aligns with the nature of Gen-AI systems: many issues require model adjustments, infrastructure changes, or updates to safety and moderation mechanisms, none of which can be addressed immediately or within the constraints of a review thread.
+As a result, developers are more likely to acknowledge problems and communicate that work is underway than to confirm the issue has been resolved.
+Table 8. Summary of developer response types, including descriptions, number of reviews, and percentage distribution (Representative subset considered).
+Topic Description # Reviews %
+Acknowledgment and Appreciation Generic responses that thank users for their feedback, often including a brief note about considering future updates or improvements. 30 31%
+Request for Additional Information Responses indicating that the reported issue lacks sufficient detail and requesting that the user contact the developer directly for further clarification or resolution.
+29 30%
+Clarification or Policy Rebuttal Responses that clarify or refute user concerns, often explaining app policies or correcting misunderstandings.
+Most common for reviews concerning free versus premium features.
+13 14%
+Guidance and Troubleshooting Tips Responses that address reported issues by offering troubleshooting advice, resources, or specific steps to help users resolve the problem. 12 13%
+Acknowledged and Under Investigation Responses confirming that the development team is aware of the issue and is currently investigating or working on a solution. 10 10%
+Issue Resolved and Update Advised Responses informing the user that the reported issue has been resolved and recommending that they update or reinstall the app. 2 2%
+Total 96 100%
+Hassan et al.
+(developer_response_safwat) analyzed developer replies across a broad set of mobile apps and identified recurring response drivers such as “thank the user”, “ask for more details”, “provide guidance”, “ask for endorsement”, “advertise other products”, wait for updates, and “justify advertisements”.
+Many of these practices align closely with our taxonomy.
+For example, their “thank the user” category maps to our “Acknowledgment and Appreciation,” their “ask for more details” category maps to our “Request for Additional Information,” and their “justify the advertisements” category maps onto our broader Clarification or Policy Rebuttal, which includes explanations about subscription models, premium features, and content restrictions.
+Notably, their work does not treat “issue resolved” as a dominant response driver, and we similarly observe that confirmed resolutions are relatively rare in our Gen-AI dataset (2%).
+We report the category explicitly for completeness of Gen-AI developer behaviors, but its low prevalence mirrors prior findings rather than contrasting with them.
+Overall, the strong correspondence between the two taxonomies suggests that general response practices extend to Gen-AI apps, while our domain-specific categories capture the additional emphasis developers place on clarifying AI-related behavior and managing expectations around the evolving nature of Gen-AI features.
+Observation 2.5: App categories show highly distinctive developer response patterns.
+Table 9 reveals variation in how developers across different app categories respond to user feedback.
+Some categories exhibit exceptionally high engagement, while others show very limited interaction.
+For example, Games stands out with both the highest reply ratio (56%) and the longest average response delay (44.7 days).
+In contrast, Education demonstrates the shortest response times, with developers replying in less than a day on average (0.7 days).
+Meanwhile, Entertainment apps show the lowest reply ratio among categories with meaningful activity (2%).
+These contrasting patterns suggest that developer responsiveness is shaped not only by category size but also by the operational demands, user expectations, and support norms unique to each domain.
+Table 9. Developer response activity across app categories, including total reviews, total developer replies, overall reply ratio, and average response delay (Full dataset considered).
+Category Total Reviews Total Replies Reply Ratio Avg Delay (days)
+Productivity 169,119 8,407 5% 8.4
+Photography 126,277 31,070 25% 4.5
+Art & Design 125,362 39,593 32% 5.5
+Entertainment 118,669 2,783 2% 1.7
+Education 41,905 16,034 38% 0.7
+Video Players & Editors 36,507 11,123 30% 2.2
+Tools 35,046 905 3% 32.45
+Music & Audio 9,150 2,127 23% 3.4
+Health & Fitness 7,849 338 4% 7.7
+Games 5,070 2,848 56% 44.7
+Books & Reference 1,112 1 0% 1.0
+All Categories 676,066 115,164 20% 10.2
+Our manual analysis of developer responses in the Games category helps explain why it exhibits both the highest reply ratio and the longest response delay.
+Responses in which developers report that an issue has been fixed account for 30% of all replies in this category, far exceeding the 2% observed in the overall dataset.
+This emphasis on confirmed fixes aligns with the long response delay shown in Table 9.
+This likely reflects that developers often wait until a patch is deployed before responding, resulting in slower but more substantive communication.
+Responses indicating that an issue is acknowledged and under investigation represent 13% of replies, slightly higher than the overall 10%.
+Together, these categories show that Games developers place strong emphasis on technical issue resolution and ongoing investigation.
+Generic thank-you messages appear in only 12% of replies, considerably lower than the overall 31%.
+The remaining responses highlight active engagement with user issues: 28% request additional information, 10% provide troubleshooting guidance, and 8% offer clarifications or policy rebuttals.
+Overall, developer communication in Games is markedly problem-oriented and resolution-driven, which leads to both high engagement and extended response timelines.
+A closer look at Education apps helps explain why this category exhibits the fastest average developer response time.
+In our manual sample, the majority of Education responses fall under Acknowledgment and Appreciation (73%), which is markedly higher than both the overall rate of 31% and the 12% observed in Games.
+These replies are typically short, template-based thank-you messages that require minimal effort to generate, allowing developers to respond quickly and consistently.
+Education apps also show a comparatively high rate of Request for Additional Information (16%).
+This proportion is notably larger than the overall sample and reflects a common pattern in Gen-AI Education reviews.
+Users frequently report that the app produced incorrect answers, yet they often omit the original prompt, the model’s response, or screenshots of the interaction.
+Without these details, developers cannot diagnose or reproduce the issue within the review thread.
+As a result, they quickly reply by directing users to contact support channels where complete context can be shared securely.
+This response type is therefore both common and fast, since it does not require problem investigation within the review platform itself.
+By contrast, response types requiring deeper engagement occur far less frequently in Education, including Guidance and Troubleshooting Tips (2%), Acknowledged and Under Investigation (6%), and Issue Resolved and Update Advised (1%).
+Together, these patterns show that Education apps tend to rely primarily on quick acknowledgments or requests for additional information, which contributes to the category’s exceptionally short response delay.
+Entertainment apps show the lowest reply ratio in our dataset, and the distribution of response types helps explain this pattern.
+Acknowledgment and Appreciation responses account for only 14%, far below the overall rate of 31%.
+This indicates that developers in this category do not typically reply to routine or positive reviews and instead reserve responses for situations where user action or developer follow-up is required.
+Correspondingly, a large share of replies fall under Request for Additional Information (38%) and Guidance and Troubleshooting Tips (22%), suggesting that developers primarily intervene when users report crashes, playback issues, or account problems that require diagnostic detail.
+The relatively high proportion of Acknowledged and Under Investigation (19%) further shows that many of these reports involve technical issues that cannot be resolved immediately, which may discourage broader engagement with lower-priority feedback.
+Together, these patterns indicate that Entertainment developers focus their limited response efforts on actionable problem reports rather than general user engagement, resulting in a lower overall reply ratio.
+Observation 2.6: Developer responsiveness varies substantially across Gen-AI topic categories.
+Table 10 summarizes developer reply behavior for each Gen-AI topic category.
+The topic with the highest reply ratio is Accessibility & Inclusivity (47%), despite having the smallest number of reviews.
+This indicates that developers treat accessibility-related feedback as a high-impact area that warrants immediate and consistent attention.
+Such reviews often report barriers to use for specific user populations, making them critical from both ethical and product-quality perspectives.
+The short average delay (0.2 days) further reflects the urgency developers assign to this category.
+Beyond accessibility, the next highest reply ratios are found in Utility & Use Cases (18%) and Content Quality (14%), suggesting that developers are highly responsive to feedback about how users apply AI features and whether the generated outputs meet expectations.
+These topics directly affect user satisfaction and daily usability, which likely motivates more frequent developer engagement.
+In contrast, the categories with the lowest reply ratios are Customer Support & Community (1%), Human–AI Interaction & Emotional Connection (2%), and Content Policy & Censorship (5%).
+These topics often involve complaints about communication quality, emotional tone, safety constraints, or moderation decisions.
+Such issues are typically systematic, tied to model behavior or policy rules that cannot be quickly adjusted, which may explain the lower response frequency and the longer average delays seen in these categories.
+This contrast between highly actionable categories and structurally constrained categories reinforces the idea that responsiveness is shaped by whether an issue is immediately fixable or requires deeper system changes.
+Overall, these patterns show that developers prioritize issues that are actionable in the short term (e.g., accessibility barriers, inaccurate outputs, task-level failures) while responding less frequently to concerns that require deeper model changes or policy revisions.
+This reflects a practical triaging strategy driven by the nature of Gen-AI systems and the feasibility of immediate fixes.
+Table 10. Developer response patterns across Gen-AI topic categories, including total reviews, total developer replies, reply ratio, and average response delay (Full dataset considered).
+Topic Category Total Reviews Total Replies Reply Ratio Avg Delay (days)
+AI Performance 61,071 7,152 12% 4.2
+Utility & Use Cases 56,505 9,895 18% 0.6
+Content Quality 44,594 6,275 14% 4.3
+Content Policy & Censorship 27,521 1,397 5% 7.9
+Human–AI Interaction & Emotional Connection 25,142 542 2% 7.5
+Features & Functionality 16,275 1,892 12% 5.5
+Customer Support & Community 3,361 28 1% 10.3
+Comparison to Other Apps 2,315 287 12% 6.0
+Personalization & Customization 1,283 106 8% 2.1
+Accessibility & Inclusivity 96 45 47% 0.2
+All Topics 238,168 27,619 12% 4.9
+Observation 2.7: Developers respond less frequently to Gen-AI reviews than to non-Gen-AI reviews as illustrated in Figure 18.
+We compared per-app developer response rates for Gen-AI and non-Gen-AI reviews using a paired non-parametric Wilcoxon signed-rank test ($n=164$ apps).
+Developers replied to a median of 9.6% of Gen-AI reviews versus 17.7% of non-Gen-AI reviews.
+Although the median per-app difference is modest (0.7-0.7 percentage points), the effect is statistically significant ($W=2911$, $p<0.001$), indicating a consistent pattern across apps.
+This gap aligns with the qualitative patterns observed earlier.
+Gen-AI reviews often describe failures rooted in model behavior, data quality, or moderation mechanisms, which are difficult to diagnose from short review text.
+By contrast, non-Gen-AI reviews commonly involve UI issues, feature requests, or account problems that are easier to acknowledge, triage, or resolve within the constraints of the review interface.
+As a result, developers appear slightly more willing to engage with non-AI feedback where their replies can directly influence user experience.
+In addition to analyzing user perspectives, we also examine how developers respond to Gen-AI reviews, capturing patterns of engagement, resolution, and support practices across app categories.
+Observation 2.8: Developers respond to Gen-AI and non-Gen-AI reviews at similar speeds.
+Figure 18 compares developer response delays for Gen-AI and non-Gen-AI reviews across all apps.
+While Gen-AI reviews show a slightly lower median delay (a difference of 0.15-0.15 days), this difference is not statistically significant.
+In practical terms, once developers choose to reply, they tend to respond at comparable speeds regardless of whether the review discusses Gen-AI functionality or non-AI aspects of the app.
+This suggests that the lower engagement with Gen-AI reviews observed in Observation 2.7 reflects differences in response likelihood rather than response promptness.
+Figure 18.
+Comparison of developer responsiveness across apps.
+Left: Relative likelihood of developer response to Gen-AI reviews compared to non–Gen-AI reviews, expressed as the ratio of per-app response probabilities (values below 1 indicate lower responsiveness to Gen-AI reviews).
+Right: Distribution of mean developer response delays per app for Gen-AI and non–Gen-AI reviews; the two delay distributions do not differ significantly.
+(Full dataset considered.)
+$\Rightarrow$ Implication 6.
+Taken together, our developer response findings show that review threads in Gen-AI apps primarily serve as an initial point of contact rather than a channel for resolving AI-specific failures.
+Developers face structural barriers when responding to Gen-AI feedback, since many issues require information that cannot be shared in public reviews or involve model- or policy-level changes that cannot be deployed quickly.
+As a result, developers tend to reserve detailed engagement for issues that are actionable in the short term and redirect users elsewhere when deeper diagnostic context is required.
+These patterns highlight a broader platform-level limitation: while app stores surface user feedback effectively, they do not provide the depth of interaction required to investigate or clarify Gen-AI behaviours.
+There is therefore an opportunity for platform owners to introduce more structured, built-in support mechanisms that complement review sections and enable richer, privacy-preserving communication channels.
+Such improvements would better support both developers and users as Gen-AI apps become more complex and more central to mobile ecosystems.
 Summary of RQ 2 From our analysis of user reviews across diverse Gen-AI app categories, we find that users are increasingly experimenting with prompts, offering feedback, and forming emotional connections with non-human agents.
@@
 For developers, this means moving beyond baseline functionality to support user control (e.g., ability to disable Gen-AI features) and differentiated value compared with AI tools (e.g., ChatGPT).
+Our developer-response analysis reinforces this point: developers most often engage with actionable, user-reported issues and struggle to address Gen-AI failures directly within the review platform.
+This suggests that better support channels will be essential for meeting rising user expectations.
 6. RQ3: What temporal trends emerge in user feedback on Gen-AI topics?
@@
     Features extraction.
-Feature selection in clustering involves identifying the most relevant features to uncover patterns in unlabeled data (Dash and Koot, 2009).
-Feature selection enhances interpretability, aligns with research goals  (Kreienkamp et al., 2025), and enables grouping of topics with similar temporal trends (e.g., increasing, fluctuating, or stable).
+Feature selection in clustering involves identifying the most relevant features to uncover patterns in unlabeled data (Dash_Koot_2009).
+Feature selection enhances interpretability, aligns with research goals  (Kreienkamp_Agostini_Monden_Epstude_DeJonge_Bringmann_2025), and enables grouping of topics with similar temporal trends (e.g., increasing, fluctuating, or stable).
 For each topic category $(T)$, we compute the slope $m$ and the standard deviation $\sigma$ for both $Avg_{(T,P)}$ and $Pr_{(T,P)}$.
@@
     Clustering algorithm.
-K-Means clustering is a partitioning method that groups unlabeled data into ($k$) clusters by minimizing within-cluster variance (Warren Liao, 2005).
+K-Means clustering is a partitioning method that groups unlabeled data into ($k$) clusters by minimizing within-cluster variance (Warren_Liao_2005).
 We apply K-Means clustering to group topics based on their slope $m$ and standard deviation $\sigma$ values.
-We select K-Means because it is efficient, flexible, and widely used for clustering in data mining tasks (Ikotun et al., 2023; Assi et al., 2025b).
-K-Means is suitable for low-dimensional feature spaces (Ikotun et al., 2023) and hence enables us to identify generalizable trend patterns across different topics and apps.
+We select K-Means because it is efficient, flexible, and widely used for clustering in data mining tasks (Ikotun_Ezugwu_Abualigah_Abuhaija_Heming_2023; 10.1145/3721125).
+K-Means is suitable for low-dimensional feature spaces (Ikotun_Ezugwu_Abualigah_Abuhaija_Heming_2023) and hence enables us to identify generalizable trend patterns across different topics and apps.
   3. 3.
     Optimal number of clusters.
-To determine the optimal number of clusters ($k$), we compute Silhouette Scores, one of the most widely adopted methods for evaluating clustering quality (Ashari et al., 2023), across multiple values of $k$.
+To determine the optimal number of clusters ($k$), we compute Silhouette Scores, one of the most widely adopted methods for evaluating clustering quality (Ashari_DwiNugroho_Baraku_NovriYanda_Liwardana_2023), across multiple values of $k$.
 Based on the silhouette results, we select the $k$ value closest to 1.
-The optimal silhouette scores correspond to $k$=3 for both the average rating trend series $Avg_{(T,P)}$ and the review percentage trend series $Pr_{(T,P)}$.
-We complement this with qualitative validation through manual inspection, confirming that $k$=3 yields the most interpretable and distinct groupings.
+The optimal silhouette scores correspond to kk=3 for both the average rating trend series $Avg_{(T,P)}$ and the review percentage trend series $Pr_{(T,P)}$.
+We complement this with qualitative validation through manual inspection, confirming that kk=3 yields the most interpretable and distinct groupings.
 Step 3: manual trend investigation.
-To interpret the identified clusters, we perform qualitative analysis on random review samples from key time periods.
+To interpret the identified clusters, we perform qualitative analysis on statistically significant random review samples from key time periods.
 For each selected topic category, we identify three representative intervals, typically from the beginning, a period of fluctuation, and the end of the timeline.
@@
 The percentage of Gen-AI reviews increases by 25%, from 18% in 2022-H1 to 43% in 2024-H2, reflecting a growing user focus on Gen-AI features.
+This growth indicates that users are increasingly evaluating apps through the lens of their Gen-AI features.
+As AI becomes a central part of the user experience, developers can no longer treat it as an add-on; users expect it to perform reliably and meaningfully enhance the app’s core value.
+The rising share of Gen-AI reviews therefore signals that user satisfaction and engagement increasingly depend on how well these features are designed, integrated, and maintained.
 Despite this surge in volume, the average rating for Gen-AI topics remains stable over time, as shown in Figure 20 with only a slight change of ± 0.2 stars across the observed period.
@@
 These opposing directional trends counterbalance each other when aggregated, leading to the observed overall stability in Gen-AI topic ratings.
-Figure 18. Evolution of the Percentage of Gen-AI Reviews
-Figure 19. Evolution of the Average Ratings of Gen-AI vs Non-Gen-AI Topics
+The following observations examine these diverging topic-specific trajectories in greater detail.
+Figure 19. Evolution of the Percentage of Gen-AI Reviews (Full dataset considered).
+Figure 20. Evolution of the Average Ratings of Gen-AI vs Non-Gen-AI Topics (Full dataset considered).
 Gen-AI topics analysis.
 As established in our experimental setup, $k=3$ yields the most interpretable and well-separated clusters.
-These clusters reflect three general trends: increasing, decreasing, and stable as shown in Figure 20 and 21 for the average ratings and the percentage of reviews, respectively.
+These clusters reflect three general trends: increasing, decreasing, and stable as shown in Figure 21 and 22 for the average ratings and the percentage of reviews, respectively.
 To deepen our understanding of the observed rating trends, we perform a manual trend investigation, i.e., a qualitative analysis, on three representative topic categories, each corresponding to a distinct trend direction.
@@
 For the Increasing trend, we analyze Content Policy & Censorship to investigate how user perceptions of policy changes have evolved and to uncover the factors contributing to rising satisfaction in this area.
-Figure 20. Trend Clusters for Average Ratings of Gen-AI Topics
-Figure 21. Trend Clusters for Percentage of Gen-AI Topics
+Figure 21. Trend Clusters for Average Ratings of Gen-AI Topics (Full dataset considered).
+Figure 22. Trend Clusters for Percentage of Gen-AI Topics (Full dataset considered).
 Observation 3.2: The Content Quality topic category shows a decline in rating over time, despite improvements in Gen-AI technology.
@@
 This evolution in review content presents a shift from initial awe to critical evaluation, as users become more experimental and less tolerant of limitations.
-We refer to users who review the app in 2022 as early adopters because 2022 is the year ChatGPT was launched (Akhtar, 2024), and we refer to users in 2023 and 2024 as mainstream users as they started using the app as it rose in popularity.
-Early adopters are typically motivated by innovation and are more tolerant of uncertainties and imperfections in emerging technologies than mainstream users (Paulino and Gudmundsson, 2024).
+We refer to users who review the app in 2022 as early adopters because 2022 is the year ChatGPT was launched (Akhtar_2024), and we refer to users in 2023 and 2024 as mainstream users as they started using the app as it rose in popularity.
+Early adopters are typically motivated by innovation and are more tolerant of uncertainties and imperfections in emerging technologies than mainstream users (Paulino2024).
 Hence, the decline in the average rating of Content Quality can be explained by the change in observed characteristics between early adopters and mainstream users.
-$\Rightarrow$ Implication 12.
-Developers entering a competitive Gen-AI market must ensure that their apps are robust, possess refined capabilities and are capable of meeting the higher expectations of mainstream users.
-Over-promising performance can lead to user dissatisfaction and reputational harm, whereas transparent communication about capabilities and limitations can foster trust.
 Observation 3.5: The AI Performance topic exhibits a stable trend in average user ratings over time, characterized by a slight increase in 2023-H1, and followed by a modest decline in 2024-H1.
-To account for this overall stability, we analyze fluctuations at the app category level.
-We identify a compensatory pattern revealing that improvements in user-perceived AI Performance in some categories are offset by declines in others.
-This inter-category balancing effect contributes to the relatively flat trend observed in the aggregate ratings.
+To account for this overall stability, we analyze fluctuations at the app-category level.
+We identify a compensatory pattern revealing that improvements in user-perceived AI Performance in some categories are offset by declines in others, producing the flat global average.
+However, this apparent stability conceals divergent user experiences.
+Categories such as Education and Games show clear improvement, reflecting that users in these domains perceive AI as directly enhancing learning and task performance.
+In contrast, most other categories (e.g., Productivity, Entertainment, Books & Reference) show gradual declines.
+This may suggest that developers in these areas are integrating AI features prematurely, before the technology is fully optimized for their use cases.
+When users accustomed to fast and accurate AI systems encounter slower or less reliable integrations, their satisfaction drops.
+These declines therefore point to gaps in seamless integration and contextual adaptation rather than outright failure of AI Performance.
+For developers, this indicates that AI performance has become a key competitive factor.
+Users now expect fast, accurate, and context-aware AI behavior.
+When AI integrations fall short or introduce lag, inconsistency, or irrelevant responses, users quickly express dissatisfaction.
+Sustaining trust therefore depends on ensuring that AI features operate seamlessly within the app’s core experience rather than as add-ons.
 Figure 24. Examples of reviews discussing AI Performance at different time periods.
@@
 Creative tools are often evaluated for imagination and output quality, while Productivity apps are judged more on efficiency and reliability.
-$\Rightarrow$ Implication 13.
+$\Rightarrow$ Implication 7.
+Developers entering a competitive Gen-AI market must ensure that their apps are robust, possess refined capabilities and are capable of meeting the higher expectations of mainstream users.
+Over-promising performance can lead to user dissatisfaction and reputational harm, whereas transparent communication about capabilities and limitations can foster trust.
 AI integration should be purpose-driven and context-aware.
 Raji et al.
-(Raji et al., 2022) discussed how AI systems are often deployed without sufficient consideration of their actual functionality and utility.
+(Raji_Kumar_Horowitz_Selbst_2022) discussed how AI systems are often deployed without sufficient consideration of their actual functionality and utility.
 The authors argue that the assumption of AI’s inherent value can lead to ineffective or even harmful implementations, stressing the need for critical assessment before integration.
@@
 Observation 3.9: The average user ratings for the Content Policy & Censorship topic follow an increasing trend, with fluctuations in 2022-H2 and 2024-H1.
-We observe that user ratings decrease in 2022-H2, rise steadily through 2023-H2, and decrease slightly in 2024-H1.
 To understand this trend, we examine the raised user concerns in three random samples of user reviews in 2022-H2, 2023-H2, and 2024-H1.
 We find that content policy complaints discuss three main concerns: 1) ethical AI training practices, 2) the effectiveness of content filters, and 3) false-positive censorship.
+This pattern reflects how content policy concerns were especially prominent in the early stages of Gen-AI deployment, when moderation systems were still limited and users frequently encountered unsafe or poorly filtered outputs.
+Over time, developers appear to have become increasingly aware of these issues and implemented more effective safeguards, leading to gradual improvement in user satisfaction.
+The upward trajectory in ratings thus signals a maturing ecosystem where content safety and moderation mechanisms are continuously refined.
+For developers, this emphasizes the importance of iterative moderation design.
+Continuous refinement of content filters not only improves safety but also builds long-term user trust.
 Figure 25. Examples of reviews discussing Content Policy at different time periods.
@@
 Minor setbacks and fluctuations are to be expected as systems evolve and adapt to user needs.
-$\Rightarrow$ Implication 14.
+$\Rightarrow$ Implication 8.
 Developers must strike a careful balance between safety and creative freedom.
@@
 Reviewers provide constructive feedback to co-create better experiences with developers.
-As Gen-AI becomes more embedded in users’ lives, particularly in the emotional or educational contexts, users come to expect empathy, personalization, and reliability (Implications 2, 7, and 10).
+As Gen-AI becomes more embedded in users’ lives, particularly in the emotional or educational contexts, users come to expect empathy, personalization, and reliability (Implications 1, 3, and 5).
 This shift highlights the importance of treating users not merely as consumers, but as stakeholders who help shape and refine Gen-AI apps.
@@
 Users increasingly describe these systems as trusted companions or emotional supports.
-Understanding the psychological and social bonds users form will be crucial for guiding the ethical design, deployment, and regulation of AI in emotionally sensitive contexts (Implication 21).
+Understanding the psychological and social bonds users form will be crucial for guiding the ethical design, deployment, and regulation of AI in emotionally sensitive contexts (Implication 5).
 7.2. Implications for Platform Owners
@@
 One possible approach is to prompt users to answer specific questions or rate distinct topics, such as usability, design, or content accuracy, separately from the overall rating.
+Beyond improving review quality, our findings also reveal structural limitations in how app stores currently support developer–user interaction.
+As shown in our developer response analysis (Implication 6), review threads in Gen-AI apps function primarily as an initial point of contact rather than a space for resolving AI-specific failures.
+Many issues require diagnostic information that cannot be shared publicly or involve model- and policy-level changes that cannot be deployed quickly.
+Platform owners can play an important role by offering richer, privacy-preserving support channels that complement review sections and allow developers to investigate Gen-AI issues more effectively.
+Strengthening these communication pathways would improve user experience and reduce the friction developers face when addressing complex AI behaviours.
 7.3. Implications for Developers
-Developers face growing pressure to meet user expectations that are increasingly shaped by leading models, such as ChatGPT (Implication 8).
-Users demand fast, accurate, and context-aware outputs across a wide range of use cases, including education, professional assistance, and emotional support (Implications 6 and 10).
-Our findings highlight several critical areas for improvement: fine-tuning censorship filters to avoid false positives (Observation 3.9; Implication 4), supporting accessibility and inclusivity concerns (Implication 11), and enabling more personalized and flexible user experiences (Implication 7).
+Developers face growing pressure to meet user expectations that are increasingly shaped by leading models, such as ChatGPT (Implication 4).
+Users demand fast, accurate, and context-aware outputs across a wide range of use cases, including education, professional assistance, and emotional support (Implications 3 and 5).
+Our findings highlight several critical areas for improvement: fine-tuning censorship filters to avoid false positives (Observation 3.9; Implication 2), supporting accessibility and inclusivity concerns (Implication 5), and enabling more personalized and flexible user experiences (Implication 3).
 Developers should also consider offering modular features and customization options, such as adjustable censorship settings, adjustable AI functionalities, and adaptive content formats.
-Importantly, developers must think critically about when and how AI is used, ensuring that integration adds meaningful value, as not every problem or task requires an AI solution (Observation 3.7; Implication 13).
+Importantly, developers must think critically about when and how AI is used, ensuring that integration adds meaningful value, as not every problem or task requires an AI solution (Observation 3.7; Implication 7).
 7.4. Implications for Testing Tool Designers
@@
 7.5. Implications for Requirements Engineers
-Requirements engineers can benefit from the structured insights our topic analysis approach provides.
+Requirements engineers can benefit from the structured insights our topic analysis framework provides.
 The ability to track how user expectations evolve by topic over time can inform roadmap planning and feature prioritization (Observation 3.3).
@@
 Engineers should leverage such review data to validate or revise assumptions about user needs, especially in light of shifting demographics and rising expectations for Gen-AI-powered functionality.
+Beyond Gen-AI apps, the same framework and analytical tools can be applied to user reviews in other domains, enabling engineers to systematically identify evolving requirements and usability concerns across a wide range of software systems.
 7.6. Implications for Policy Makers and Regulators
@@
 While some early users voiced concerns about data ethics and copyright in training data, these concerns faded in later samples (Observation 3.10), which may indicate normalization rather than resolution.
-Policy makers must ensure that ethical standards around content ownership, training transparency, and accountability are formalized and enforced (Implication 4).
+Policy makers must ensure that ethical standards around content ownership, training transparency, and accountability are formalized and enforced (Implication 2).
 Furthermore, content moderation must strike a balance between safety and usability.
-Regulators should develop protocols to mitigate the harm caused by unregulated AI generations (Observation 3.11; Implication 4), particularly as Gen-AI becomes increasingly pervasive in sensitive domains such as education and emotional support.
+Regulators should develop protocols to mitigate the harm caused by unregulated AI generations (Observation 3.11; Implication 2), particularly as Gen-AI becomes increasingly pervasive in sensitive domains such as education and emotional support.
 8. Threats to Validity
@@
 Our analysis focuses on publicly available user reviews from the Google Play Store, which may not fully reflect user experiences on other platforms, such as the Apple App Store.
+This single-platform scope may limit the generalizability of results to other ecosystems with different user bases or review practices.
 We take several steps to enhance the generalizability of our results.
@@
 The rapid integration of Gen-AI features into mobile apps has transformed how users interact with technology.
-In this study, we present a user-centered analysis of Gen-AI mobile apps by examining over 676,000 reviews from 173 apps on the Google Play Store.
+In this study, we present a user-centered analysis of Gen-AI mobile apps by examining over 676,066 reviews from 173 apps on the Google Play Store.
 Using our proposed four-phase SARA methodology and prompt-based LLM techniques, we uncover actionable insights into how users experience and evaluate Gen-AI functionalities.
@@
 expert users) as Gen-AI tools continue and mature.
-Data Availability. We provide a replication package, including manually labelled data, and Python scripts to replicate the analyses at https://github.com/Safwat-UofT/Gen-AI_User_Reviews.
+Data Availability. We provide a replication package, including manually labeled data, and Python scripts to replicate the analyses at omitted.
 Acknowledgements.

### v4

@@
-Understanding the Challenges and Promises of Developing Generative AI Apps: An Empirical Study
+Understanding the Challenges and Opportunities of Generative AI Apps: An Empirical Study
 Abstract.
-The release of ChatGPT in 2022 triggered a rapid surge in generative artificial intelligence mobile apps (i.e., Gen-AI apps).
-Despite widespread adoption, little is known about how end users perceive and evaluate these Gen-AI functionalities in practice.
-In this work, we conduct a user-centered analysis of 676,066 automatically labeled reviews from 173 Gen-AI apps on the Google Play Store.
-We propose a structured four-phase framework, SARA (Selection, Acquisition, Refinement, and Analysis), which integrates and extends state-of-the-art techniques for large-scale review collection, filtering, and analysis using prompt-based LLMs.
-First, we empirically validate the reliability of LLMs for topic extraction and assignment, achieving 91% accuracy through five-shot prompting and LLM-based filtering of non-informative reviews.
-We then apply the framework to informative reviews to identify the ten most discussed topics (e.g., AI Performance, Content Quality, and Content Policy & Censorship) and analyze the key challenges and emerging opportunities.
-Finally, we examine how these topics evolve over time, offering insight into shifting user expectations and engagement patterns with Gen-AI apps.
-Based on our findings and observations, we present actionable implications for developers and researchers.
+The release of ChatGPT in 2022 triggered a rapid surge in generative artificial intelligence mobile apps (Gen-AI apps).
+Despite widespread adoption, little is known about how end users perceive and evaluate these Gen-AI functionalities.
+We conduct a user-centered analysis of 1,035,342 reviews from 171 Gen-AI apps from the Google Play Store.
+We propose SARA (Selection, Acquisition, Refinement, and Analysis), a four-phase framework that leverages prompt-based LLMs for large-scale review analysis.
+We validate the reliability of LLM-based topic extraction and assignment using 4,353 manually evaluated reviews, achieving 91% accuracy with five-shot prompting and filtering of non-informative reviews.
+We identify the top ten topics (e.g., AI Performance and Emotional Connection) and perform a cross-platform comparison with Apple App Store reviews.
+Through qualitative analysis of 762 reviews, we uncover three opportunities (AI for Accessibility and Wellbeing, AI as a Collaborative Creative Tool, and AI Versatility) and three challenges (Managing User Expectations and AI Limitations, Balancing Content Moderation and Creative Freedom, and Strategic Integration of Gen-AI Features).
+Finally, we analyze temporal trends, revealing how user concerns shift as users mature.
+Our findings enable researchers and developers to better leverage the capabilities of Gen-AI apps and address potential challenges.
 Generative AI, Large Language Models, User Review Analysis, Mobile Apps
@@
 1. Introduction
-Mobile applications (apps) are a crucial aspect of people’s lives, and their global data usage continues to rise each year (Rathod_Agal_2023).
-A major shift in the mobile app landscape occurred with the release of ChatGPT [^1] https://play.google.com/store/apps/details?id=com.openai.chatgpt&hl=en, which achieved 7.4 million downloads within its first ten days.
-This rapid adoption signaled strong consumer interest in generative artificial intelligence (Gen-AI) and sparked a surge of similar artificial intelligence (AI) powered apps across app marketplaces.
-In 2024, this trend contributed to a 112% year-over-year increase in downloads of AI chatbot apps (Ceci2025), such as Microsoft Copilot [^2] https://play.google.com/store/apps/details?id=com.microsoft.copilot and DeepSeek [^3] https://play.google.com/store/apps/details?id=com.deepseek.chat.
-Gen-AI refers to a class of AI that generates new and meaningful content, including text, images, audio and video (Feuerriegel_Hartmann_Janiesch_Zschech_2024; 10628027).
-Unlike traditional AI models, which rely on historical data to perform predictions and analytical tasks, Gen-AI models synthesize information from diverse sources to create original outputs (gozalobrizuela2023surveygenerativeaiapplications).
-Among the most prominent forms of Gen-AI are large language models (LLMs), i.e., deep learning models trained on massive text corpora to understand and generate human-like language (Hadi_Tashi_Qureshi_Shah_Muneer_Irfan_Zafar_Shaikh_Akhtar_Wu_et_al_2023).
+Mobile applications (apps) are a crucial aspect of people’s lives, and their global data usage continues to rise each year (Rathod and Agal, 2023).
+A major shift in the mobile app landscape followed the release of ChatGPT [^1] https://play.google.com/store/apps/details?id=com.openai.chatgpt&hl=en, which achieved 7.4 million downloads within its first ten days.
+This rapid adoption reflects strong consumer interest in generative artificial intelligence (Gen-AI) and sparked a surge of similar artificial intelligence (AI) powered apps across app marketplaces.
+In 2024, this trend contributed to a 112% year-over-year increase in downloads of AI chatbot apps (Ceci, 2025), such as Microsoft Copilot [^2] https://play.google.com/store/apps/details?id=com.microsoft.copilot and DeepSeek [^3] https://play.google.com/store/apps/details?id=com.deepseek.chat.
+Gen-AI refers to a class of AI that generates new and meaningful content, including text, images, audio and video (Feuerriegel et al., 2024; Zhang et al., 2024).
+Unlike traditional AI models, which rely on historical data to perform predictions and analytical tasks, Gen-AI models synthesize information from diverse sources to create original outputs (Gozalo-Brizuela and Merchan, 2024).
+Among the most prominent forms of Gen-AI are large language models (LLMs), which are deep learning models trained on massive text corpora to understand and generate human-like language (Hadi et al., 2023).
 LLMs power a wide range of natural language processing (NLP) tasks, including text generation, summarization, translation, and dialogue.
-Developers are increasingly integrating Gen-AI models into mobile apps to enable novel capabilities, such as conversational interfaces, voice interaction, and content generation (Hau_Hassan_Zhou_2025).
-When embedded into apps, these models can deliver personalized and adaptive experiences by providing real-time, context-aware feedback that enhances user engagement and motivation (Yuen_Schlote_2024).
+Developers are increasingly integrating Gen-AI models into mobile apps to enable novel capabilities, such as conversational interfaces, voice interaction, and content generation (Hau et al., 2025).
+When embedded into apps, these models can deliver personalized and adaptive experiences by providing real-time, context-aware feedback that enhances user engagement and motivation (Yuen and Schlote, 2024).
 A key enabler of this integration is the growing availability of Gen-AI-powered application programming interfaces (APIs).
-APIs serve as software intermediaries that enable mobile apps to access language understanding and generation functionalities without having to host or fine-tune models locally (Chen_Liu_Zhou_Zhao_Wang_Wang_Chen_Bissyandé_Klein_Li_2024).
-While the integration of Gen-AI into mobile apps offers clear benefits in functionality and user experience, it also introduces a range of technical and practical challenges.
-Implementing Gen-AI features requires consideration of API integration, prompt engineering, token limitation, inference cost, and the unpredictable model outputs (Chen_Gao_Chen_Zhang_Liu_2025).
-These issues can degrade the quality of Gen-AI-powered apps, leading to functional failures (e.g., crashes, incorrect behaviour, or suboptimal user interfaces), reduced efficiency (e.g., latency or excessive token consumption), and potential security vulnerabilities (Shao_Huang_Shen_Ma_Su_Wan_2024).
+APIs serve as software intermediaries that enable mobile apps to access language understanding and generation functionalities without having to host or fine-tune models locally (Chen et al., 2025a).
+While integrating Gen-AI into mobile apps offers clear benefits to functionality and user experience, it also introduces a range of technical and practical challenges.
+Implementing Gen-AI features requires consideration of API integration, prompt engineering, token limitations, inference costs, and the unpredictable model outputs (Chen et al., 2025b).
+These issues can degrade the quality of Gen-AI-powered apps, leading to functional failures (e.g., crashes, incorrect behavior, or suboptimal user interfaces), reduced efficiency (e.g., latency or excessive token consumption), and potential security vulnerabilities (Shao et al., 2025).
 Additionally, beyond implementation, quality assurance for Gen-AI-powered apps remains a concern.
-Unlike traditional software systems, Gen-AI components often lack formal specifications, making it challenging to evaluate correctness or performance in a deterministic way (Nahar_Kästner_Butler_Parnin_Zimmermann_Bird_2024).
+Unlike traditional software systems, Gen-AI components often lack formal specifications, making it challenging to evaluate correctness or performance in a deterministic way (Nahar et al., 2024).
 As a result, ensuring the reliability and robustness of Gen-AI-integrated mobile apps continues to be a significant challenge for developers.
@@
 For example, Nahar et al.
-(Nahar_Kästner_Butler_Parnin_Zimmermann_Bird_2024) employed interviews and surveys to investigate how developers address Gen-AI integration issues, including latency, energy consumption, cost, and fairness.
+(Nahar et al., 2024) employed interviews and surveys to investigate how developers address Gen-AI integration issues, including latency, energy consumption, cost, and fairness.
 While informative, their work is limited to developer perspectives and does not consider how these technical challenges impact end-user experience.
 Chen et al.
-(Chen_Gao_Chen_Zhang_Liu_2025) analyzed posts from an OpenAI developer forum to uncover challenges in prompting, API use, and plugin development.
-This study offers insight into implementation bottlenecks but also neglects user-facing concerns.
+(Chen et al., 2025b) analyzed posts from an OpenAI developer forum to uncover challenges in prompting, API use, and plugin development.
+Their study offers insight into implementation bottlenecks but also neglects user-facing concerns.
 Similarly, Shao et al.
-(Shao_Huang_Shen_Ma_Su_Wan_2024) examined open-source mobile apps with Gen-AI integration and found that 77% experienced integration challenges, which impacted functionality, performance, or security issues.
+(Shao et al., 2025) examined open-source mobile apps with Gen-AI integration and found that 77% experienced integration challenges, which impacted functionality, performance, or security issues.
 However, their analysis is limited to code-level integration problems and does not investigate whether users are aware of or affected by these issues.
 In contrast, our work examines the user-perceived impact of Gen-AI-enabled apps, using large-scale review data to understand how users perceive Gen-AI features in practice.
-Existing studies on user perspectives of Gen-AI apps tend to focus on specific domains, such as education (Golding_Lippert_Neuschatz_Salomon_Burke_2024; Kim_Yu_Detrick_Li_2025; Lee_Arnold_Srivastava_Plastow_Strelan_Ploeckl_Lekkas_Palmer_2024; Shata_Hartley_2025; Kim_Klopfer_Grohs_Eldardiry_Weichert_Cox_Pike_2025; Yuen_Schlote_2024), creative image generation (Tang_Zhang_Ciancia_Wang_2024; haase2023art; Oppenlaender_Silvennoinen_Paananen_Visuri_2023), or mental health (Jin_Kim_Han_2025; Heinz_Mackin_Trudeau_Bhattacharya_Wang_Banta_Jewett_Salzhauer_Griffin_Jacobson_2025).
+Existing studies on user perspectives of Gen-AI apps tend to focus on specific domains, such as education (Golding et al., 2024; Kim et al., 2025a; Lee et al., 2024a; Shata and Hartley, 2025; Kim et al., 2025b; Yuen and Schlote, 2024), creative image generation (Tang et al., 2024; Haase et al., 2023; Oppenlaender et al., 2023), or mental health (Jin et al., 2025; Heinz et al., 2025).
+In addition, prior work is often limited in scale or scope.
+For example, Alabduljabbar (Alabduljabbar, 2024) analyzes 11,549 reviews from five Gen-AI apps focusing on usability, and Meng et al.
+(Meng et al., 2026) study sentiment across 100,010 reviews from nine apps, while Massenon et al.
+(Massenon et al., 2025) analyze approximately three million reviews from 90 AI-powered apps without focusing specifically on Gen-AI.
 Our work takes a systematic view by analyzing user reviews across a wide range of Gen-AI tools and application areas.
 To our knowledge, this is the first study to systematically examine the full landscape of topics users discuss across diverse Gen-AI domains.
-We identify the technical pain points surfaced by users, in addition to emerging behavioural patterns, expectations, and sources of satisfaction or frustration.
+We identify the technical pain points surfaced by users, in addition to emerging behavioral patterns, expectations, and sources of satisfaction or frustration.
 This user-centered perspective complements developer-focused studies, providing critical insights for improving the usability and effectiveness of Gen-AI mobile apps.
-Figure 1. Examples of reviews of Gen-AI apps
 To gain these insights, we analyze user reviews that offer valuable perspectives on the challenges, expectations, and satisfaction associated with Gen-AI apps.
 For example, Figure 1 illustrates representative reviews from three Gen-AI apps, highlighting both praise and frustration.
-Building on prior work in user review mining and LLM-based topic analysis (Maram_LLM_Cure; Chen_Lin_Hoi_Xiao_Zhang_2014), we propose a structured four-phase framework called Selection, Acquisition, Refinement, and Analysis (SARA).
+Building on prior work in user review mining and LLM-based topic analysis (Assi et al., 2025a; Chen et al., 2014), we propose a structured four-phase framework called Selection, Acquisition, Refinement, and Analysis (SARA).
 This framework enables the systematic collection, cleaning, and analysis of user reviews from Gen-AI apps, demonstrating the feasibility of large-scale, reproducible studies using LLMs.
+While validated in the context of Gen-AI apps, SARA’s modular design makes it adaptable to the analysis of user reviews extracted from app store platforms across multiple mobile app domains.
+Figure 1. Examples of reviews of Gen-AI apps.
+A figure showing sample user reviews from Gen-AI mobile apps.
+Review 1 (from the ChatGPT app, rated 4 stars): “This app works fantastic!
+The only drawback i see in this is the memory of the AI.
+Chat only remembers the commands for the next two conversation after you give it a command.” Review 2 (from the AI Mirror: AI Photo & Video app, rated 2 stars): “It showed my dark skined friend as white and we tried 5 times its really annoying.” Review 3 (from GoArt - Ghibli Style AI Image app, rated 1 stars): “It’s terrible the art style didn’t came out the way it should’ve and the details didn’t came out as I expect it and it looks like something it picks up from a Google search.”
 First, we identify relevant Gen-AI apps (i.e., Selection) and collect user reviews from the Google Play Store (i.e., Acquisition).
@@
 Finally, we employ prompt-based LLM techniques to analyze the reviews (i.e., Analysis).
-Using this integrative framework, we automatically label and quantitatively analyze 676,066 user reviews from 173 Gen-AI apps using LLMs, complemented by qualitative analyses on statistically representative samples.
-Before analyzing Gen-AI user reviews (RQ2–RQ3), we first evaluate the accuracy and reliability of our LLM-based framework (RQ1) to establish its validity for large-scale analysis.
-This ensures that the topic extraction and assignment process used in the subsequent analyses is accurate and trustworthy.
+Using this integrative framework, we automatically label and quantitatively analyze 1,035,342 user reviews from 171 Gen-AI apps using LLMs, complemented by qualitative analyses on statistically representative samples.
+Our research questions (RQs) are organized as follows: we first evaluate the accuracy and reliability of our LLM-based framework for extracting and assigning topics from Gen-AI app reviews (RQ1).
+We then apply the framework to analyze review topics and their evolution over time (RQ2 and RQ3).
+We study the following RQs.
 RQ1: How accurately can LLMs identify topics in user reviews of Gen-AI apps?
 In this RQ, we investigate how prompt configuration (e.g., using 0-shot, 3-shot, and 5-shot) and input quality (e.g., the proportion of informative reviews) affect the performance of LLMs in extracting topics from user reviews.
-Our results show that increasing the number of few-shot examples and filtering non-informative reviews significantly improve accuracy.
+Our results show that increasing the number of few-shot examples and filtering non-informative reviews significantly improves accuracy.
 The best-performing setup achieves an overall accuracy of 91%, confirming the reliability of our LLM framework for downstream analysis.
-RQ2: What are the most prominent topics discussed in user reviews and in developer responses of Gen-AI apps?
-To answer this RQ, we apply the LLM-based topic labeling framework from RQ1 to informative reviews across 173 Gen-AI apps.
-We analyze the top 10 topics in Gen-AI app reviews to understand user behaviour, key challenges and opportunities, and derive actionable insights for developers and researchers.
-While our primary focus is on user reviews, we also include a small supplementary examination of developer responses to better contextualize how developers engage with Gen-AI feedback.
-Our findings reveal that users tend to rate Gen-AI topics positively and expect high performance, often providing constructive feedback from a collaborative mindset.
-Notably, some users form emotional bonds with conversational agents and value personalized interactions.
-These insights have practical implications for developers aiming to improve the usability, engagement, and trustworthiness of Gen-AI features.
+RQ2: What are the most prominent topics discussed in user reviews?
+To answer this RQ, we analyze the top 10 topics in Gen-AI app reviews to understand user behavior, key challenges and opportunities, and derive actionable insights for developers and researchers.
+Our findings reveal three key opportunities: AI for Accessibility and Wellbeing, AI as a Collaborative Creative Tool, and AI Versatility, and three challenges: Managing User Expectations and AI Limitations, Balancing Content Moderation and Creative Freedom, and Strategic Integration of Gen-AI Features.
 RQ3: What temporal trends emerge in user feedback on Gen-AI topics?
-In this RQ, we explore how user feedback on Gen-AI topics evolves as the technology matures.
-We cluster topics to identify evolutionary trends and support the analysis by qualitative review interpretation.
-Our analysis reveals that while the average rating for Gen-AI topics remains high, topic trends diverge: Content Quality ratings decline due to rising user expectations, AI Performance remains stable due to balancing effects, and Content Policy & Censorship ratings improve with better filtering mechanisms with time.
-These shifts reflect an increasingly critical and experienced user base that raises expectations for Gen-AI apps over time.
+In this RQ, we explore how user feedback on Gen-AI topics evolves as the technology matures for two Gen-AI topics; Emotional Connection, and Content Policy & Censorship.
+Our analysis reveals that reviews evolve from initial novelty in emotional interactions to dependency and addiction risks, and from early trust and privacy issues to broader tensions around content censorship and bias.
 Our study provides valuable insights for app developers and researchers by uncovering how users perceive and evaluate Gen-AI features in mobile apps. Specifically, our contributions are as follows:
   • We propose a structured four-phase framework, SARA, that builds upon established techniques in user review analysis and LLM-based topic modeling.
-Using SARA, we conduct a large-scale empirical study of 676,066 reviews from 173 Gen-AI apps on the Google Play Store to 1) identify the most frequently discussed Gen-AI-related topics, 2) uncover user expectations and pain points, and 3) analyze temporal trends to reveal how user perceptions evolve.
-While validated in the context of Gen-AI apps, SARA’s modular design makes it generalizable to analyzing user feedback in other mobile app domains and AI-integrated systems.
+Using SARA, we conduct a large-scale empirical study of 1,035,342 reviews from 171 Gen-AI apps on the Google Play Store, complemented by a cross-platform comparison with Apple App Store (App Store) apps to assess generalizability.
   • We evaluate the effectiveness of LLMs for topic extraction and assignment of user reviews, demonstrating that increasing the number of labeled examples and filtering out non-informative reviews significantly improves accuracy.
-  • We provide actionable feedback for developers to maintain competitiveness, for policymakers to explore the ethical concerns surrounding content policy and censorship, and for user requirement engineers to better understand user needs.
+  • We conduct a large-scale quantitative analysis of user reviews across Gen-AI app categories, extracting 10 representative topics to identify three key opportunities and three challenges.
+  • We provide actionable insights for developers to maintain competitiveness, for policymakers to examine ethical concerns such as content policy and censorship, and for requirements engineers to better understand user needs.
+  • We conduct a temporal analysis of user reviews to examine how user perceptions and behaviors toward Gen-AI apps evolve over time, revealing a shift from early novelty and trust toward more complex concerns such as emotional dependency, system limitations, and content moderation.
   • We provide a replication package that includes the dataset of Gen-AI apps, three manually labeled samples for evaluating the filtering prompt, and the complete set of code we used to run all prompt-based experiments presented in this paper.
@@
 Section 3 presents our methodology.
-Sections 46 describe the experiment design and the findings obtained for each RQ.
-Section 7 synthesizes our key findings and implications of our work.
+Sections 4-6 describe the findings obtained for each RQ.
+Section 7 presents a platform comparison and developers’ response analysis.
+Section 8 synthesizes our key findings and frames them into implications and actionable insights.
 Section 9 discusses potential threats to the validity of our findings.
@@
 2.1. User Perspective of Gen-AI Apps
-The user’s perspective on utilizing Gen-AI tools in education has been well studied (Golding_Lippert_Neuschatz_Salomon_Burke_2024; Kim_Yu_Detrick_Li_2025; Lee_Arnold_Srivastava_Plastow_Strelan_Ploeckl_Lekkas_Palmer_2024).
-Shata and Hartley (Shata_Hartley_2025) used surveys grounded in the Technology Acceptance Model (TAM) and Social Cognitive Theory (SCT) to examine faculty attitudes, identifying trust as the key driver of Gen-AI adoption.
+The user’s perspective on the use of Gen-AI tools in education has been well studied (Golding et al., 2024; Kim et al., 2025a; Lee et al., 2024a).
+Shata and Hartley (Shata and Hartley, 2025) used surveys grounded in the Technology Acceptance Model (TAM) and Social Cognitive Theory (SCT) to examine faculty attitudes and identified trust as the key driver of Gen-AI adoption.
 Kim et al.
-(Kim_Klopfer_Grohs_Eldardiry_Weichert_Cox_Pike_2025) compared student and faculty perceptions on the use of Gen-AI in learning.
+(Kim et al., 2025b) compared student and faculty perceptions on the use of Gen-AI in learning.
 Their study revealed that students and faculty used Gen-AI at similar rates, while students found Gen-AI tools easier to learn.
 Both faculty and students believed that Gen-AI will have more negative than positive effects on learning.
-Yuen and Schlote (Yuen_Schlote_2024) surveyed adult learners to explore their views on the integration of AI in language learning apps.
-Their participants expressed a desire for in-app social features to assist in learning a language, and the authors suggested AI integration in the form of AI tutors that act as conversation partners.
+Yuen and Schlote (Yuen and Schlote, 2024) surveyed adult learners to explore their views on integrating AI into language learning apps.
+Their participants expressed a desire for in-app social features to assist with language learning, and the authors suggested integrating AI tutors that act as conversation partners.
 In the creative image domain, Tang et al.
-(Tang_Zhang_Ciancia_Wang_2024) surveyed users of Gen-AI image tools focusing on satisfaction, challenges, acceptance, perceptions, and applicability of Gen-AI image tools.
+(Tang et al., 2024) surveyed users of Gen-AI image tools focusing on satisfaction, challenges, acceptance, perceptions, and applicability of Gen-AI image tools.
 They found that users most frequently rely on these tools during the early ideation stage rather than for producing final outputs.
 In support of that, Haase et al.
-(haase2023art) examined how Gen-AI images influence creative idea generation, and demonstrated that Gen-AI art effectively supports ideation.
+(Haase et al., 2023) examined how Gen-AI images influence creative idea generation, and demonstrated that Gen-AI art effectively supports ideation.
 Oppenlaender et al.
-(Oppenlaender_Silvennoinen_Paananen_Visuri_2023) focused more on the user’s understanding of technology, fears, and thoughts on the risks of Gen-AI image tools in their survey.
+(Oppenlaender et al., 2023) focused more on the user’s understanding of technology, fears, and thoughts on the risks of Gen-AI image tools in their survey.
 Their participants voiced concerns over a societal risk, as people could use these tools to generate images that spread misinformation.
-In mental health, applications of AI include detection, diagnosis, treatment, and ongoing condition management (10842666; 10963316; 10985281; 10911418).
-Our focus is on Gen-AI-driven treatment through conversational agents.
+In mental health, applications of AI include detection, diagnosis, treatment, and ongoing condition management (Sharma and Patel, 2024; Madhuri et al., 2025; S et al., 2025; Kheterpal and Gill, 2024).
 Jin et al.
-(Jin_Kim_Han_2025) reviewed barriers to sustained engagement in mental health apps and suggested integrating AI in the form of AI agents that can detect emotional changes in users, build strong relationships with users, and promote timely intervention.
+(Jin et al., 2025) reviewed barriers to sustained engagement in mental health apps and suggested integrating AI in the form of AI agents that can detect emotional changes in users, build strong relationships with users, and promote timely intervention.
 Heinz et al.
-(Heinz_Mackin_Trudeau_Bhattacharya_Wang_Banta_Jewett_Salzhauer_Griffin_Jacobson_2025) studied the efficacy of a mental health AI chatbot.
+(Heinz et al., 2025) studied the efficacy of a mental health AI chatbot.
 They found that it significantly reduced users’ symptoms, maintained high engagement, and achieved user acceptance, with a reported therapeutic alliance comparable to that of human therapists.
-Prior studies on user perceptions of Gen-AI apps have primarily relied on interviews and surveys to capture attitudes toward adoption, trust, and usefulness within specific domains such as education (Shata_Hartley_2025; Kim_Klopfer_Grohs_Eldardiry_Weichert_Cox_Pike_2025) and image generation (Tang_Zhang_Ciancia_Wang_2024; Oppenlaender_Silvennoinen_Paananen_Visuri_2023).
+Prior studies on user perceptions of Gen-AI apps have primarily relied on interviews and surveys to capture attitudes toward adoption, trust, and usefulness within specific domains such as education (Shata and Hartley, 2025; Kim et al., 2025b) and image generation (Tang et al., 2024; Oppenlaender et al., 2023).
 These works offer valuable insight into how users conceptualize Gen-AI, but they often reflect processed opinions rather than immediate experiences.
@@
 This perspective complements survey-based research by revealing practical concerns and responses as they naturally occur.
-For example, while prior studies in mental health (Jin_Kim_Han_2025; Heinz_Mackin_Trudeau_Bhattacharya_Wang_Banta_Jewett_Salzhauer_Griffin_Jacobson_2025) demonstrated that conversational AI can foster engagement and therapeutic rapport, our findings extend this understanding by showing that users explicitly value Gen-AI’s perceived neutrality, emotional support, and human-like realism.
+For example, while prior studies in mental health (Jin et al., 2025; Heinz et al., 2025) demonstrated that conversational AI can foster engagement and therapeutic rapport, our findings extend this understanding by showing that users explicitly value Gen-AI’s perceived neutrality, emotional support, and human-like realism.
 Furthermore, while earlier work examined single domains in isolation, our study provides a holistic view across multiple app categories, revealing shared topics and concerns that transcend context.
-By organizing insights by functional features rather than by domain, our findings highlight common usability, quality, and ethical challenges that characterize Gen-AI technologies more broadly.
+By organizing insights by opportunities and challenges rather than by domain, our findings highlight common usability, quality, and ethical challenges that characterize Gen-AI technologies more broadly.
 2.2. User Review Analysis Techniques
 In this section, we summarize prior work related to 1) filtering user reviews, 2) pre-LLM approaches to topic extraction, 3) LLM-based topic extraction and assignment methods, 4) single-task topic analysis, and 5) domain-specific applications of LLMs in user review analysis.
-These studies collectively form the foundation for our own framework SARA, which builds upon their methods but advances them by integrating filtering, topic extraction, and topic assignment into a unified, scalable framework for analyzing Gen-AI app reviews.
+These studies collectively form the foundation for our own framework, SARA, which builds upon their methods while advancing them by integrating filtering, topic extraction, and topic assignment into a unified, scalable framework for analyzing Gen-AI app reviews.
 Filtering user reviews.
@@
 Chen et al.
-(Chen_Lin_Hoi_Xiao_Zhang_2014) introduced a framework called AR-miner that used Latent Dirichlet Allocation (LDA) to first filter informative reviews with a hit-rate of 70%, and second to group the informative reviews using topic modeling.
+(Chen et al., 2014) introduced a framework called AR-miner that used Latent Dirichlet Allocation (LDA) to first filter informative reviews with a hit-rate of 70%, and second to group the informative reviews using topic modeling.
 Ghosh et al.
-(Ghosh_Pargaonkar_Eisty_2024) evaluated fine-tuned LLMs for filtering non-informative user reviews on the Google Play Store and found them highly effective, achieving accuracies of 92% for both BERT and DistilBERT, and 91% for Gemma.
+(Ghosh et al., 2024) evaluated fine-tuned LLMs for filtering non-informative user reviews on the Google Play Store and found them highly effective, achieving accuracies of 92% for both BERT and DistilBERT, and 91% for Gemma.
 Al Wahshat et al.
-(Al_Wahshat_Abu_ulbeh_Yusoff_Zakaria_Amir_Fazamin_Wan_Hamzah_P_2023) highlighted the potential of using GPT-4 in filtering manipulated reviews due to its powerful language processing capabilities.
+(Al Wahshat et al., 2023) highlighted the potential of using GPT-4 in filtering manipulated reviews due to its powerful language processing capabilities.
 Building on this line of work, our study integrates automated filtering as the first step of a unified, LLM-assisted framework that combines filtering with topic extraction and assignment.
 While prior work primarily evaluated filtering accuracy in isolation, we demonstrate that incorporating this process within a broader analysis framework improves data quality and supports large-scale review interpretation.
-Our filtering component achieves comparable accuracy (90%) while operating at a much larger scale across diverse Gen-AI app categories.
+Our filtering component achieves comparable accuracy (91%) while operating at a much larger scale across diverse Gen-AI app categories.
 Pre-LLM approaches to topic extraction.
 Before the advancement of LLMs, researchers used various methods to analyze user reviews.
-Genc Nayebi and Abran (Genc-Nayebi_Abran_2017) conducted a literature review in 2016 that explored approaches for mining opinions from user reviews.
+Genc Nayebi and Abran (Genc-Nayebi and Abran, 2017) conducted a 2016 literature review exploring approaches to mining opinions from user reviews.
 They identified three studies that relied on manual analysis, while others used automated techniques, such as Classification and Regression Trees (CART), word frequency statistics, correlation analysis, LDA, and Hexbin plots.
@@
 For example, Arambepola et al.
-(10499727) used the Valence Aware Dictionary and sEntiment Reasoner (VADER) for sentiment analysis to categorize reviews into positive, neutral, and negative categories.
+(Arambepola et al., 2024) used the Valence Aware Dictionary and sEntiment Reasoner (VADER) for sentiment analysis to categorize reviews as positive, neutral, or negative.
 They then used an LDA model for topic modeling of user reviews of educational apps.
@@
 Assi et al.
-(Assi_Hassan_Tian_Zou_2021) proposed the Global-Local sensitive Feature Extractor (GLFE) to identify high-level features in user reviews, achieving 79–82% precision and 74–77% recall on annotated review data.
+(Assi et al., 2021) proposed the Global-Local sensitive Feature Extractor (GLFE) to identify high-level features in user reviews, achieving 79-82% precision and 74-77% recall on annotated review data.
 These early techniques established the foundation for automated review analysis but relied heavily on statistical or lexicon-based models.
@@
 Arambepola et al.
-(Arambepola_Lalendra_Wimalasena_Munasinghe_2025) conducted a recent systematic review of mobile app review analysis techniques and found that integrating LLMs into app user review analysis provides a more holistic and context-rich approach to evaluating reviews and understanding users.
+(Arambepola et al., 2025) conducted a recent systematic review of mobile app review analysis techniques and found that integrating LLMs into app user review analysis provides a more holistic and context-rich approach to evaluating reviews and understanding users.
 Pham et al.
-(Pham_Hoyle_Sun_Resnik_Iyyer_2024) introduced TopicGPT, a framework to perform topic extraction by prompting Open AI’s GPT (openai2023gpt) to generate topics and assign them to documents.
+(Pham et al., 2024) introduced TopicGPT, a framework to perform topic extraction by prompting Open AI’s GPT (OpenAI, 2023) to generate topics and assign them to documents.
 Their framework outperformed traditional topic models, such as LDA, SeededLDA, and BERTopic.
 Prakash et al.
-(Prakash_Wang_Hoang_Hee_Lee_2023) introduced PromptMTopics, a prompt-based ChatGPT model set to extract topics from textual meme captions that outperformed baseline models.
+(Prakash et al., 2023) introduced PromptMTopics, a prompt-based ChatGPT model set to extract topics from textual meme captions that outperformed baseline models.
 While these studies evaluated LLM-based extraction on textual datasets such as documents or memes, our work applies similar techniques directly to large-scale user reviews.
 Assi et al.
-(Maram_LLM_Cure) introduced LLM-Cure, an approach that extracts and assigns features from user reviews to make suggestions for feature improvement.
+(Assi et al., 2025a) introduced LLM-Cure, an approach that extracts and assigns features from user reviews to make suggestions for feature improvement.
 Their approach outperformed baseline models, achieving an F1-score of 85%.
@@
 For example, Wei et al.
-(Wei_Courbis_Lambolais_Xu_Bernard_Dray_2023) focused on review assignment using predefined labels rather than extracting topics.
+(Wei et al., 2023) focused on review assignment using predefined labels rather than extracting topics.
 They assigned reviews using fixed categories such as feature requests, problem reports, and irrelevant feedback.
@@
 Dos Santos et al.
-(Dos_Santos_Oliveira_De_Jesus_Aljedaani_Eler_2023) and Ren et al.
-(Ren_Nakagawa_Tsuchiya_2024) extracted topics from user reviews without performing the assignment task.
+(Dos Santos et al., 2023) and Ren et al.
+(Ren et al., 2024) extracted topics from user reviews without performing the assignment task.
 While this approach is effective for uncovering topical insights, it does not account for the frequency or distribution of each topic, limiting its usefulness for trend or prevalence analysis.
@@
 Beyond general topic extraction, some studies applied LLMs to narrow domains of app analysis.
-Sorathiya and Ginde (Sorathiya_Ginde_2024) employed a hybrid approach combining Natural Language Inference (NLI) and LLMs to identify privacy concerns in mental health app reviews, uncovering ethically relevant feedback that keyword-based methods often miss.
+Sorathiya and Ginde (Sorathiya and Ginde, 2024) employed a hybrid approach that combined Natural Language Inference (NLI) with LLMs to identify privacy concerns in mental health app reviews, uncovering ethically relevant feedback that keyword-based methods often miss.
 Rezaei Nasab et al.
-(RezaeiNasab_Dashti_Shahin_Zahedi_Khalajzadeh_Arora_Liang_2025) examined fairness concerns in AI-based mobile apps, using LLM and machine learning based classifiers to detect fairness mentions and cluster them into six categories.
+(Rezaei Nasab et al., 2025) examined fairness concerns in AI-based mobile apps, using LLM and machine learning-based classifiers to detect fairness mentions and cluster them into six categories.
 While these studies reveal how LLMs can surface ethics-related insights within specific domains, our dataset and focus differ substantially.
-We analyze user reviews from a large-scale collection of Gen-AI apps spanning multiple functional categories, including but not limited to health & fitness, education, productivity, and entertainment.
+We analyze user reviews from a large-scale collection of Gen-AI apps across multiple functional categories, including education, productivity, and entertainment.
 Unlike prior domain-specific datasets centered on single themes such as privacy or fairness, our corpus captures a broad range of emergent Gen-AI topics (i.e.
-AI performance, content quality, and content policy) that are unique to this new class of applications.
+AI performance, content quality, and content policy) that are unique to this new class of apps.
 This wider scope allows us to identify both category-specific differences and cross-domain similarities, showing how users across varied contexts often express comparable concerns and expectations about Gen-AI functionality.
@@
 To study user feedback on Gen-AI apps, we propose and validate a structured four-phase framework we call SARA, which stands for Selection, Acquisition, Refinement, and Analysis.
-SARA encompasses 1) the selection of relevant Gen-AI apps, 2) the acquisition of user reviews, 3) a refinement data cleaning framework to ensure data quality for LLM analysis, and finally, 4) the LLM-based analysis of user reviews through topic extraction and assignment.
-SARA integrates and extends state-of-the-art methods in user review analysis (Maram_LLM_Cure; Chen_Lin_Hoi_Xiao_Zhang_2014), combining automated filtering, topic extraction, and topic assignment into a unified, empirically validated workflow that supports large-scale, LLM-driven analysis of user feedback.
-Although evaluated within the Gen-AI domain, SARA’s modular design enables its application to other mobile or AI-enabled systems for analyzing large-scale user feedback.
-We first evaluate the reliability of SARA in accurately labeling review topics (RQ1) and then apply it in subsequent analyses to explore user concerns and temporal trends (RQ2–RQ3).
 We present an overview of our research framework in Figure 2.
 3.1. Selection of Gen-AI apps
-Our app selection process involves three steps: 1) generating keywords from prior work, 2) searching the Google Play Store and scraping relevant apps, and 3) manually screening and validating each app to confirm it is Gen-AI.
+Our app selection process involves four steps: 1) generating keywords from prior work, 2) searching the Google Play Store and scraping relevant apps, 3) manually screening and validating each app to confirm that it satisfies our definition of a Gen-AI app as introduced earlier, and 4) finding the App Store version of that app.
+Specifically, we retain only apps whose core functionality involves generating novel content (e.g., text, images, audio, or video) and exclude apps that rely solely on AI for prediction, analysis, or classification without generating new content.
+Figure 2. Overview of our framework.
+An overview diagram of the SARA framework.
+The process begins with the Google Play Store, where Gen-AI apps are selected through four steps: (1) generating keywords, (2) scraping Google Play Store apps, (3) screening and validating apps, and (4) identifying corresponding App Store apps.
+This selection process results in 194 confirmed Gen-AI apps.
+User reviews are then collected, yielding 4,312,168 reviews from 193 apps.
+These reviews are refined through a multi-stage filtering process, resulting in 1,035,342 informative reviews from 171 apps.
+Finally, the refined dataset is analyzed through three research questions: (RQ1) designing prompts and evaluating accuracy, (RQ2) identifying and analyzing Gen-AI topics, and (RQ3) examining the evolution of these topics over time.
   1. 1.
-    Keyword generation.
-We compile a list of keywords derived from a Gen-AI taxonomy by Gozalo-Brizuela and Garrido-Merchán (gozalobrizuela2023surveygenerativeaiapplications).
-The authors organized Gen-AI apps into 14 categories, each with a taxonomy of the relevant technology.
-We exclude four categories, i.e.,“biotech,” “brain,” “drug discovery,” and “other,” because an Android mobile app user does not typically use them.
-We create a list of ten categories relevant to mobile apps and repurpose the taxonomy as keywords, as shown in Table 1.
-    Figure 2. Overview of our framework
+    Generate Keywords.
+We compile a list of keywords derived from a Gen-AI taxonomy proposed by Gozalo-Brizuela and Garrido-Merchán (Gozalo-Brizuela and Merchan, 2024).
+We adapt this taxonomy to mobile app contexts by selecting relevant categories and deriving corresponding keywords.
+The complete list of categories and keywords is provided in Appendix A.
   2. 2.
-    App scraping.
+    Scrape Google Play Store Apps.
 Using our list of keywords, we scrape the Google Play Store to identify all matching apps.
-Some keywords (e.g.,“Image Editing” and “Business”) are too broad, so we append “Generative AI” for all keywords when searching the Google Play Store.
 For each identified app, we collect descriptive metadata, including app ID, app name, Google Play Store category, star rating, number of downloads, number of users who rated the app, and the app description.
 In this paper, we refer to Google Play Store app categories as app categories.
+Our app data collection occurred in October 2024.
   3. 3.
-    Manual screening and validation.
+    Screen and Validate Apps.
 The initial search yields 303 apps.
-To ensure the inclusion of only genuine Gen-AI apps, the first and second authors independently review the apps by looking at the title, description, and images of the apps.
-We assess the inter-rater reliability, using Cohen’s Kappa (cohen1960kappa) ($\kappa=0.61$), which indicates substantial agreement.
-In case of disagreement, the two authors discuss and resolve the discrepancies for the ambiguous cases.
+To ensure the inclusion of only genuine Gen-AI apps, the first and second authors independently review the apps by examining their titles, descriptions, and images.
+We assess the inter-rater reliability, using Cohen’s Kappa (Cohen, 1960) ($\kappa=0.61$), which indicates substantial agreement.
+In case of disagreement, the two authors discuss and resolve discrepancies in ambiguous cases.
 This rigorous process results in a final list of 194 confirmed Gen-AI apps.
-Table 1. Categories and keywords for app search in the Google Play Store
-Category Keywords
-Text Conversational AI, Text-to-Science, Text-to-Author Simulation, Text-to-Medical Advice, Text-to-Itinerary, Doc-to-Text
-Image Image Editing, Artistic Images, Realistic Images
-Video Text-to-Video, Video Production, Image-to-Video
-3D Text-to-3D
-Code and Software Text-to-Code, Text-to-Multilingual Code
-Speech Text-to-Speech, AI Understanding, Speech-to-Text
-Business Business
-Marketing Marketing, Advertisement, Accounting
-Gaming Video game Creation
-Music Music Generator
+  4. 4.
+    Identify the Corresponding App Store Apps.
+Our empirical study primarily focuses on user reviews from the Google Play Store.
+However, we also identify the corresponding App Store versions of these apps in order to examine whether the identified themes generalize across platforms.
+Out of the 194 Gen-AI apps from the Google Play Store, we identify the App Store version using AppMagic  [^4] https://appmagic.rocks/top-charts/apps?tag=243926&store=1&date=2024-10-01&topDepth=1000.
+We identify 81 App Store apps in January 2026.
 3.2. Acquisition of User Reviews
-Following app selection, we collect user reviews using the Google Play Store Scraper (googleplayscraper).
-For each retrieved review, the crawler extracts a comprehensive set of metadata, including reviewId, userName, userImage, content (i.e., review text), score (i.e., user rating from 1 to 5 stars), thumbsUpCount, reviewCreatedVersion, at (i.e., timestamp of the review), replyContent, repliedAt, and appVersion.
-Our subsequent analysis primarily focuses on the content, score, and at fields.
+Following app selection, we collect user reviews using the Google Play Store Scraper (Mingyu, n.d.).
+For each retrieved review, the crawler extracts a comprehensive set of metadata, including content (i.e., review text), score (i.e., user rating from 1 to 5 stars), at (i.e., timestamp of the review), replyContent (developer’s response), and repliedAt (i.e., timestamp of response).
 We organize the collected reviews based on the Google Play Store categories (app categories) of their respective apps.
-At the end of this step, we collect 2,778,643 reviews of 186 apps, as eight apps from our list have no textual reviews.
+At the end of this step, we collect 4,312,168 reviews of 193 apps.
+User reviews from the App Store apps are collected using the Appbot platform (Appbot, 2026).
+We extract the same review metadata as Google Play Store reviews.
+At the end of this step, we collect 540,119 reviews from 70 apps.
+We map App Store reviews to Google Play Store app categories to enable later cross-platform contextualization.
+Review acquisition occurred in February 2026.
 3.3. Multi-Stage User Review Refinement
-To ensure the robustness and reliability of our LLM-based analysis, we implement a multi-stage data cleaning framework designed to enhance the quality of the user review data progressively.
+To ensure the robustness and reliability of our LLM-based analysis, we implement a multi-stage data-cleaning framework that progressively improves the quality of user review data.
 This multi-stage refinement process comprises four distinct stages.
 We refine data in stages because we are interested in experimenting with different levels of data cleaning in RQ1.
-We summarize the number of reviews retained after each cleaning stage in Table 2.
+We summarize the number of reviews retained after each cleaning stage in Table 1 for Google Play Store reviews and in Table 2 for App Store reviews.
 Each stage builds upon the previous one as follows:
-Table 2. Number of reviews by app category and stage (Full dataset considered).
-App category Stage 0 Stage 1 Stage 2 Stage 3
-Productivity 362,210 354,829 233,311 169,119
-Photography 1,035,946 430,490 188,382 126,277
-Art & Design 332,881 238,291 157,320 125,362
-Entertainment 205,467 205,467 156,636 118,669
-Education 101,812 95,590 57,680 41,905
-Video Players & Editors 364,124 109,610 51,182 36,507
-Tools 231,865 104,636 48,807 35,046
-Music & Audio 17,875 17,875 11,781 9,150
-Health & Fitness 41,537 10,368 8,747 7,849
-Games 29,922 6,254 5,555 5,070
-Books & Reference 2,071 2,071 1,423 1,112
-Total 2,725,710 1,575,481 920,824 676,066
+Table 1. Number of Google Play Store reviews by app category and stage (Full dataset considered).
+App category Processing Stages
+Stage 0 Stage 1 Stage 2 Stage 3
+Basic cleaning Temporal filtering short review exclusion LLM filtering
+Productivity 903,719 817,968 599,382 402,055
+Photography 1,349,480 613,680 284,065 170,446
+Entertainment 431,740 257,237 205,108 152,139
+Art & Design 509,505 302,899 174,911 125,375
+Education 156,435 142,047 95,254 65,699
+Video Players & Editors 476,730 136,542 62,807 39,972
+Tools 271,218 94,374 52,070 36,694
+Games 35,853 34,850 29,648 25,211
+Music & Audio 48,667 44,487 25,781 17,751
+Total 4,183,347 2,444,084 1,529,026 1,035,342
+Table 2. Number of App Store reviews by app category and stage (Full dataset considered).
+App category Processing Stages
+Stage 0 Stage 1 Stage 2 Stage 3
+Basic cleaning Temporal filtering short review exclusion LLM filtering
+Productivity 147,288 136,807 122,580 78,767
+Photography 109,752 26,639 23,028 15,375
+Entertainment 60,939 37,987 36,522 26,777
+Art & Design 74,311 51,700 48,062 33,556
+Education 117,752 115,121 92,002 56,944
+Video Players & Editors 9,763 7,608 6,848 4,420
+Tools 4,329 3,826 3,604 2,504
+Games 5,328 5,083 4,910 3,774
+Music & Audio 7,317 7,016 5,930 3,352
+Total 536,779 391,787 343,486 225,469
 Stage 0: Basic cleaning.
-Text preprocessing, as demonstrated in prior work (Ghosh_Pargaonkar_Eisty_2024; Prakash_Wang_Hoang_Hee_Lee_2023; Roumeliotis_Tselikas_Nasiopoulos_2024; Lee_Lee_Seo_2024), is a standard practice for eliminating noise and ensuring more effective analysis.
-It involves the removal of punctuation, special characters, extra white spaces, emojis, and irrelevant data such as usernames and URLs.
-In the initial stage, we exclude empty reviews and apps with no reviews, as we are interested in analyzing the textual content of the reviews.
-Then, we perform fundamental text preprocessing, including replacing commas with white spaces and removing special characters and emojis from the review content.
-We exclude five categories that contain fewer than 1,000 reviews, as small sample sizes tend to produce repetitive and low-quality topics.
-To validate this threshold, we experiment with categories of varying sizes and find that categories with fewer than 1,000 reviews yield topics that lack diversity and informativeness.
-For instance, the Personalization app category, which contains only 23 reviews, produces two nearly identical and non-informative topics labeled Positive Feedback and Praise.
-Additionally, some topics focus on non-informative content, such as Personal Comments or Signatures, which capture examples like “hello marlyn” and “Placau!!!!” that are not relevant to the app’s functionality or performance.
-Based on these observations, we adopt 1,000 reviews as the minimum threshold and exclude any app categories that fall below this cutoff.
+Text preprocessing, as demonstrated in prior work (Ghosh et al., 2024; Prakash et al., 2023; Roumeliotis et al., 2024; Lee et al., 2024b), is a standard practice for eliminating noise and ensuring more effective analysis.
+In the initial stage, following prior work, we normalize whitespace and Unicode characters; remove emojis, URLs, and usernames; replace commas with whitespace; convert to lowercase; and drop empty reviews.
 Stage 1: Temporal filtering.
-This stage filters reviews written prior to 2022 to ensure Gen-AI relevancy.
+This stage filters reviews written prior to Gen-AI integration to ensure Gen-AI relevancy.
 Some apps included in our initial screening are not originally designed as Gen-AI tools but have integrated Gen-AI functionalities at a later stage.
-For instance, in the case of Microsoft Bing Search [^4] https://play.google.com/store/apps/details?id=com.microsoft.bing, Bing Chat is a recently introduced feature that extends the capabilities of the search engine by incorporating Gen-AI models (Iorliam_Ingio_2024).
+For instance, in the case of Microsoft Bing Search [^5] https://play.google.com/store/apps/details?id=com.microsoft.bing, Bing Chat is a recently introduced feature that extends the capabilities of the search engine by incorporating Gen-AI models (Iorliam and Ingio, 2024).
 Consequently, it is essential to exclude user reviews published before the integration of such features, as they reflect experiences with the app before the adoption of Gen-AI.
-While foundational work on AI models began as early as 2012, notable advancements in Gen-AI only started to gain momentum after 2018 (Sengar_Hasan_Kumar_Carroll_2024).
-The year 2018 also marked a rise in academic interest and publications related to Gen-AI (Dwivedi_Elluri_2024).
-However, the field witnessed a major breakthrough in 2022, particularly with the public release and widespread adoption of ChatGPT, which significantly accelerated the prominence and deployment of Gen-AI technologies across various platforms (Akhtar_2024).
-To maximize the relevance and accuracy of the user feedback analyzed in our study, we adopt 2022 as the cutoff year for including reviews.
+We outline our temporal filtering procedure below.
+  • Integration Date Identification.
+To approximate the Gen-AI integration date for each app, we adopt a conservative, multi-source identification strategy and record the earliest verifiable date across all available evidence.
+Specifically, we consult the following sources: (i) developer websites, blog posts, press releases, and public announcements; and (ii) historical release notes and version histories obtained from AppMagic [^6] https://appmagic.rocks/.
+When multiple sources report different dates, we retain the earliest date supported by documentary evidence.
+This approach prioritizes evidentiary certainty over speculative precision.
+  • Fallback Strategy.
+For Google Play Store apps, if no explicit integration date can be identified, or if the earliest documented evidence postdates October 2024, we conservatively assign October 2024 as the reference date.
+This date corresponds to the time at which we independently screened and verified the presence of Gen-AI functionality in the app.
+Therefore, the recorded date reflects the earliest verifiable confirmation of Gen-AI features rather than the exact deployment date.
+For App Store apps, we do not apply the October 2024 fallback, as App Store descriptions were not archived at the time of app selection.
+Consequently, if no verifiable integration date could be identified using release notes or developer sources, the app is excluded from temporal analysis.
+  • Summary of Identified Dates.
+For Google Play Store apps, we identify integration dates for 4 apps through public articles, for 125 apps through release notes, and 64 apps are assigned to October 2024 based on independent verification.
+For App Store apps, integration dates are identified for 2 apps via developer websites, 5 through public articles, and 63 through release notes.
+We exclude 10 App Store apps due to the absence of verifiable integration dates.
+  • Reliability and Coding Procedure.
+To assess reliability, we randomly sample 65 apps from our 193-app list (95% confidence level, 10% margin of error) for independent annotation.
+The first and second authors independently review release notes and supporting sources to assign integration dates.
+Inter-rater agreement is high (Cohen’s $k$ = 0.8524), indicating near-perfect agreement.
+Six disagreements are identified.
+These primarily arise from differing interpretations of whether certain features or naming conventions constituted evidence of generative functionality.
+For instance, references to “AI” in app names (e.g., “AI Song”) are not considered sufficient evidence unless accompanied by explicit indications of generative capabilities (e.g., “AI Image Generator”) or supporting descriptions.
+Additionally, features such as “AI effects” or “AI enhancement” are clarified as AI-assisted functionalities rather than Gen-AI because they modify content rather than generate novel content.
+Following the discussion, we refine the coding criteria accordingly and resolve all disagreements through consensus.
+The first author then applies the finalized criteria to label the remaining apps.
 Stage 2: Short review exclusion.
-Building upon the cleaned data from Stage 1, this stage employs a simple heuristic to further reduce noise by excluding reviews containing three words or less.
-Short reviews of three words or less do not contain context for meaningful topics (Assi_Hassan_Tian_Zou_2021; 10499727).
+Building upon the cleaned data from Stage 1, this stage employs a simple heuristic to further reduce noise by excluding reviews containing three words or fewer.
+Short reviews of three words or less do not contain context for meaningful topics (Assi et al., 2021; Arambepola et al., 2024).
 This step aims to eliminate overly brief and potentially non-informative feedback.
 Stage 3: LLM-assisted informative review filtration.
-Non-informative reviews bring noise to user reviews analysis and influence the performance of topic extraction (Chen_Lin_Hoi_Xiao_Zhang_2014).
+Non-informative reviews bring noise to user reviews analysis and influence the performance of topic extraction (Chen et al., 2014).
 To improve the quality of extracted topics, we exclude non-informative reviews.
@@
 While Chen et al.
-(Chen_Lin_Hoi_Xiao_Zhang_2014) focused on informativeness from a developer’s perspective, we expand this definition to consider the informational needs of multiple stakeholders, including users, platform owners, developers and researchers.
+(Chen et al., 2014) focused on informativeness from a developer’s perspective, we expand this definition to consider the informational needs of multiple stakeholders, including users, platform owners, developers and researchers.
 We carry out the LLM-assisted process in the following steps.
@@
     Prompt design.
-To filter non-informative reviews, we utilize Open AI’s (openai2023gpt) GPT-4o-mini model with zero temperature for reproducibility.
-We use gpt-4o-mini instead of the full gpt-4o model because it offers significantly faster response times while maintaining reasonable output quality for our task, which does not require complex reasoning (OpenAI_O4Mini_2024).
-    Prior research comparing explicit, implicit, and creative prompting (Giray_2023; Ho_Mayberry_Nguyen_Dhulipala_Pallipuram_2024; Deshmukh_Raut_Bhavsar_Gurav_Patil_2025) shows that explicit, well-structured prompts produce more factual and task-aligned outputs, while underspecified ones invite unpredictable reasoning.
-Consistent with these findings and with approaches such as LLM-Cure (Maram_LLM_Cure), we design an explicit, structured prompt that classifies each review as either informative or non-informative.
+To filter non-informative reviews, we utilize Open AI’s (OpenAI, 2023) GPT-4o-mini model with zero temperature for reproducibility.
+We use gpt-4o-mini instead of the full gpt-4o model because it offers significantly faster response times while maintaining reasonable output quality for our task, which does not require complex reasoning (OpenAI, 2024).
+Prior research comparing explicit, implicit, and creative prompting (Giray, 2023; Ho et al., 2024; Deshmukh et al., 2025) shows that explicit, well-structured prompts produce more factual and task-aligned outputs, while underspecified ones invite unpredictable reasoning.
+Consistent with these findings and with approaches such as LLM-Cure (Assi et al., 2025a), we design an explicit, structured prompt that classifies each review as either informative or non-informative.
 The prompt includes representative examples, a fixed binary label set, and a CSV-style output format to constrain the model’s responses.
@@
     Iterative prompt refinement.
-Prior studies highlight that prompt structure and iterative refinement (“dynamic prompting”) substantially influence LLM performance  (Marvin_Hellen_Jjingo_Nakatumba-Nabende_2024; He_Rungta_Koleczek_Sekhon_Wang_Hasan_2024).
+Prior studies highlight that prompt structure and iterative refinement (“dynamic prompting”) substantially influence LLM performance  (Marvin et al., 2024; He et al., 2024).
 In our setup, we adopt this principle by refining prompts through multiple pilot runs to improve label consistency and clarity.
@@
 Then, the first and second authors manually label the reviews in the samples as informative or non-informative.
-We calculate the inter-rater reliability using Kappa (cohen1960kappa) ($\kappa=0.91$), which indicates almost perfect agreement.
+We calculate the inter-rater reliability using Kappa (Cohen, 1960) ($\kappa=0.91$), which indicates almost perfect agreement.
     Next, we design and test different versions of the filtering prompt against the manually labeled data, evaluating the model’s filtering accuracy and iteratively refining the prompt.
@@
 We repeat this iterative process until we achieve a peak filtering accuracy of 91% on held-out validation samples.
-Figure 3 shows the final version of the filtering prompt, which includes ten examples (five informative and five non-informative) and yields optimal performance.
-    Figure 3. Template of the filtering prompt used to filter out non-informative reviews.
+The final version of the filtering prompt, which includes ten examples (five informative and five non-informative) and yields optimal performance, is provided in Appendix B.1.
   3. 3.
@@
 To assess the generalizability of the final filtering prompt and reduce the risk of overfitting to specific review instances, we manually evaluate its performance on a separate set of previously unseen reviews of the same categories and the same size.
-The model demonstrates strong generalization capabilities, achieving an average classification accuracy of 90%.
+The model demonstrates strong generalization capabilities, achieving an average filtering accuracy of 90%.
   4. 4.
@@
 Any unreturned reviews are consolidated into new batches and resent.
-This process reduces 920,824 reviews to 676,066 informative reviews, as shown in Table 2.
-Our data collection, spanning September and October 2024, resulted in an initial dataset of 2,778,643 user reviews across 186 unique Gen-AI apps.
-Following our multi-stage cleaning process, we obtain a dataset of 676,066 informative reviews across 173 Gen-AI apps, which forms the basis for our subsequent analysis.
+At the end of Stage 3, the dataset consists of informative reviews across 15 Google Play Store app categories.
+We then assess topic quality across categories and exclude eight categories with fewer than 1,000 reviews.
+Small sample sizes do not support the extraction of distinct topics, as reviews tend to overlap in content.
+Prior work shows that a fixed number of topics may be too high for some datasets, leading to over-clustering and insufficiently distinct topics (Greene et al., 2014).
+To validate this threshold, we perform topic extraction across all app categories and observe that categories with fewer than 1,000 reviews consistently produce redundant and poorly differentiated topics.
+For example, in the Books & Reference category (342 informative reviews), multiple topics capture overlapping concepts, such as “Content Generation Features” and “Quality of Generated Content,” or “User Experience” and “User Interface.” A similar pattern is observed in the Health & Fitness category (729 reviews), where topics remain insufficiently distinct.
+These redundancies introduce ambiguity during topic assignment, making it difficult to reliably determine which topic a review belongs to.
+Therefore, we adopt the 1,000-review threshold to ensure topic diversity, stability, and reliable labeling.
+Our data collection, spanning October 2024 to February 2026, results in an initial dataset of 4,183,347 Google Play Store user reviews across 172 unique Gen-AI apps and 536,779 App Store reviews from 70 apps.
+Following our multi-stage cleaning process, we obtain a dataset of 1,035,342 informative Google Play Store reviews across 171 Gen-AI apps, and 225,469 App Store reviews from 59 apps, which form the basis for our subsequent analysis.
 3.4. LLM-powered user review analysis
@@
 For RQ1, we design and evaluate our topic extraction and assignment prompts.
-We experiment with different data noise levels (i.e., different stages) and numbers of few-shot examples.
+We experiment with different levels of data noise (i.e., different stages) and numbers of few-shot examples.
 For RQ2, we extract the top 10 topics per app category and assign them to user reviews.
-We conduct manual thematic analysis for each Gen-AI topic to derive user insights from reviews.
-For RQ3, we analyze how Gen-AI topics evolve over time using clustering and manual review.
+We conduct manual thematic analysis for each Gen-AI topic to identify challenges and opportunities derived from user review insights.
+For RQ3, we analyze how Gen-AI topics evolve over time using temporal clustering and manual review.
 In the following sections, we describe the experiment design and the findings for each RQ.
+3.5. Adaptability of SARA Framework
+In this study, SARA is empirically evaluated only on Gen-AI mobile apps.
+However, SARA is designed as a modular analysis framework in which the overall analytical framework remains fixed, while domain-specific configurations can be adapted.
+In particular, SARA can be applied to analyze user reviews of multiple mobile apps across different app categories and app stores by customizing domain-specific components.
+Because the framework operates on user-generated review data rather than Gen-AI-specific content, it can be adapted to analyze user perceptions of a wide range of mobile apps.
+In the first stage, app selection is domain-agnostic: researchers may define and select target apps based on any domain-specific criteria or definitions relevant to their study.
+The second stage, review acquisition, follows the same procedure regardless of domain, as user reviews are collected from app store platforms using consistent metadata and retrieval mechanisms.
+The review refinement stage is largely reusable across domains.
+Basic preprocessing steps (e.g., removal of white space, emojis, and short reviews) remain unchanged, and our findings in RQ1 show that LLMs are effective at filtering non-informative reviews using a general definition of informative reviews that is not tied to Gen-AI-specific functionality.
+As such, the non-informative review-filtering prompt can be applied without modification to other domains.
+In the analysis stage, SARA employs topic extraction followed by topic assignment.
+While the overall prompt structure for topic extraction is reusable, the examples and topic labels must be adapted to reflect domain-specific concepts and functionalities.
+This customization is expected and enables the framework to capture contextually relevant themes.
+Similarly, topic assignment relies on a reusable prompt template, with domain-specific examples supplied by the researcher when applying the framework to a new domain.
+Together, these design choices allow SARA to be adapted to multiple mobile apps across any app category and across different app stores by modifying domain-specific inputs and topic definitions, while preserving the framework’s core analytical stages and processing logic.
 4. RQ1: How accurately can LLMs identify topics in user reviews of Gen-AI apps?
 Effective topic extraction and assignment of user reviews are essential for understanding user needs and improving app features for Gen-AI apps.
-Prior work has shown that prompt engineering techniques can help tailor LLMs to specific tasks (Marvin_Hellen_Jjingo_Nakatumba-Nabende_2024; 10.1145/3736408).
-Applying LLMs to topic extraction and assignment requires well-designed prompts to achieve high accuracy.
-In our setup, we adopt this principle by refining prompts through multiple pilot runs on statistically significant samples to improve label consistency and clarity.
-In this RQ, we investigate two key aspects of prompt design: 1) the extent to which cleaning the input data prior to prompting improves accuracy, and 2) the effect of varying the number of few-shot examples in the prompt.
-The goal is to identify the optimal combination of preprocessing and few-shot-based prompting that enables more accurate and reliable topic extraction and assignment from user reviews.
+We identify the optimal combination of preprocessing and few-shot prompting to enable more accurate and reliable topic extraction and assignment from user reviews.
 4.1. Experiment Setup
-To design an effective topic extraction prompt, we experiment with varying levels of data cleaning and different numbers of few-shot examples.
-As described in section 3, we design a four-stage data cleaning framework.
-Stage 0 includes basic textual cleaning.
-Stage 1 excludes reviews before the launch of ChatGPT in 2022 (Akhtar_2024).
-Stage 2 excludes short reviews with three words or less.
-Stage 3 excludes non-informative reviews using an LLM-based filtering approach.
-We conduct experiments on Stage 1, Stage 2, and Stage 3, but not on Stage 0.
-We exclude Stage 0 because while it performs essential preprocessing, it does not remove irrelevant data as done in Stage 1.
+To design an effective topic extraction prompt, we conduct experiments on Stage 1, Stage 2, and Stage 3, but not on Stage 0.
+We exclude Stage 0 because, although it performs essential preprocessing, it does not remove irrelevant data as in Stage 1.
 We approach this problem in four steps: 1) we generate random samples of user reviews for topic extraction and assignment drawing from each stage of data cleaning, 2) we apply three prompting strategies with varying numbers of few-shot examples (i.e., 0-shot, 3-shot, and 5-shot) to extract topics from these samples, 3) assign the extracted topics to a new set of sample reviews using the same prompts, and (4) we evaluate the accuracy of the topic assignment to determine the optimal combination of data cleaning and few-shot examples.
@@
 This sample is used for topic assignment, where manual evaluation is feasible and interpretable.
-Figure 5.
-Template of the topic extraction prompts for (E1: 0-shot, E2: 3-shot, and E3: 5-shot) used to extract the top topics from a large sample of stage $x$ and app category $y$.
-Insert category name is where we insert the name of the app category $y$.
-Insert sample of reviews is where we list our large sample of reviews separated by a new line escape.
 Topic extraction.
-Because prior work reported mixed effects of few-shot examples (Wang_Yang_Wei_2023; Prakash_Wang_Hoang_Hee_Lee_2023), we design three experiments, i.e., E1: 0 shot, E2: 3 shot and E3: 5 shot, with three prompt configurations differing in the number of few-shot examples, i.e., 0-shot, 3-shot and 5-shot.
-These values are chosen based on prior research and practical considerations in prompt-based learning  (Dos_Santos_Oliveira_De_Jesus_Aljedaani_Eler_2023; Maram_LLM_Cure; Pham_Hoyle_Sun_Resnik_Iyyer_2024).
+Because prior work reported mixed effects of few-shot examples (Wang et al., 2023; Prakash et al., 2023), we design three experiments, i.e., E1: 0 shot, E2: 3 shot and E3: 5 shot, with three prompt configurations differing in the number of few-shot examples, i.e., 0-shot, 3-shot and 5-shot.
+These values are chosen based on prior research and practical considerations in prompt-based learning  (Dos Santos et al., 2023; Assi et al., 2025a; Pham et al., 2024).
 In E1: 0 shot, the prompt contains only the instructions.
@@
 This results in a set of extracted top topics, denoted as $T(x,y,n)$, where $n\in\{0,3,5\}$ indicates the number of few-shot examples used in the prompt.
-Figure 5 shows the three prompt configuration templates.
-To mitigate hallucination during topic extraction, we anchor the prompt in actual user reviews, instruct the model to output only ten distinct high-level topics that describe app features, functionalities, and utility, and provide few-shot examples that illustrate the expected level of abstraction and discourage free-form or overly creative topic invention.
+The final version of the topic extraction prompt and the few-shot examples can be found in Appendix B.2.
+To mitigate hallucinations during topic extraction, we anchor the prompt in actual user reviews, instruct the model to output only ten distinct high-level topics that describe app features, functionalities, and utility, and provide few-shot examples that illustrate the expected level of abstraction and discourage free-form or overly creative topic invention.
 To further adapt the LLM to the Gen-AI domain, the few-shot examples are drawn from real reviews of Gen-AI apps, exposing the model to characteristic terminology and contexts (e.g., discussions of AI Performance, Technical Difficulties, and Content Quality).
-Figure 6.
-Template of the P5: topic assignment prompt used to assign topics to a list of reviews.
-We insert the set of top topics $T(x,y,n)$, five few-shot examples, and the small sample $S_{\text{small}}(x,y)$.
 Topic assignment.
 To evaluate the effectiveness of the extracted topic sets $T(x,y,n)$ produced by each topic extraction experiment with $n\in\{0,3,5\}$, we perform a follow-up topic assignment task using LLM.
-Specifically, for each experiment $n$, we design a dedicated assignment prompt, illustrated in Figure 6, that takes as input: 1) the category-specific set of top topics $T(x,y,n)$ extracted for that app category in that experiment, 2) five few-shot example reviews manually selected and labeled from the corresponding large sample $S_{\text{large}}(x,y)$, and 3) the target reviews from the small sample $S_{\text{small}}(x,y)$ to which topics are assigned.
-We manually select and assign topics to five few-shot examples to improve assignment accuracy as done in prior work (Maram_LLM_Cure; Deshmukh_Raut_Bhavsar_Gurav_Patil_2025).
+Specifically, for each experiment $n$, we design a dedicated assignment prompt that takes as input: 1) the category-specific set of top topics $T(x,y,n)$ extracted for that app category in that experiment, 2) five few-shot example reviews manually selected and labeled from the corresponding large sample $S_{\text{large}}(x,y)$, and 3) the target reviews from the small sample $S_{\text{small}}(x,y)$ to which topics are assigned.
+We manually select and assign topics to five few-shot examples to improve assignment accuracy as done in prior work (Assi et al., 2025a; Deshmukh et al., 2025).
+The final topic assignment prompt is provided in Appendix B.3.
 To mitigate hallucination during topic assignment, we constrain the model’s output to a predefined set of deterministic labels, i.e., the extracted topics $T(x,y,n)$.
@@
 To assess the accuracy of the LLM-assigned topics, we conduct a manual annotation of the assignments.
-The first and second authors manually and independently label the correctness of LLM assignment.
+The first and second authors manually and independently label the correctness of the LLM assignment.
 An assignment is considered correct if the LLM assigned topic accurately reflects the content of the review.
-The Cohen’s Kappa agreement score (cohen1960kappa) is computed, yielding a score of 0.56, indicative of moderate agreement.
+The Cohen’s Kappa agreement score (Cohen, 1960) is computed, yielding a score of 0.56, indicative of moderate agreement.
 The two authors discuss and resolve the initial disagreements to align on labeling criteria.
+We analyze cases of annotator disagreement to better understand ambiguity in topic classification.
+Disagreements primarily occur in reviews with limited context or broad, non-informative statements (e.g., the “brief expressions of satisfaction” topic identified in earlier stages), where the relevance of a specific topic is unclear.
+In these cases, annotators differ in their strictness: one may accept the assigned topic as valid, while another may judge the review as too vague or incoherent to support that assignment.
+Upon discussion, we label such assignments as incorrect, as the reviews do not provide sufficient or interpretable evidence for the assigned topic, and therefore should have been marked as “other” by the LLM.
+These disagreements highlight the challenges of topic classification in earlier stages, where noisy and non-informative reviews can lead to the extraction of generic topics.
+As a result, assigning topics becomes difficult due to overlapping themes and the absence of clear semantic signals in the underlying reviews.
 In total, 4,353 LLM topic assignments are manually evaluated for correctness for five categories, three stages, and three experiments.
 Table 3 shows the accuracy for each cleaning stage, app category and shot examples.
-To determine whether the observed accuracy differences are statistically significant, we perform statistical tests.
-Specifically, we apply the non-parametric test, Kruskal–Wallis H test (Chan_Walmsley_1997), to evaluate the effect of few-shot prompt configurations on overall accuracy.
-Additionally, post-hoc pairwise comparisons are conducted using the Mann–Whitney U test (Mann_Whitney_1947) with Bonferroni correction to adjust for multiple testing (3b6386c9-eeae-33a7-8c59-c0604aa87bef).
-We also use one-way ANOVA (Judd_McClelland_Ryan_2017) to test for significant differences across cleaning stages and few-shot settings.
+We supplement this with Table 4, which reports accuracy on informative reviews across cleaning stages, app categories, and shot settings.
+This separation enables a fair comparison, as non-informative reviews are easier to classify.
 Table 3.
@@
 E3: 5 shot 93% 81% 94% 90% 92% 90% 90% 84% 92% 88% 93% 89% 95% 86% 86% 95% 91% 91%
-4.2. RQ1 Findings
-Observation 1.1: Providing more examples significantly improves classification and assignment accuracy.
-The average accuracy increases across experiments for each cleaning stage, with E3: 5-shot consistently achieving the highest accuracy in all three stages, as shown in Table 3.
-Notably, E3: 5-shot is the only setting to achieve average accuracies above 80% across all samples.
-The Kruskal–Wallis test indicates that the number of few-shot examples has a statistically significant effect on overall accuracy ($H=7.778$, $p=0.0205$).
-Post hoc Mann–Whitney U tests with Bonferroni correction show that the accuracy improvement from E1: 0 shot to E3: 5 shot is statistically significant ($p=0.0181$), whereas differences between E1: 0 shot and E2: 3 shot ($p=1.000$) and E2: 3 shot and E3: 5 shot ($p=0.2519$) are not.
-These results suggest that increasing the number of few-shot examples, specifically 5-shot, significantly enhances overall model accuracy.
-Observation 1.2: The proportion of informative reviews varies by cleaning stage, which impacts overall accuracy.
-As shown in Table 3, accuracy does not follow a consistent trend across the three cleaning stages (i.e., Stage 1, Stage 2, and Stage 3).
-For instance, the average accuracy for E1: 0 shot and E2: 3 shot increase from Stage 1 to Stage 2, but decrease from Stage 2 to Stage 3.
-In contrast, E3: 5 shot shows a slight decrease from Stage 1 to Stage 2, followed by a slight increase in Stage 3.
-This inconsistency may be explained by differences in the proportion of informative reviews across stages: 43% in Stage 1, 66% in Stage 2, and 88% in Stage 3.
-Since assigning topics to non-informative reviews is generally easier and less dependent on nuanced understanding, variations in their prevalence can skew overall accuracy.
-As a result, overall accuracy may not accurately reflect the model’s effectiveness in handling more meaningful and content-rich reviews across different cleaning stages.
 Table 4.
@@
 E3: 5 shot 87% 85% 94% 82% 76% 85% 80% 79% 90% 82% 89% 84% 94% 92% 90% 95% 97% 94%
-Observation 1.3: Stage 3 consistently yields the highest accuracy on informative reviews across all app categories and experiments.
-Given the inconsistencies observed in overall accuracy explained in observation 1.2, we isolate informative reviews to better evaluate model performance.
+To determine whether the observed accuracy differences are statistically significant, we perform statistical tests.
+Specifically, we apply the non-parametric test, Kruskal-Wallis H test (Chan and Walmsley, 1997), to evaluate the effect of few-shot prompt configurations on overall accuracy.
+Additionally, post-hoc pairwise comparisons are conducted using the Mann-Whitney U test (Mann and Whitney, 1947) with Bonferroni correction to adjust for multiple testing (Cabin and Mitchell, 2000).
+We also use one-way ANOVA (Judd et al., 2017) to test for significant differences across cleaning stages and few-shot settings.
+4.2. RQ1 Findings
+Finding 1.1: Providing more examples significantly improves classification and assignment accuracy.
+The average accuracy increases across experiments for each cleaning stage, with E3: 5-shot consistently achieving the highest accuracy in all three stages, as shown in Table 3.
+Notably, E3: 5-shot is the only setting to achieve average accuracies above 80% across all samples.
+The Kruskal-Wallis test indicates that the number of few-shot examples has a statistically significant effect on overall accuracy ($H=7.778$, $p=0.0205$).
+Post hoc Mann-Whitney U tests with Bonferroni correction show that the accuracy improvement from E1: 0 shot to E3: 5 shot is statistically significant ($p=0.0181$), whereas differences between E1: 0 shot and E2: 3 shot ($p=1.000$) and E2: 3 shot and E3: 5 shot ($p=0.2519$) are not.
+These results suggest that increasing the number of few-shot examples, specifically 5-shot, significantly enhances overall model accuracy.
+Finding 1.2: The proportion of informative reviews varies by cleaning stage, which impacts overall accuracy.
+As shown in Table 3, accuracy does not follow a consistent trend across the three cleaning stages (i.e., Stage 1, Stage 2, and Stage 3).
+For instance, the average accuracy for E1: 0 shot and E2: 3 shot increase from Stage 1 to Stage 2, but decrease from Stage 2 to Stage 3.
+In contrast, E3: 5 shot shows a slight decrease from Stage 1 to Stage 2, followed by a slight increase in Stage 3.
+This inconsistency may be explained by differences in the proportion of informative reviews across stages: 43% in Stage 1, 66% in Stage 2, and 88% in Stage 3.
+Since assigning topics to non-informative reviews is generally easier and less dependent on nuanced understanding, variations in their prevalence can skew overall accuracy.
+As a result, overall accuracy may not accurately reflect the model’s effectiveness in handling more meaningful and content-rich reviews across different cleaning stages.
+Finding 1.3: Stage 3 consistently yields the highest accuracy on informative reviews across all app categories and experiments.
+Given the inconsistencies observed in overall accuracy explained in finding 1.2, we isolate informative reviews to better evaluate model performance.
 Table 4 shows that, across all experiments and app categories, average accuracy on informative reviews is highest in Stage 3.
@@
 The ANOVA for few-shot experiments does not reveal significant differences ($F(2,42)=2.25$, $p=0.118$).
-Observation 1.4: The best performing setup combines filtering non-informative reviews, i.e., Stage 3 and more few shot examples, i.e., E3: 5-shot.
+Finding 1.4: The best performing setup combines filtering non-informative reviews, i.e., Stage 3 and more few shot examples, i.e., E3: 5-shot.
 Our findings suggest that using LLMs to filter the non-informative reviews and increasing the number of few-shot examples to five improves the accuracy of the model.
-Based on these findings, we adopt the Stage 3 – E3: 5-shot configuration for topic classification and assignment in RQ2, as it yielded the highest overall accuracy of 91% and the highest informative accuracy of 94%.
-Observation 1.5: Under the best-performing configuration (E3: 5-shot), only 9% of reviews were misclassified.
+Based on these findings, we adopt the Stage 3 - E3: 5-shot configuration for topic classification and assignment in RQ2, as it yielded the highest overall accuracy of 91% and the highest informative accuracy of 94%.
+Finding 1.5: Under the best-performing configuration (E3: 5-shot), only 9% of reviews were misclassified.
 Among these, 2% of all reviews involved incorrect topic assignments, 3% were missed topics (valid reviews marked as “Other”), and 4% were over-assignments (non-topical reviews given a label).
@@
 Accordingly, we adopt this configuration for topic extraction in RQ2.
-5. RQ2: What are the most prominent topics discussed in user reviews and in developer responses of Gen-AI apps?
-Having validated the accuracy of our LLM-based topic extraction and assignment framework in RQ1, we now apply it to the full dataset to examine what users discuss in their reviews of Gen-AI apps.
-Understanding the issues that users care about is essential for improving the design and evaluation of Gen-AI apps.
-In RQ2, we uncover the top 10 topics that users discuss in reviews of Gen-AI apps across different app categories.
-We then analyze topic frequency and associated review ratings to interpret common user behaviours and needs, supporting the identification of actionable insights that inform the design of more adaptive and user-centered Gen-AI experiences.
+5. RQ2: What are the most prominent topics discussed in user reviews?
+We identify the top 10 topics discussed in user reviews across app categories and analyze their frequency and associated ratings.
+This analysis uncovers common user concerns, expectations, and usage patterns, from which we derive key opportunities, challenges, and actionable insights for designing more effective, user-centered Gen-AI systems.
 5.1. Experiment Setup
-In RQ1, we experiment with varying levels of data cleaning and numbers of few-shot examples.
-We find that filtering out non-informative reviews using an LLM and providing five-shot examples improves topic extraction and assignment accuracy.
-In RQ2, we apply the optimal topic extraction prompt identified in RQ1 to the remaining app categories and assign topics to all reviews in Stage 3.
-In this section, we provide a detailed description of the implementation of RQ2.
+We apply RQ1’s optimal configuration to identify and analyze topics from user reviews.
+This section describes the topic extraction process, the classification of extracted topics, and the quantitative and qualitative analyses used to interpret user feedback.
+We implement our framework to both Google Play Store and App Store reviews and discuss platform comparison in Section 7.
+In this section, our analysis is focused on the Google Play Store dataset because it is larger and more substantive.
 Topic extraction and assignment.
-The results of RQ1 indicate that the optimal dataset for analysis is Stage 3 data, which underwent four stages of cleaning: basic cleaning, temporal filtering, short review exclusion, and LLM-assisted informative review filtration.
-Our dataset spans 11 app categories, totaling 676,066 user reviews as shown in Table 2.
-As in RQ1, we sample a statistically significant set of informative reviews (large sample) per app category, using a 95% confidence level and 2% margin of error.
-The resulting sample sizes range from 761 to 2,368 reviews, with an average of 2,031 per category.
-For each app category $y$, we extract the top 10 topics from the large sample using the validated 5-shot topic extraction prompt shown in Figure 5.
-Next, we use the topic assignment prompt shown in Figure 6 to assign one of the extracted topics to each review in the dataset.
-For each app category $y$, the corresponding top 10 topics are used as input to label reviews with a single topic.
+Following RQ1, we draw a statistically significant large sample of informative reviews from stage 3 (2,291 average reviews per app categories) to extract the top 10 topics from each app category.
+These topics are then used to assign a single topic label to each review in the full dataset.
 Classification and categorization.
-As our primary interest lies in understanding user feedback related to Gen-AI functionality, we identify and analyze reviews that discuss such features.
-To do so, we manually examine all extracted topics and label them as either Gen-AI topics, which reference Gen-AI capabilities, or non-Gen-AI topics, which pertain to general app features outside the scope of Gen-AI.
-All reviews originate from Gen-AI apps; the distinction here is between review types, not between Gen-AI and non-Gen-AI apps.
-We further group semantically related topics into high-level topic categories based on shared themes or functionality, facilitating higher-level interpretation of user concerns.
-In line with prior studies (Maram_LLM_Cure; Pagano_Maalej_2013; Chen_Lin_Hoi_Xiao_Zhang_2014), we also classify reviews with 1 or 2-star ratings as negative, 3-star rating as neutral, and 4 or 5-star rating as positive.
+Our primary interest is user feedback related to Gen-AI functionality.
+We therefore manually examine the extracted topics and label them as either Gen-AI topics, which reference Gen-AI capabilities, or non-Gen-AI topics, which refer to general app functionality.
+All reviews originate from Gen-AI apps; the distinction therefore concerns the type of feedback rather than the app type.
+To facilitate higher-level interpretation, we group semantically related topics into broader topic categories.
+In line with prior studies (Assi et al., 2025a; Pagano and Maalej, 2013; Chen et al., 2014), we also categorize reviews by star ratings: 1-2 stars as negative, 3 stars as neutral, and 4-5 stars as positive.
 Analysis.
-We perform large-scale quantitative analysis on all automatically labeled reviews across both app categories and topic categories, followed by qualitative analysis on a representative sample.
-Specifically, for the quantitative component, we compute the following metrics:
+We perform large-scale quantitative analysis on all labeled reviews across both app categories (e.g., Productivity and Photography) and topic categories (e.g., AI Performance and Creative Potential).
+Specifically, we compute the following metrics:
   1. (1)
-    Average Gen-AI Rating (AGR): the average star rating of reviews labeled with Gen-AI topics.
+    Average Gen-AI Rating (AGR): the average rating of reviews discussing Gen-AI topics.
   2. (2)
-    Average Non-Gen-AI Rating (ANR): the average rating of reviews labeled with non-Gen-AI topics.
+    Average Non-Gen-AI Rating (ANR): the average rating of reviews discussing non-Gen-AI topics.
   3. (3)
@@
     Non-Gen-AI Review Percentage (NRP): the proportion of non-Gen-AI reviews relative to the total number of reviews.
-We conduct the Mann–Whitney U test to examine whether reviews labeled as Gen-AI-related receive significantly higher ratings than those that are not.
-For developer-response analysis, we compute reply ratios, response delays, and response-type distributions for both Gen-AI and non–Gen-AI reviews, and we compare per-app reply behavior using a Wilcoxon signed-rank test (wilcoxon1992).
-We complement our quantitative analyses with qualitative examination by manually coding statistically significant samples (95% confidence level, 10% margin of error) of both user reviews and developer responses.
-For user reviews we draw samples from every Gen-AI topic category resulting in a total of 901 reviews.
-We manually code and analyze all sampled reviews to provide richer, contextualized insights that may surface actionable implications for both researchers and developers.
-For developer responses, we perform two qualitative analyses.
-First, we code a representative sample of 96 replies drawn across all app categories to characterize the overall distribution of developer responses.
-Second, we manually code an additional 281 replies from three focal categories (Games, Education, and Entertainment) to capture deeper, category-specific response practices.
-Together, these qualitative analyses provide contextual grounding for the quantitative findings and offer a more complete understanding of both user concerns and developer engagement within Gen-AI apps.
+To examine whether reviews discussing Gen-AI functionality receive higher ratings than other reviews, we compute AGR and ANR for each app by aggregating review scores at the app level.
+We apply the Wilcoxon signed-rank test (Wilcoxon, 1992) to compare the two ratings.
+To complement the quantitative results, we conduct qualitative coding of statistically significant samples (95% confidence level, 10% margin of error).
+We sample reviews from each Gen-AI topic category, resulting in 762 reviews that are manually coded to identify recurring user concerns that derive opportunities, challenges, and actionable insights.
 5.2. RQ2 Findings
-Observation 2.1: Gen-AI reviews consistently receive higher user ratings (AGR) than non-Gen-AI reviews (ANR) across all app categories of Gen-AI apps.
-As shown in Table 5, the average rating of reviews discussing Gen-AI functionality (i.e., AI Performance and Utility & Use Cases) exceeds the average rating of non-Gen-AI reviews (i.e., Technical Difficulties and Monetization Methods & Structure) in every app category of Gen-AI apps.
-This indicates that, when users explicitly comment on Gen-AI features, they tend to rate their experiences more positively than when they comment on other aspects of the same apps.
-To assess the statistical significance of this difference, we apply the Mann–Whitney U test.
-The results reveal a significant difference in the rating distributions between the two groups (U=65,113,513,850U=65{,}113{,}513{,}850, $p<0.0001$).
-Gen-AI reviews tend to receive higher ratings, with a median rating of 5.0 and a mean of 4.10, compared to a median of 4.0 and a mean of 3.32 for non-Gen-AI reviews.
-These findings suggest that Gen-AI functionality is a major contributor to positive user experiences across a wide range of app categories.
-This pattern is especially pronounced in categories such as Education and Books & Reference, which show some of the largest gaps between AGR and ANR.
-In Books & Reference, Gen-AI topics such as AI Performance and Human AI Interaction & Emotional Connection receive high ratings, while non-Gen-AI topics like Technical Difficulties and Customer Support are rated much lower due to crashes, generation failures, and unresolved support issues.
-Similarly, in Education, users highly rate Gen-AI topics such as Content Quality and Utility & Use Cases, but express dissatisfaction with Monetization Methods & Structure and Technical Difficulties.
-Together, these results indicate that strong Gen-AI features can substantially boost user satisfaction, but poor reliability, monetization, or support can still pull overall ratings down.
-Developers should therefore treat Gen-AI capabilities as amplifiers of perceived value, not as substitutes for robust app quality, stability, and fair pricing.
+Finding 2.1: Gen-AI reviews consistently receive higher ratings than non-Gen-AI reviews across app categories, suggesting that negative feedback often relates to app design issues rather than the Gen-AI functionality itself.
+Table 5 shows that reviews discussing Gen-AI functionality receive higher ratings than those discussing other app aspects, with an average of 4.2 compared to 2.8.
+A Wilcoxon signed-rank test confirms that AGR is significantly higher than ANR across apps ($p<0.001$), with significant differences observed in 7 of 9 categories.
+This pattern is particularly pronounced in categories such as Art & Design and Music & Audio, where Gen-AI topics related to Creative Potential receive very high ratings (4.8-4.9), while non-Gen-AI topics such as Monetization Methods & Structure receive substantially lower ratings (1.8).
 Table 5.
@@
 App Category AGR ANR GRC GRP # Gen-AI Topics
-Productivity 4.2 3.5 84,122 50% 6
-Entertainment 4.0 3.2 56,672 48% 6
-Education 4.6 3.3 31,465 75% 6
-Art & Design 3.4 3.2 26,585 21% 3
-Photography 3.9 3.3 15,567 12% 2
-Tools 4.0 3.4 7,782 22% 3
-Health & Fitness 4.8 4.0 6,896 88% 7
-Video Players & Editors 4.6 3.9 5,179 14% 2
-Games 3.6 2.7 1,788 35% 5
-Music & Audio 3.0 2.6 1,679 18% 4
-Books & Reference 4.1 2.3 428 39% 4
-Average 4.0 3.2 21,651 38% 4
-Observation 2.2: Gen-AI functionality is a prominent focus in user feedback, comprising an average of 38% of reviews across app categories and reaching as high as 88% for the Health & Fitness category.
-As shown in Table 5, the prevalence of Gen-AI–focused reviews varies by app category, reflecting the differing roles Gen-AI plays across app domains.
-Categories such as Health & Fitness, Education, and Productivity contain the highest proportions of Gen-AI reviews, with 88%, 75%, and 50%, respectively, indicating that users in these domains directly engage with AI features to achieve personal or task-oriented goals such as self-improvement, learning, or workflow assistance.
+Productivity 4.3 3.0 257,434 64% 5
+Photography 4.5 2.8 66,066 39% 3
+Entertainment 3.6 2.8 61,803 41% 5
+Art & Design 4.3 2.3 54,227 43% 3
+Education 4.5 4.0 46,942 72% 6
+Games 3.9 2.4 15,695 62% 6
+Tools 4.3 3.0 14,923 41% 5
+Video Players & Editors 4.5 3.2 12,741 32% 3
+Music & Audio 4.1 2.1 6,992 39% 5
+Average 4.2 2.8 59,647 48% 5
+These results suggest that lower ratings are often driven by non-Gen-AI concerns, particularly monetization and technical issues, rather than dissatisfaction with Gen-AI functionality itself.
+For developers, these results suggest that strong Gen-AI capabilities alone are insufficient, as poor app design, reliability issues, and monetization practices can override their benefits.
+Finding 2.2: The percentage of reviews discussing Gen-AI varies across app categories, reflecting whether Gen-AI is a core or supplementary feature.
+Table 5 shows variation in the proportion of Gen-AI-related reviews across categories.
+Categories such as Education (72%) and Productivity (64%) exhibit the highest percentages, indicating that users primarily evaluate these apps through their AI capabilities.
+In these domains, discussions frequently center on AI performance and practical use cases, such as studying, writing, and content generation, highlighting strong user interest in the applicability of Gen-AI features.
+In contrast, categories such as Video Players & Editors, Music & Audio, and Photography show much lower proportions of Gen-AI reviews, where AI functionality typically serves as a supplementary feature within broader tools.
+As a result, user feedback is more focused on non-Gen-AI concerns such as technical issues, monetization, and usability.
+Figure 4 further shows variation within categories, indicating that apps differ in how they position Gen-AI functionality.
+For developers, these results suggest that when Gen-AI is a core feature, users focus on its utility and performance, whereas when it is supplementary, overall app design and reliability play a larger role in shaping user perceptions.
 Figure 4. Box plot showing the distribution of the percentage of reviews classified as Gen-AI at the app level, grouped by app category (Full dataset considered).
-In contrast, categories such as Photography and Video Players & Editors exhibit lower percentages of Gen-AI reviews, where 58% of feedback still focuses on non-Gen-AI topics such as usability, technical issues, and subscription or pricing models.
-This suggests that in creative or media-focused apps, users remain more concerned with traditional performance and value-for-money factors than with AI-driven functionality.
-Figure 7 illustrates the distribution of Gen-AI review percentages across individual apps within various categories, highlighting how consistently users discuss Gen-AI features within each domain.
-Categories such as Entertainment and Health & Fitness exhibit the widest spreads, suggesting substantial variation among apps; some are highly AI-centric, while others receive minimal user attention to Gen-AI features.
-In contrast, categories like Video Players & Editors and Photography show narrow spreads concentrated at low percentages, indicating that across these domains, Gen-AI features consistently play an assistive or complementary role rather than being the central focus of user experience.
-Together, these patterns suggest that while certain app categories demonstrate uneven user emphasis on Gen-AI, others achieve relatively uniform user engagement of Gen-AI features.
-Observation 2.3: Users discuss a wide range of Gen-AI topics, revealing where developers excel and where improvement is needed.
-Table 6 and Table 7 present the top Gen-AI and non-Gen-AI topics, respectively.
-Developers should pay particular attention to topics that appear most frequently across apps because they represent the aspects users discuss and evaluate most.
-Excelling in these areas is critical for maintaining competitiveness and user satisfaction.
-The Gen-AI topics with the highest numbers of reviews are AI Performance, Utility & Use Cases, and Content Quality.
-The topics most commonly mentioned across different apps are AI Performance, Content Quality, and Content Policy & Censorship.
-Similarly, topics with high average ratings (e.g., Human AI Interaction & Emotional Connection and Utility & use cases) serve as benchmarks for best user experiences that other apps can emulate.
-Conversely, lower-rated or frequently criticized topics (e.g., Content Policy & Censorship and Content Quality) highlight recurring weaknesses across Gen-AI apps.
-Developers should treat these as priority areas for refinement to stay ahead of the competition and sustain user trust.
-We further examine each of the Gen-AI categories in the subsections below.
-Table 6.
-Summary of Gen-AI topic categories, including the Gen-AI Reviews Count (GRC), the average Gen-AI rating (AGR), the number of apps, and the number of app categories in which each topic category appears (Full dataset considered).
-Gen-AI Topic Category GRC AGR # Apps # App Categories
-AI Performance 61,071 4.0 127 10
-Utility & Use Cases 56,505 4.8 45 3
-Content Quality 44,594 3.6 131 8
-Content Policy & Censorship 27,521 2.9 76 5
-Human AI Interaction & Emotional Connection 25,142 4.8 13 4
-Features & Functionality 16,275 4.0 58 4
-Customer Support & Community 3,361 4.7 12 3
-Comparison to Other Apps 2,315 3.9 13 3
-Personalization & Customization 1,283 4.0 30 3
-Accessibility & Inclusivity 96 4.1 4 1
-Table 7.
-Summary of non-Gen-AI topic categories, including the non-Gen-AI Review Count (NRC), the average non-Gen-AI rating (ANR), the number of apps, and the number of app categories in which each topic category appears (Full dataset considered).
-Non-Gen-AI Topic Category NRC ANR # Apps # App Categories
-Technical Difficulties 99,621 2.2 157 10
-Monetization Methods & Structure 85,756 2.3 159 11
-User Interface & Experience 80,827 4.5 156 11
-Utility & Use Cases 21,464 4.8 34 1
-Updates & Evolution 21,100 2.8 92 7
-Personalization & Customization 6,860 4.8 18 1
-Customer Support & Community 6,848 1.8 85 6
-Limitations of Usage 5,971 2.5 45 4
-Features & Functionality 1,810 3.0 11 1
-Comparison to Other Apps 1,493 3.4 17 1
-Accessibility & Inclusivity 401 2.6 5 2
-Other (unassigned) 105,752 4.2 165 11
-1) AI Performance.
-AI Performance refers to how well the AI model performs in terms of 1) accuracy (27%), 2) speed (22%), 3) understanding (15%), 4) memory (5%), 5) reliability (5%), and 6) creativity & randomness (3%).
-AI accuracy describes the precision of the AI response to the user’s prompt.
-AI speed represents how quickly the AI responds to a user’s prompt.
-AI understanding indicates the AI’s ability to comprehend the meaning, intent, and context behind the user’s prompt.
-AI memory signifies how well the AI remembers prior conversations and links them to the current one.
-AI can have both short-term memory, which spans a single conversation, and long-term memory, which spans multiple conversations.
-Reliability refers to the consistency and dependability of the AI system in performing its intended tasks without crashing, freezing, timing out, or failing to generate output.
-Creativity captures the degree of originality, uniqueness, or unpredictability in the AI’s responses.
-Figure 8. Examples of reviews discussing AI Performance
-Insights.
-We find that 79% of the users who discussed AI Performance provided a positive review with a rating of 4 or 5.
-Reviews 1 and 2 in Figure 8 show examples of positive reviews.
-Despite this overall satisfaction, 16% of reviews mentioned limitations in AI Performance, with the most frequently cited issue being AI understanding (5%), followed by inconsistency in accuracy (4%), reliability (3%), speed (2%), and AI memory (1%).
-A recurring theme in the less favourable reviews is poor AI understanding, which leads to irrelevant or partially correct responses.
-Some users show awareness of prompt engineering and attempt to guide the AI with examples, yet still encounter issues.
-For instance, review 3 in Figure 8 illustrates a user struggling to communicate effectively with the AI, despite providing clear examples.
-Similarly, review 4 in Figure 8 highlights inconsistent performance for complex requests.
-While the AI excels at handling simple tasks, it often struggles with more complex ones.
-Although memory-related concerns (e.g., forgetting prior context or repeating content) are occasionally mentioned as shown in review 5 in Figure 8, they tend to be rated more leniently, indicating that users may be more forgiving of these limitations.
-2) Content Quality. Content Quality refers to users’ assessment of AI-generated content in terms of usefulness, coherence, creativity, engagement, structure and presentation.
-Figure 9. Examples of reviews discussing Content Quality
-Insights.
-In our sample, 67% of reviews express satisfaction, praising the creative and polished outputs such as producing “radio-ready” songs and appending useful resources to textual answers as illustrated in reviews 1 and 2 in Figure 9 respectively.
-However, 33% of reviews reveal concerns in three domains.
-First, users report that image-generation features frequently fail to produce accurate or faithful visual outputs, particularly for complex or creative prompts (e.g., failed attempts to depict two people and an eagle wearing denim shorts in reviews 3 and 4 in Figure 9, respectively.) This finding aligns with broader evidence that AI image generators often collapse under high-complexity tasks (Petsiuk_Siemenn_Surbehera_Chin_Tyser_Hunter_Raghavan_Hicke_Plummer_Kerret_et_al_2022).
-Second, 8% of reviews highlight concerns regarding insufficient diversity and inclusion.
-For instance, review 5 in Figure 9 highlights a disparity in image quality between indigenous women and people of colour.
-Third, a minority of users (3% of reviews) highlight formatting and user interface issues in text outputs, such as grammatical errors, excessive length, inconsistent structure, indicating a need for improved polish as illustrated in review 6 in Figure 9.
-$\Rightarrow$ Implication 1.
-Gen-AI reviews consistently yield higher user ratings across app categories.
-Developers must recognize Gen-AI as a crucial driver of overall user satisfaction and strategically incorporate these features to remain competitive.
-However, they should treat Gen-AI capabilities as amplifiers of perceived value, not as substitutes for robust app quality, stability, and fair pricing.
-Even the most advanced AI functionality cannot compensate for poor technical reliability, weak customer support, or frustrating monetization structures.
-While users are generally satisfied with the overall performance of Gen-AI apps, a portion of their concerns relate to limitations in human-AI communication, particularly when the AI struggles to interpret nuanced or complex prompts.
-To better meet user expectations and improve usability, developers should prioritize enhancing the Gen-AI’s ability to understand and respond appropriately to more complex interactions.
-Although users generally regard Gen-AI outputs as high quality, enhancing image accuracy, diversity, inclusion, formatting, and user interface presentation remains essential for delivering truly reliable and user-centered content.
-Developers should prioritize these areas to refine trust, inclusivity, and satisfaction across varied use cases.
-3) Content Policy & Censorship.
-Content policy refers to rules and guidelines that define what type of AI-generated content is permissible.
-Censorship refers to automated or manual mechanisms used to block, modify, or suppress inappropriate, offensive, harmful, or illegal content based on those rules.
-Figure 10. Examples of reviews discussing Content Policy & Censorship
-Insights.
-In our review sample, 73% of users comment on censorship experience.
-Of these 23% report frustration with the filtering that often block benign content, adversely affecting usability.
-For example, the author of review 1 in Figure 10 reports getting censored for saying “hi.” On the other hand, 10% of the raised concerns relate to lack of effective filtering such as unethical imagery involving children bypassing filters as expressed in review 2 in Figure 10.
-Regarding adult content, 67% of concerned users express dissatisfaction with the excessive restriction of access.
-While 19% explicitly request an optional censorship toggle for adult users as illustrated in review 3 in Figure 10.
-Additionally, users express concern over unclear content ownership.
-Some users report that their creations were reused or uploaded by the platform without consent, as shown in review 4 in Figure 10.
-$\Rightarrow$ Implication 2.
-Developers need to strike the right balance in filtering content to avoid blocking genuinely harmful material while preventing harmless content.
-Ideally, developers could offer optional and adjustable censorship settings for mature or context-specific use.
-At the same time, platforms must clearly state who owns AI-generated outputs and how they can be reused, to maintain transparency and protect user trust.
-4) Features & Functionality.
-Features & Functionality refer to the Gen-AI-driven capabilities offered by the app, such as voice interaction and prompt autocompletion, which shape the user experience and determine which features users find valuable or lacking.
-Insights.
-In our sample, 75% of user reviews praise specific Gen-AI features, while 25% criticize them.
-For example, some users react negatively when a “stop-generation” button is removed and replaced with a voice mode feature, as illustrated in review 1 in Figure 11.
-While 11% of users enjoy the voice features, some describe them as useless.
-This indicates that even well‑received Gen-AI additions (such as voice interaction) can alienate users if they disrupt existing valued controls.
-Figure 11. Examples of reviews discussing Features & Functionality
-5) Utility & Use Cases.
-Utility & Use Cases refer to the practical roles Gen-AI plays in users’ daily lives, including enhancing accessibility, supporting professional tasks, and enabling creative expression across diverse contexts.
-Figure 12. Examples of reviews discussing Utility & Use Cases
-Insights.
-In our sample, 64% of reviews emphasize AI’s educational utility, such as helping with homework, essay writing, tutoring, and breaking down complex topics as shown in review 1 in Figure 12.
-Additionally, some users specifically praise how Gen-AI can act as an accessible on-demand learning assistant for dyslexic users, as commented in review 2 in Figure 12.
-Beyond education, 11% of reviews describe how the Gen-AI tools facilitate routine work tasks, such as inspiring new ideas, structuring project workflows, and even acting as fitness or guitar instructors, as illustrated in review 3 in Figure 12.
-6) Personalization & Customization.
-Personalization & Customization describe how Gen-AI systems adapt to user preferences and allow users to tailor outputs.
-This includes customizing output formats, editing generated content, and adopting preferred AI personas, e.g., voices or AI character roles.
-Figure 13. Examples of reviews discussing Personalization & Customization
-Insights.
-In our sample, 38% of reviews emphasize the AI’s inability to create a personalized experience.
-For example, review 1 in Figure 13, discusses the app’s inability to maintain shopping and to-do lists.
-Additionally, 10% of users request inline editing capabilities to support customization by refining generated content directly, rather than having to regenerate entirely.
-For instance, review 2 in Figure 13 suggests the option to edit lyrics in an already generated song rather than regenerate a new one entirely.
-Meanwhile, 9% of reviews request more diversity in AI characters and custom voices.
-This reflects a growing interest in personalizing AI identities to match preferences such as female voice options.
-For example, review 3 in Figure 13 describes the absence of customized voice features as a “deal breaker.”
-$\Rightarrow$ Implication 3.
-Developers should design Gen-AI features, such as voice mode, in a way that complements existing controls rather than replacing them.
-For instance, it is crucial to ensure new AI-driven tools coexist with familiar functionalities, so that developers can both innovate and maintain intuitive user experiences.
-Developers should implement customizable output options, enabling users to format, export, and save AI-generated content (e.g., lists, summaries) in a manner that suits their workflows.
-They should also add inline editing tools, so users can refine content directly without needing to regenerate responses.
-Additionally, offering persona customization, such as selectable character roles or voice styles, will enhance user engagement and make interactions feel more personalized.
-Despite the launch of tools like ChatGPT Edu (openaiChatGPTedu), which is tailored for academic contexts, Gen-AI platforms must go further by prioritizing accessibility-first design, for example, enabling assistive note-taking, voice transcription, and dyslexia support features.
-Developers should also integrate customizable, role-based modes or templates, such as “legal advisor,” “fitness coach,” or “creative brainstorming partner,” to make these tools more adaptable and effective across diverse professional and personal applications.
-7) Comparison to Other Apps. Comparison to Other Apps refers to comparing the AI app to other tools users have previously tried.
-Figure 14. Examples of reviews discussing Comparisons to Other Apps
-Insights.
-We find that 66% of the reviews draw comparisons to other AI-powered apps (e.g., comparing Microsoft Bing Search with Google).
-We also observe that 9% of the reviews compare the app with non-Gen-AI alternatives.
-For instance, review 1 in Figure 14 criticizes the forced integration of the AI chat feature, expressing a strong preference for traditional search functionality.
-Finally, 18% of reviews draw comparisons between the apps they review and general-purpose AI tools (e.g., ChatGPT), highlighting expectations shaped by user experiences with AI tools, as shown in review 2 in Figure 14.
-77% of the reviews that compare the app with ChatGPT believe that ChatGPT outperforms the app in question, while 15% prefer the performance of the domain-specific app.
-These comparisons highlight a key challenge for domain-specific Gen-AI app developers.
-Their apps must offer distinct value that goes beyond the general-purpose functionality of ChatGPT.
-8) Customer Support & Community.
-Customer Support & Community refers to reviews that address developers directly to express their experience, such as reporting user enjoyment with specific features or requesting to fix app shortcomings.
-Figure 15. Examples of reviews discussing Customer Support & Community
-Insights.
-In our sample, we find that 51% of users express enjoyment with specific features (e.g., review 1 in Figure 15), 19% suggest new features to enhance their experience (e.g., review 2 in Figure 15), 9% request to fix app shortcomings (e.g., review 3 in Figure 15), and 21% discuss generic enjoyment of the app without relating it to a specific feature.
-$\Rightarrow$ Implication 4.
-ChatGPT is a baseline comparator to users, in order to succeed developers must offer value beyond what ChatGPT offers.
-Developers should also consider creating a toggle button to allow users to opt in or out of AI-enabled features, such as AI search.
-The regular features may be sufficient for users (e.g., regular search can satisfy users’ searching needs, and AI search is not always necessary).
-User reviews highlight key features that developers should preserve when updating their apps.
-Prior work provided approaches that facilitate competitor analysis (Maram_LLM_Cure; PeerApps_SANER; Assi_Hassan_Tian_Zou_2021).
-For competing developers, analyzing user feedback from similar apps can also uncover desirable features worth adopting to remain competitive in the market.
-9) Human AI Interaction & Emotional Connection.
-Human AI Interaction & Emotional Connection reviews highlight the emotional and relational dimensions of human–AI interaction, particularly with conversational Gen-AI apps.
-Figure 16. Examples of reviews discussing Human AI Interaction & Emotional Connection
-Insights.
-We observe that users enjoy their interactions with conversational Gen-AI apps for multiple reasons.
-First, users enjoy a realistic interaction (i.e., closely similar to human dialogue).
-This perceived realism fosters a sense of companionship and emotional connection (e.g., review 1 in Figure 16).
-Second, reviews highlight the comfort in the emotional support provided by Gen-AI apps, particularly because the interaction lacks real human judgment, as shown in review 2 in Figure 16.
-This illustrates the importance of AI’s perceived neutrality and non-judgmental nature, especially when discussing sensitive or emotionally charged topics.
-Finally, users point to the therapeutic potential of conversational AI in providing mental and emotional support, such as managing mental health challenges (e.g., review 3 in Figure 16) and developing strategies to achieve personal goals.
-10) Accessibility & Inclusivity.
-Accessibility and inclusivity refer to the degree to which the Gen-AI apps accommodate users for affordability, multilingual support, age appropriateness, and usability for individuals with cognitive or physical impairments.
-Figure 17. Examples of reviews discussing Accessibility & Inclusivity
-Insights.
-Users discuss different accessibility and inclusivity topics including: 1) financial barriers (raised in 31% of the reviews, such as review 1 in Figure 17), 2) greater language support (raised in 22% of the reviews, such as review 2 in Figure 17), 3) disability-related needs (raised in 8% of the reviews, such as reviews 3 and 4 in Figure 17), 4) age-related concerns (raised in 10% of the reviews, such as Reviews 5 and 6 in Figure 17), and 5) general accessibility concerns (raised in 31% of the reviews).
-$\Rightarrow$ Implication 5.
-The perception of conversational Gen-AI apps as emotionally supportive and non-judgmental suggests their growing role as trusted companions, particularly in contexts involving emotional vulnerability.
-Developers should account for these evolving user expectations by ensuring that Gen-AI interfaces maintain a tone of empathy and neutrality, while also incorporating safeguards to avoid over-reliance or the illusion of professional psychological care.
-Furthermore, this trend highlights the need for interdisciplinary collaboration with mental health professionals to responsibly shape the affective design of Gen-AI systems used for emotionally sensitive interactions.
-The findings further highlight the importance of designing Gen-AI apps with diverse accessibility needs in mind.
-Developers should prioritize inclusive design by integrating features such as read-aloud support, customizable text presentation, and adaptive interfaces tailored to neurodiverse users (e.g., Attention Deficit Hyperactivity Disorder, Dyslexia and Dyspraxia).
-Observation 2.4: Developer Responses to Gen-AI Reviews Are Primarily Courteous or Diagnostic, With Few Immediate Resolutions.
-Table 8 shows that for Gen-AI apps, the most common developer response type is Acknowledgment and Appreciation (31%), which consists largely of brief thank-you messages.
-These replies most often accompany positive reviews and serve as courteous engagement rather than substantive technical follow-up.
-The second most frequent category, Request for Additional Information (30%), reflects the practical limits of troubleshooting Gen-AI issues through review threads.
-Even when space is not a constraint, users rarely provide the detailed context needed to diagnose problems that relate to model behavior, device configuration, or account-specific issues.
-Developers therefore redirect users to email or in-app support channels where logs, device information, or credentials can be shared securely.
-By contrast, Issue Resolved and Update Advised appears rarely (2%), whereas Acknowledged and Under Investigation is more common (10%).
-This pattern aligns with the nature of Gen-AI systems: many issues require model adjustments, infrastructure changes, or updates to safety and moderation mechanisms, none of which can be addressed immediately or within the constraints of a review thread.
-As a result, developers are more likely to acknowledge problems and communicate that work is underway than to confirm the issue has been resolved.
+Finding 2.3: User discussions of Gen-AI apps reveal key opportunities, including AI for Accessibility and Wellbeing, AI as a Collaborative Creative Tool, and AI Versatility, as well as challenges related to Managing User Expectations and AI Limitations, Balancing Content Moderation and Creative Freedom, and Strategic Integration of Gen-AI Features.
+Table 6 lists the Gen-AI topics with their descriptions and examples.
+Descriptive statistics on Gen-AI and non-Gen-AI topic categories can be found in Tables 12 and  13 in Appendix C.
+Building on the topic landscape identified, user feedback reveals several recurring themes that highlight both the potential and the limitations of Gen-AI apps.
+We synthesize these themes into three key opportunities and three challenges for developers designing Gen-AI apps as follows.
+Opportunity 1: Discussions of accessibility and wellbeing in Gen-AI apps highlight opportunities for inclusive and responsible design with appropriate safeguards for vulnerable users.
+User discussions reveal that Gen-AI apps support accessibility needs, learning assistance, and emotional wellbeing.
+These use cases are primarily reflected in the Utility & Use Cases and Emotional Connection topics, which receive some of the highest average ratings in our dataset (4.8).
+Reviews in the Utility & Use Cases topic show that users rely on Gen-AI tools for educational and everyday assistance.
+In our sample, 59% of reviews discussing this topic describe educational use cases such as tutoring, explaining difficult concepts, or assisting with homework.
+Users also highlight accessibility benefits, describing how Gen-AI systems help individuals with learning disabilities such as dyslexia or Attention-deficit/hyperactivity disorder (ADHD) by providing personalized explanations and on-demand assistance.
+Similarly, the Emotional Connection topic reveals that users frequently interact with Gen-AI systems as conversational companions.
+Reviews in this category are overwhelmingly positive (95%), with users describing their interaction with AI as a form of companionship, including having someone to talk to (21%), forming friendly relationships with the AI (6%), or using the system as a romantic companion (11%).
+Users report that interacting with the AI helps alleviate loneliness (14%), boredom (10%), or mental health concerns like anxiety and depression (5%).
+Although relatively rare, a small number of reviews (6%) raise concerns about potentially unhealthy interactions.
+For example, users describe the experience of speaking with AI characters as addictive or isolating, reporting that extended interactions reduced their engagement with real-world activities and relationships.
+Together, these discussions illustrate how Gen-AI apps extend beyond productivity tools to support accessibility and wellbeing.
+For developers, these findings highlight the opportunity to design Gen-AI systems that intentionally support educational accessibility and responsible conversational companionship while ensuring appropriate safeguards for vulnerable users.
+For policymakers and regulators, our findings highlight the need to examine the formation of human-AI relationships and to establish guidelines that address risks of over-attachment and dependency, particularly among vulnerable populations such as individuals with disabilities or mental health concerns.
+Opportunity 2: Discussions of Gen-AI as a collaborative creative tool highlight opportunities to design systems that support iterative co-creation and human-AI collaboration.
+User reviews describe Gen-AI systems as collaborative partners that assist users in ideation and creative exploration rather than fully automating tasks.
+This opportunity is primarily reflected in the Creative Potential and Utility & Use Cases topics.
+The Creative Potential topic receives overwhelmingly positive sentiment, with 98% of reviews expressing positive ratings.
+Users describe employing Gen-AI tools for a wide range of creative activities including graphic design, storytelling, roleplaying, music creation, and social media content production.
+For example, users report generating short videos for platforms such as TikTok or Instagram, experimenting with character creation, and using AI-generated content as inspiration for creative projects.
+Evidence of collaborative workflows also emerges in the Utility & Use Cases topic.
+Within this topic, reviews describe users working alongside AI to support tasks such as idea generation (4%), research assistance (4%), cooking guidance (1%), and programming support (1%).
+In these cases, the AI does not complete the entire task independently; instead, users interact iteratively with the system, refining prompts and incorporating AI-generated suggestions into their own work.
+These discussions suggest that users increasingly perceive Gen-AI systems as collaborative partners that augment human creativity and problem solving rather than replacing human effort entirely.
+For developers, these findings highlight the importance of designing Gen-AI systems that support iterative interaction, ideation support, and flexible co-creation workflows where users and AI contribute jointly to task completion.
+Gen-AI Topic Category Description (D)
+Example (E)
+AI Performance D: AI Performance captures user feedback on how effectively the AI generates responses, including aspects such as accuracy, speed, conversational quality, versatility across tasks in different domains, memory, and understanding.
+E: “Nice art but not really responding to my text well.
+It shows basic pictures not the complex combination I asked for so not that worth it for some more crazy experimentation ok for basic generation I guess.”
+Utility & Use Cases D: Utility & Use cases refers to reviews discussing the different ways they utilize Gen-AI in their lives.
+E: “ChatGPT is my new personal trainer and guitar teacher.”
+Content Quality D: Content Quality captures users’ assessments of the usefulness, coherence, creativity, engagement, structure, and presentation quality of AI-generated outputs.
+E: “Great pictures just not what I asked for.
+I have tried 7 times and not one was correct.
+Meaning I ask for two people with specific details to clothes etc.
+I get either one great drawing of one person or a distorted drawing of two people but only three legs Hilarious but I was not aiming for humor”
+Creative Potential D: Creative Potential captures how users leverage Gen-AI tools for creative expression, including content creation, design, storytelling, and interactive conversations.
+E: “Great for TikTok and Instagram short videos. ”
+Content Policy & Censorship D: Content policy refers to rules and guidelines governing the use of training and user data in Gen-AI systems, including how data is collected, used, and regulated.
+Censorship refers to automated or manual mechanisms used to block, modify, or suppress inappropriate, offensive, harmful, or illegal content based on those policies.
+E: “I’ve been having this app for almost a year now and I was enjoying it a lot.
+But recently whenever I’ve been having the most normal conversations with the AI bots it always gives me the info that somehow the messages don’t meet the guideline or something like that.
+I don’t know why it does that but I’m pretty disappointed.
+Just to clarify it again my conversations are absolutely normal.”
+Features & Functionality D: Features & Functionality captures user feedback on specific AI-enabled capabilities offered within the app, particularly voice interaction features and their usability.
+E: “Can’t even set an alarm by voice.”
+Emotional Connection D: Emotional Connection captures user experiences where Gen-AI systems function as conversational companions, sources of emotional support, or social interaction partners.
+E: “This app shouldn’t exist at all.
+It ruins lifes harms the environment.
+the addiction is a real thing.
+At first it’s fun but the more you use it the more it pulls you in to the point where you can’t put it down and I wonder how many people struggle with this in silence.
+It’s disgusting that you profit off of people’s well being.”
+Comparison to Other Apps D: Comparison to Other Apps captures instances where users evaluate the Gen-AI app by explicitly comparing it with alternative tools they have previously used.
+E: “This is by far the best app out there it’s better than Socratic, Chegg or any other app claiming to be able to help you with your homework.
+It gives you a detailed answer and it gives you why that is the answer.
+I love this app, I cannot believe i had just found it”
+Table 6. Gen-AI topics with descriptions and example reviews.
+Opportunity 3: Discussions of Gen-AI versatility highlight opportunities to design flexible systems that support evolving and user-driven use cases.
+User discussions demonstrate that Gen-AI systems are applied across a wide variety of domains and tasks, highlighting the versatility of generative capabilities.
+Evidence for this opportunity emerges across several topics, particularly AI Performance and Utility & Use Cases.
+Reviews discussing AI Performance indicate that users evaluate the AI based on its ability to handle diverse tasks and domains.
+In this topic, users comment on aspects such as accuracy (19%), response speed (18%), conversational quality (15%), and versatility across tasks (10%).
+Positive sentiment toward this topic is high, with 80% of reviews expressing satisfaction.
+Similarly, discussions in Utility & Use Cases demonstrate that users continuously discover new ways to apply Gen-AI tools in everyday contexts, including cooking advice, fitness guidance, studying religion, learning musical instruments, programming, general problem solving, and for life advice.
+These findings suggest that many Gen-AI use cases emerge organically from user experimentation rather than from predefined developer intentions.
+For developers, this highlights the importance of monitoring how users adapt AI systems to new tasks and designing flexible systems that can accommodate evolving user needs.
+Challenge 1: Discussions of user expectations and AI limitations highlight challenges in managing expectation gaps and communicating system capabilities transparently.
+Despite generally positive perceptions of Gen-AI capabilities, reviews highlight frustration when AI systems fail to meet user expectations.
+These concerns are primarily reflected in the AI Performance and Content Quality topics.
+While 80% of reviews discussing AI Performance are positive, approximately 18% of reviews report performance limitations, particularly related to poor AI memory, inaccurate responses, inconsistent conversational behavior, and slow response times.
+Similarly, 21% of reviews discussing Content Quality report dissatisfaction with generated outputs, often describing cases where the AI fails to follow prompts accurately or produces unexpected results.
+These issues become particularly pronounced when users submit complex prompts requiring multiple elements or relationships, where generated outputs may become partially incorrect or inconsistent.
+These findings highlight a key challenge for developers: managing the gap between user expectations and the actual capabilities of Gen-AI systems.
+Developers may benefit from designing mechanisms that communicate system limitations more transparently, such as providing feedback when prompts exceed model capabilities or offering suggestions for refining prompts.
+Challenge 2: Discussions of content moderation and creative freedom highlight challenges in balancing safety requirements with flexible and transparent moderation systems.
+Content moderation in regards to censorship and restriction of content emerges as one of the most controversial aspects of Gen-AI apps.
+This challenge is captured in the Content Policy & Censorship topic.
+User sentiment toward moderation mechanisms is mixed.
+While 28% of reviews express positive sentiment toward content policies, 55% express dissatisfaction.
+Among reviews mentioning filtering mechanisms, 77% report frustration with the system’s behavior.
+The most common complaint involves overly restrictive filtering, where harmless prompts are blocked unexpectedly (29%).
+Additionally, 23% of users express frustration that censorship limits their ability to generate mature or violent narratives, suggesting that current moderation policies restrict creative freedom.
+Conversely, a smaller portion of users (6%) argue that moderation systems fail to block inappropriate content effectively.
+These findings highlight the difficulty of balancing safety requirements with users’ desire for creative freedom.
+Users often hold conflicting expectations regarding acceptable content boundaries, particularly when engaging in creative storytelling or roleplaying scenarios.
+For developers and platform providers, these results highlight the importance of designing transparent and adaptable moderation systems that maintain safety standards while accommodating diverse creative use cases.
+Challenge 3: Discussions of Gen-AI feature integration highlight challenges in determining when AI capabilities add meaningful value while maintaining reliability.
+User feedback also reveals challenges related to how Gen-AI capabilities are integrated into mobile apps.
+This challenge is reflected in the Features & Functionality and Comparison to Other Apps topics.
+Reviews discussing Features & Functionality show that users appreciate conversational AI capabilities such as voice interaction, with 68% of reviews expressing positive sentiment.
+However, some users report that integrating AI features into existing systems can introduce reliability issues, particularly when AI-powered assistants fail to perform simple tasks that traditional systems previously handled effectively.
+Additionally, reviews in the Comparison to Other Apps topic demonstrate that users frequently evaluate Gen-AI apps relative to well-known AI tools such as ChatGPT and Photomath.
+These comparisons suggest that users expect new Gen-AI apps to offer capabilities that match or exceed existing systems.
+Together, these findings highlight an important design challenge: not every feature benefits from the integration of Gen-AI.
+Developers must carefully determine when AI capabilities meaningfully enhance the user experience and when simpler solutions may be more appropriate.
+Successfully integrating Gen-AI therefore requires thoughtful design decisions that balance innovation, reliability, and clear value for users.
+Summary of RQ 2 Our analysis shows that user discussions of Gen-AI apps center on generative capabilities and practical use cases, while negative feedback is primarily driven by broader app design issues.
+Reviews of Gen-AI features receive higher ratings, indicating generally positive user perceptions.
+Qualitative findings highlight opportunities in accessibility, creative collaboration, and diverse use cases, alongside challenges in managing user expectations, balancing content moderation with creative freedom, and determining effective integration of Gen-AI features.
+6. RQ3: What temporal trends emerge in Google Play Store user feedback on Gen-AI topics?
+To understand how users perceive Gen-AI features over time, we explore the temporal dynamics of user feedback and ratings in app reviews.
+Our goal is to uncover how user opinions about Gen-AI topics shift as Gen-AI capabilities mature and become more integrated into mobile apps.
+6.1. Experiment Setup
+We track the evolution of user feedback over time by conducting a two-level trend analysis.
+First, we conduct a high-level comparison of overall trends between Gen-AI and non-Gen-AI topics by examining how 1) average ratings and 2) the percentage of reviews evolve over time.
+This perspective helps highlight the increasing user attention to Gen-AI features.
+Second, we perform a more fine-grained, category-level analysis of Gen-AI topics to uncover detailed temporal patterns.
+We track the evolution of individual Gen-AI topic categories using clustering techniques and qualitative review analysis.
+The following steps outline how we conduct the analysis:
+Step 1: data preparation.
+We group reviews according to their topic category $(T)$ and time period $(P)$.
+Specifically, for every topic category $T$, we organize reviews into yearly intervals based on their submission dates.
+Then, for each topic category $T$, we construct two time series across all time periods $P$: 1) $Avg_{(T,P)}$ as the average rating of reviews in category $T$ posted during period $P$, and 2) $Pr_{(T,P)}$ as the percentage of Gen-AI reviews assigned to category $T$ within each app during period $P$, computed relative to the total number of Gen-AI reviews for that app in the same period.
+To avoid large apps dominating the results, we compute both measures at the app level and then average them across all apps that contain reviews in that period.
+Step 2: clustering methodology.
+To identify patterns in how Gen-AI topics evolve over time, we cluster the topics based on their trends.
+For each time series $Avg_{(T,P)}$ and $Pr_{(T,P)}$, we follow the three-step process listed below:
+  1. 1.
+    Clustering algorithm.
+K-Means clustering is a partitioning method that groups unlabeled data into $k$ clusters by minimizing within-cluster variance (Warren Liao, 2005).
+In our case, each topic category $T$ is represented across all time periods $P$, using either $Avg_{(T,P)}$ or $Pr_{(T,P)}$.
+These values form a time series vector for each topic, capturing how the topic evolves over time.
+We apply K-Means clustering to group topics with similar temporal trajectories.
+We select K-Means because it is efficient, flexible, and widely used for clustering in data mining tasks (Ikotun et al., 2023; Assi et al., 2025b).
+  2. 2.
+    Optimal number of clusters.
+To determine the optimal number of clusters ($k$), we use the elbow method, a widely adopted technique for evaluating clustering structure by examining the reduction in within-cluster variance as $k$ increases (Ashari et al., 2023).
+Through elbow plot evaluation and manual inspection of clusters, we identify $k=3$ as the point where additional clusters provide only marginal improvement for both Rating and Percentage.
+For $Pr_{(T,P)}$, the clusters reflect increasing, decreasing, and fluctuating patterns (Figure 6).
+For $Avg_{(T,P)}$, the clusters correspond to increasing, decreasing, and stable trends (Figure 7).
+To statistically validate the observed differences between Gen-AI and non-Gen-AI ratings over time, we perform a paired Wilcoxon signed-rank test on the yearly average ratings.
+Specifically, for each time period $P$, we compute the average rating of Gen-AI reviews and the average rating of non-Gen-AI reviews and compare the paired values across all periods.
+Step 3: manual trend investigation.
+To further examine trends in average ratings, we conduct an in-depth qualitative analysis of two representative topics.
+First, we analyze the opportunity AI for Accessibility and Wellbeing through the Emotional Connection topic to understand how users engage with Gen-AI emotionally over time.
+Second, we examine the challenge Balancing Content Moderation and Creative Freedom, through the topic Content Policy & Censorship which exhibits highly variable and often conflicting user perspectives, making it particularly informative for understanding evolving user concerns.
+To interpret the identified trends, we conduct a qualitative analysis of randomly sampled reviews from three key time periods.
+For each topic category and time period, we draw samples at a 95% confidence level with a 10% margin of error.
+We manually code a total of 407 reviews to identify recurring concerns, behavioral patterns, and shifts in user expectations, providing qualitative context for the observed quantitative trends.
+6.2. Findings
+Finding 3.1: The proportion of reviews discussing Gen-AI topics decreases over time but remains substantial.
+Figure 5 illustrates the proportion of reviews discussing Gen-AI topics over time.
+Gen-AI related reviews represent a large share of user feedback throughout the dataset, accounting for approximately 77% of reviews in 2019 and gradually declining to around 41% by 2026.
+Despite this decline, Gen-AI related discussions continue to represent a substantial portion of the overall review landscape.
+This pattern suggests that while the relative prominence of Gen-AI topics has decreased over time, Gen-AI capabilities remain an important focus of user feedback within mobile apps.
+Figure 5. Evolution of the percentage of Gen-AI reviews (left), and of the evolution of the average ratings of Gen-AI vs Non-Gen-AI topics (right) (Full dataset considered).
+Finding 3.2: Gen-AI reviews receive consistently higher ratings than non-Gen-AI reviews.
+As shown in Figure 5, reviews associated with Gen-AI topics consistently receive higher ratings than those discussing non-Gen-AI aspects across all time periods.
+While non-Gen-AI ratings decrease from 3.9 to 2.3, Gen-AI ratings remain higher, generally between 4.5 and 3.9.
+A paired Wilcoxon signed-rank test confirms this difference ($p<0.05$), indicating that Gen-AI-related feedback is rated significantly more positively than non-Gen-AI feedback.
+Figure 6. Trend Clusters for Percentage of Gen-AI Topics (Full dataset considered).
+Finding 3.3: Topic categories cluster into fluctuating, decreasing, and increasing trends based on their percentage of reviews over time.
+We cluster topic categories based on the temporal evolution of their review proportions into three trends as shown in Figure 6.
+The fluctuating cluster includes Comparison to Other Apps, Content Policy & Censorship, and Features & Functionality.
+These topics vary over time in response to external developments and emerging issues.
+Comparison to Other Apps rises sharply following the introduction of prominent Gen-AI systems (e.g., ChatGPT in 2022) (Akhtar, 2024), as users begin benchmarking apps against well-known alternatives such as GPT-based systems or domain-specific tools like Photomath.
+This topic continues to fluctuate as new technologies and competitors emerge.
+Similarly, Content Policy & Censorship fluctuates because it is driven by user dissatisfaction; users tend to discuss it primarily when encountering moderation issues, leading to intermittent spikes (notably around 2023 and 2026).
+Features & Functionality shows moderate variation, reflecting ongoing but less event-driven user feedback on app capabilities.
+The increasing cluster includes AI Performance, Content Quality, Emotional Connection, and Utility & Use Cases.
+These topics show a steady rise over time, indicating that as Gen-AI systems improve, users increasingly engage with and evaluate core system capabilities.
+In particular, the growth of AI Performance and Content Quality reflects improving model capabilities and rising user expectations.
+At the same time, the increase in Emotional Connection and Utility & Use Cases suggests that users are moving beyond novelty, increasingly discussing how they interact with AI and the diverse contexts in which they apply it.
+Overall, these trends indicate a shift from novelty-driven engagement to performance- and use-oriented evaluation, highlighting the maturation of user expectations as Gen-AI technologies evolve.
+Figure 7. Trend Clusters for Average Ratings of Gen-AI Topics (Full dataset considered).
+Finding 3.4: Emotional Connection with AI evolves from novelty to dependency and addiction concerns, while declining developer engagement with these emerging risks highlights the need for responsible design and safeguards.
+The trend is shown in Figure 7.
+We summarize the sub-topics discussed in each year in Figure 8.
+Emotional interaction with AI shifts over time from novelty-driven companionship to more sustained and structured forms of engagement.
+Early use centers on conversational companionship, but with time, users increasingly rely on AI for emotional support, social confidence, and expression.
+As adoption matures, interactions become more habitual, with users describing consistent reliance on AI for communication and coping.
+This progression introduces new risks, including over-reliance, addiction, and disengagement from real-world interactions, particularly among emotionally vulnerable users.
+Overall, the trend reflects a transition from exploratory use to dependency-related challenges, highlighting the need for responsible design and mechanisms that promote healthy usage boundaries.
+Developer responsiveness in Emotional Connection declines sharply from 55% in 2022 to below 2% after 2024, despite the emergence of more complex concerns such as addiction and emotional dependency, indicating reduced attention to longer-term user risks.
+Figure 8. Subtopics of reviews discussing Emotional Connection (Representative subset considered).
+Figure 9. Subtopics of reviews discussing Content Policy & Censorship (Representative subset considered).
+Finding 3.5: Content moderation concerns shift from trust and privacy to censorship, accompanied by increased developer responsiveness.
+The trend is shown in Figure 7 and the subtopics are summarized in Figure 9.
+Content moderation concerns evolve from early distrust around data use and privacy toward debates about censorship, bias, and platform control.
+As adoption matures, users focus less on whether their data is safe and more on how content is filtered and governed.
+A key tension emerges between users reporting excessive moderation (restrictions on regular use) and those requesting reduced moderation for greater freedom in generating mature content.
+While mentions of both increase over time, requests for reduced moderation grow more strongly than calls for stricter moderation in response to unintended explicit outputs.
+New concerns around age-related restrictions and verification further complicate user experience and trust.
+Overall, the findings point to a growing mismatch between platform governance mechanisms and evolving user expectations, highlighting the need for more transparent, flexible, and user-aware moderation strategies.
+Developer responses increase for content-related concerns, rising from 3% in 2019 to 20% in 2026, reflecting a focus on actionable issues like censorship.
+In contrast, responses to emotional connection concerns decline, leaving human-centered risks comparatively under-addressed.
+Summary of RQ 3 To Our temporal analysis shows that (1) the proportion of Gen-AI discussions decreases over time but remains substantial, (2) Gen-AI reviews consistently receive higher ratings than non-Gen-AI reviews, (3) emotional interaction with AI evolves from novelty to dependency concerns, and (4) content moderation challenges shift from early trust and privacy issues to tensions around censorship, bias, and platform control.
+7. Discussion
+In this section, we first present a platform-comparison analysis to assess the generalizability to the App Store.
+Second, we present an analysis of developers’ responses to explore how they respond to Gen-AI reviews.
+7.1. Platform Comparative Analysis
+To assess whether insights derived from Google Play Store reviews generalize to other mobile app stores, we apply SARA to the App Store versions of apps in our dataset and compare the resulting topic distributions across platforms.
+The analysis focuses on the 59 apps that are available on both the Google Play Store and the App Store, enabling a controlled comparison of user feedback for the same apps across different review environments.
+Table 7. Distribution of Gen-AI topic categories in App Store (Astore) and Google Play Store (GPS) reviews (full dataset considered).
+Topic Category Astore Reviews GPS Reviews Astore Rating GPS Rating
+Utility & Use Cases 74,089 (32.9%) 111,736 (14.3%) 4.9 4.8
+AI Performance 23,846 (10.6%) 130,492 (16.7%) 3.3 4.2
+Content Quality 16,477 (7.3%) 86,888 (11.1%) 3.7 3.6
+Content Policy & Censorship 12,299 (5.5%) 16,934 (2.2%) 2.4 2.5
+Creative Potential 11,607 (5.2%) 57,551 (7.4%) 4.9 3.6
+Emotional Connection 1,071 (0.5%) 6,446 (0.8%) 3.8 4.7
+Comparison to Other Apps 983 (0.4%) 1,154 (0.1%) 4.4 4.4
+Features & Functionality 740 (0.3%) 18,283 (2.3%) 4.0 2.3
+Inclusivity 743 (0.3%) None 2.3 None
+Topic categories remain consistent across platforms.
+Across app categories, the overlap of topic categories between Google Play Store and App Store reviews ranges from 64% to 100% (average 81%), indicating that users discuss largely similar aspects of Gen-AI apps regardless of platform.
+However, the emphasis placed on these topics differs across platforms.
+App Store users more frequently describe how they apply Gen-AI tools in practice, whereas Google Play Store users more often evaluate the technical capabilities of the underlying Gen-AI models.
+As shown in Table 7, App Store reviews are dominated by discussions of Utility & Use Cases, which account for 32.9% of Gen-AI related reviews.
+In contrast, this topic accounts for only 14.3% of Gen-AI discussions in Google Play Store reviews, whereas AI Performance emerges as the most frequently discussed topic (16.7%).
+App Store users tend to be more critical when evaluating the reliability and quality of AI-generated outputs.
+Discussions of AI Performance receive a lower average rating on the App Store (AGR = 3.3) than on Google Play Store (AGR = 4.2).
+Differences also emerge in how users discuss the Emotional Connection of Gen-AI systems.
+In Google Play Store reviews, this topic often reflects positive experiences such as companionship or emotional support (AGR = 4.7), while App Store reviews more frequently frame similar experiences in terms of potential risks, including addiction or negative effects on mental health (AGR = 3.8).
+App Store users may be more attentive to issues related to privacy, bias, and responsible AI behavior.
+Ethical concerns appear more prominently in App Store reviews.
+Inclusivity-related issues, such as gender and ethnicity bias in generated content, form a distinct topic category in App Store reviews but do not appear frequently enough in Google Play Store reviews to constitute a separate category.
+Similarly, privacy concerns captured under Content Policy & Censorship appear as a topic across two App Store app categories, while they are only discussed in one Google Play Store app category.
+Summary of Platform Analysis Overall, these results suggest that while the primary dimensions of user feedback are consistent across platforms, the emphasis placed on them differs.
+App Store users tend to highlight practical usage scenarios and ethical considerations, whereas Google Play Store users more frequently evaluate the technical performance of Gen-AI systems.
+Monitoring feedback across multiple platforms, therefore, provides complementary insights into user expectations and emerging concerns surrounding Gen-AI apps.
+7.2. Developer Reply Analysis
+This section examines how developers engage with user feedback in Gen-AI app reviews.
+While previous sections focused on user perspectives, developer replies reveal how teams assess reported issues, manage user expectations, and respond to the practical constraints of diagnosing AI-related problems through public review channels.
+For developer responses, we conduct two qualitative analyses: first, coding a representative sample of 96 replies across all app categories to characterize general response patterns, and second, analyzing 281 replies from three focal categories (Games, Education, and Entertainment) to capture category-specific developer practices.
+Developer responses are predominantly generic and template-based, with limited occurrence of substantive issue resolution.
+The most common developer response type, as shown in Table 8, is Acknowledgment and Appreciation (31%), representing typically short thank-you messages posted in response to positive reviews.
 Table 8. Summary of developer response types, including descriptions, number of reviews, and percentage distribution (Representative subset considered).
@@
 Total 96 100%
+Developers redirect users to external support channels where complete context and diagnostic information can be shared securely.
+The second most common category, Request for Additional Information (30%), reflects the practical limitations of diagnosing Gen-AI issues through review threads.
+Users report incorrect or unexpected AI outputs without providing the original prompt, generated response, or screenshots.
+By contrast, confirmed resolutions are rare: only 2% of replies indicate that an issue has been resolved, while 10% acknowledge that developers are investigating the problem.
+This pattern reflects the nature of Gen-AI systems, where many issues require model adjustments, infrastructure changes, or updates to safety mechanisms that cannot be deployed immediately.
+As a result, review replies often serve as an initial point of contact rather than a channel for resolving AI-related failures directly.
+General developer reply practices reported in prior work extend to Gen-AI apps.
 Hassan et al.
-(developer_response_safwat) analyzed developer replies across a broad set of mobile apps and identified recurring response drivers such as “thank the user”, “ask for more details”, “provide guidance”, “ask for endorsement”, “advertise other products”, wait for updates, and “justify advertisements”.
-Many of these practices align closely with our taxonomy.
-For example, their “thank the user” category maps to our “Acknowledgment and Appreciation,” their “ask for more details” category maps to our “Request for Additional Information,” and their “justify the advertisements” category maps onto our broader Clarification or Policy Rebuttal, which includes explanations about subscription models, premium features, and content restrictions.
-Notably, their work does not treat “issue resolved” as a dominant response driver, and we similarly observe that confirmed resolutions are relatively rare in our Gen-AI dataset (2%).
-We report the category explicitly for completeness of Gen-AI developer behaviors, but its low prevalence mirrors prior findings rather than contrasting with them.
-Overall, the strong correspondence between the two taxonomies suggests that general response practices extend to Gen-AI apps, while our domain-specific categories capture the additional emphasis developers place on clarifying AI-related behavior and managing expectations around the evolving nature of Gen-AI features.
-Observation 2.5: App categories show highly distinctive developer response patterns.
-Table 9 reveals variation in how developers across different app categories respond to user feedback.
-Some categories exhibit exceptionally high engagement, while others show very limited interaction.
-For example, Games stands out with both the highest reply ratio (56%) and the longest average response delay (44.7 days).
-In contrast, Education demonstrates the shortest response times, with developers replying in less than a day on average (0.7 days).
-Meanwhile, Entertainment apps show the lowest reply ratio among categories with meaningful activity (2%).
-These contrasting patterns suggest that developer responsiveness is shaped not only by category size but also by the operational demands, user expectations, and support norms unique to each domain.
+(Hassan et al., 2018) identified several common developer response drivers in mobile app ecosystems, including thanking users, requesting additional details, and providing guidance.
+Many of these patterns align closely with our taxonomy.
+For example, their “thank the user” category corresponds to our Acknowledgment and Appreciation.
+Overall, this correspondence suggests that general developer reply practices extend to Gen-AI apps, while our categories capture the additional emphasis on clarifying AI behavior and managing expectations around evolving AI capabilities.
+Developer response behavior differs across app categories, with some (i.e Education) prioritizing fast, template-based replies and others (i.e Games) slower but more substantive engagement.
+Developer response behavior also varies substantially across app categories as shown in Table 9.
+For example, Games exhibits both the highest reply ratio (56%) and the longest response delay (44.7 days).
+Our manual analysis suggests that this delay is associated with responses that confirm issues have been fixed, which represent 30% of replies in this category compared to only 2% in the overall dataset.
+Developers often appear to wait until a patch is deployed before responding, resulting in slower but more substantive communication.
 Table 9. Developer response activity across app categories, including total reviews, total developer replies, overall reply ratio, and average response delay (Full dataset considered).
@@
 All Categories 676,066 115,164 20% 10.2
-Our manual analysis of developer responses in the Games category helps explain why it exhibits both the highest reply ratio and the longest response delay.
-Responses in which developers report that an issue has been fixed account for 30% of all replies in this category, far exceeding the 2% observed in the overall dataset.
-This emphasis on confirmed fixes aligns with the long response delay shown in Table 9.
-This likely reflects that developers often wait until a patch is deployed before responding, resulting in slower but more substantive communication.
-Responses indicating that an issue is acknowledged and under investigation represent 13% of replies, slightly higher than the overall 10%.
-Together, these categories show that Games developers place strong emphasis on technical issue resolution and ongoing investigation.
-Generic thank-you messages appear in only 12% of replies, considerably lower than the overall 31%.
-The remaining responses highlight active engagement with user issues: 28% request additional information, 10% provide troubleshooting guidance, and 8% offer clarifications or policy rebuttals.
-Overall, developer communication in Games is markedly problem-oriented and resolution-driven, which leads to both high engagement and extended response timelines.
-A closer look at Education apps helps explain why this category exhibits the fastest average developer response time.
-In our manual sample, the majority of Education responses fall under Acknowledgment and Appreciation (73%), which is markedly higher than both the overall rate of 31% and the 12% observed in Games.
-These replies are typically short, template-based thank-you messages that require minimal effort to generate, allowing developers to respond quickly and consistently.
-Education apps also show a comparatively high rate of Request for Additional Information (16%).
-This proportion is notably larger than the overall sample and reflects a common pattern in Gen-AI Education reviews.
-Users frequently report that the app produced incorrect answers, yet they often omit the original prompt, the model’s response, or screenshots of the interaction.
-Without these details, developers cannot diagnose or reproduce the issue within the review thread.
-As a result, they quickly reply by directing users to contact support channels where complete context can be shared securely.
-This response type is therefore both common and fast, since it does not require problem investigation within the review platform itself.
-By contrast, response types requiring deeper engagement occur far less frequently in Education, including Guidance and Troubleshooting Tips (2%), Acknowledged and Under Investigation (6%), and Issue Resolved and Update Advised (1%).
-Together, these patterns show that Education apps tend to rely primarily on quick acknowledgments or requests for additional information, which contributes to the category’s exceptionally short response delay.
-Entertainment apps show the lowest reply ratio in our dataset, and the distribution of response types helps explain this pattern.
-Acknowledgment and Appreciation responses account for only 14%, far below the overall rate of 31%.
-This indicates that developers in this category do not typically reply to routine or positive reviews and instead reserve responses for situations where user action or developer follow-up is required.
-Correspondingly, a large share of replies fall under Request for Additional Information (38%) and Guidance and Troubleshooting Tips (22%), suggesting that developers primarily intervene when users report crashes, playback issues, or account problems that require diagnostic detail.
-The relatively high proportion of Acknowledged and Under Investigation (19%) further shows that many of these reports involve technical issues that cannot be resolved immediately, which may discourage broader engagement with lower-priority feedback.
-Together, these patterns indicate that Entertainment developers focus their limited response efforts on actionable problem reports rather than general user engagement, resulting in a lower overall reply ratio.
-Observation 2.6: Developer responsiveness varies substantially across Gen-AI topic categories.
-Table 10 summarizes developer reply behavior for each Gen-AI topic category.
-The topic with the highest reply ratio is Accessibility & Inclusivity (47%), despite having the smallest number of reviews.
-This indicates that developers treat accessibility-related feedback as a high-impact area that warrants immediate and consistent attention.
-Such reviews often report barriers to use for specific user populations, making them critical from both ethical and product-quality perspectives.
-The short average delay (0.2 days) further reflects the urgency developers assign to this category.
-Beyond accessibility, the next highest reply ratios are found in Utility & Use Cases (18%) and Content Quality (14%), suggesting that developers are highly responsive to feedback about how users apply AI features and whether the generated outputs meet expectations.
-These topics directly affect user satisfaction and daily usability, which likely motivates more frequent developer engagement.
-In contrast, the categories with the lowest reply ratios are Customer Support & Community (1%), Human–AI Interaction & Emotional Connection (2%), and Content Policy & Censorship (5%).
-These topics often involve complaints about communication quality, emotional tone, safety constraints, or moderation decisions.
-Such issues are typically systematic, tied to model behavior or policy rules that cannot be quickly adjusted, which may explain the lower response frequency and the longer average delays seen in these categories.
-This contrast between highly actionable categories and structurally constrained categories reinforces the idea that responsiveness is shaped by whether an issue is immediately fixable or requires deeper system changes.
-Overall, these patterns show that developers prioritize issues that are actionable in the short term (e.g., accessibility barriers, inaccurate outputs, task-level failures) while responding less frequently to concerns that require deeper model changes or policy revisions.
-This reflects a practical triaging strategy driven by the nature of Gen-AI systems and the feasibility of immediate fixes.
+In contrast, Education apps demonstrate the shortest response delays (0.7 days).
+In our manual sample, 73% of responses fall under Acknowledgment and Appreciation, suggesting that developers frequently rely on quick, template-based replies that require minimal investigation.
+Education apps also show a relatively high proportion of Request for Additional Information responses (16%), reflecting situations where users report incorrect AI answers without providing sufficient context for developers to reproduce and resolve the issue.
+Meanwhile, Entertainment apps show the lowest reply ratio (2%).
+Developers in this category rarely respond to routine or positive feedback and instead appear to focus their replies on actionable problem reports.
+A large proportion of replies fall under Request for Additional Information (38%), Guidance and Troubleshooting Tips (22%), and Acknowledged and Under Investigation (19%), suggesting that developers primarily intervene when users report crashes, playback failures, or account issues that require additional diagnostic detail.
+Developer responsiveness varies across Gen-AI topics, prioritizing actionable issues over systemic and human-centered concerns (Table 10).
+The highest reply ratio appears in Accessibility & Inclusivity (47%), despite this category containing the smallest number of reviews.
+This suggests that developers treat accessibility-related feedback as high-impact issues requiring immediate attention.
+Higher engagement also appears in Utility & Use Cases (18%) and Content Quality (14%), both of which directly affect user satisfaction and day-to-day usage of Gen-AI functionality.
 Table 10. Developer response patterns across Gen-AI topic categories, including total reviews, total developer replies, reply ratio, and average response delay (Full dataset considered).
@@
 Content Policy & Censorship 27,521 1,397 5% 7.9
-Human–AI Interaction & Emotional Connection 25,142 542 2% 7.5
+Emotional Connection 25,142 542 2% 7.5
 Features & Functionality 16,275 1,892 12% 5.5
@@
 All Topics 238,168 27,619 12% 4.9
-Observation 2.7: Developers respond less frequently to Gen-AI reviews than to non-Gen-AI reviews as illustrated in Figure 18.
-We compared per-app developer response rates for Gen-AI and non-Gen-AI reviews using a paired non-parametric Wilcoxon signed-rank test ($n=164$ apps).
-Developers replied to a median of 9.6% of Gen-AI reviews versus 17.7% of non-Gen-AI reviews.
-Although the median per-app difference is modest (0.7-0.7 percentage points), the effect is statistically significant ($W=2911$, $p<0.001$), indicating a consistent pattern across apps.
-This gap aligns with the qualitative patterns observed earlier.
-Gen-AI reviews often describe failures rooted in model behavior, data quality, or moderation mechanisms, which are difficult to diagnose from short review text.
-By contrast, non-Gen-AI reviews commonly involve UI issues, feature requests, or account problems that are easier to acknowledge, triage, or resolve within the constraints of the review interface.
-As a result, developers appear slightly more willing to engage with non-AI feedback where their replies can directly influence user experience.
-In addition to analyzing user perspectives, we also examine how developers respond to Gen-AI reviews, capturing patterns of engagement, resolution, and support practices across app categories.
-Observation 2.8: Developers respond to Gen-AI and non-Gen-AI reviews at similar speeds.
-Figure 18 compares developer response delays for Gen-AI and non-Gen-AI reviews across all apps.
-While Gen-AI reviews show a slightly lower median delay (a difference of 0.15-0.15 days), this difference is not statistically significant.
-In practical terms, once developers choose to reply, they tend to respond at comparable speeds regardless of whether the review discusses Gen-AI functionality or non-AI aspects of the app.
-This suggests that the lower engagement with Gen-AI reviews observed in Observation 2.7 reflects differences in response likelihood rather than response promptness.
+In contrast, topics such as Customer Support & Community (1%), Emotional Connection (2%), and Content Policy & Censorship (5%) receive fewer developer replies.
+These issues often involve systemic behavior of AI models or platform policies that cannot be easily modified, which may explain the lower engagement levels.
+Overall, this pattern suggests that developers prioritize feedback that can be addressed through short-term fixes (e.g., accessibility barriers or incorrect outputs), while responding less frequently to concerns that require deeper model changes or policy revisions.
+Developers respond less frequently to Gen-AI reviews than to non-Gen-AI reviews, despite similar response delays.
+We compare per-app developer response rates for Gen-AI and non-Gen-AI reviews using a paired Wilcoxon signed-rank test ($n=164$ apps).
+Developers replied to a median of 9.6% of Gen-AI reviews compared to 17.7% of non-Gen-AI reviews, indicating that developers engage less frequently with AI-related feedback.
+Although the per-app median difference is modest (0.7-0.7 percentage points), the effect is statistically significant ($W=2911$, $p<0.001$), suggesting a consistent pattern across apps.
+One possible explanation is that Gen-AI failures often stem from model behavior, training data limitations, or safety and moderation mechanisms.
+Diagnosing such issues typically requires additional context, such as the original prompt, the generated response, the user’s device configuration, or backend logs.
+Because this information is rarely included in short review text and cannot easily be exchanged through public review threads, developers may be less likely to engage directly with Gen-AI feedback in this setting.
+Finally, once developers decide to reply, response speed appears similar across review types.
+Response delay distributions for Gen-AI and non-Gen-AI reviews do not differ significantly, indicating that the lower engagement with Gen-AI feedback reflects differences in response likelihood rather than response promptness.
+As illustrated in Figure 10, the distribution of response ratios across apps is centered below one, indicating that developers generally respond less frequently to Gen-AI reviews.
 Figure 10.
 Comparison of developer responsiveness across apps.
-Left: Relative likelihood of developer response to Gen-AI reviews compared to non–Gen-AI reviews, expressed as the ratio of per-app response probabilities (values below 1 indicate lower responsiveness to Gen-AI reviews).
-Right: Distribution of mean developer response delays per app for Gen-AI and non–Gen-AI reviews; the two delay distributions do not differ significantly.
+Left: Relative likelihood of developer response to Gen-AI reviews compared to non-Gen-AI reviews, expressed as the ratio of per-app response probabilities (values below 1 indicate lower responsiveness to Gen-AI reviews).
+Right: Distribution of mean developer response delays per app for Gen-AI and non-Gen-AI reviews; the two delay distributions do not differ significantly.
 (Full dataset considered.)
-$\Rightarrow$ Implication 6.
-Taken together, our developer response findings show that review threads in Gen-AI apps primarily serve as an initial point of contact rather than a channel for resolving AI-specific failures.
-Developers face structural barriers when responding to Gen-AI feedback, since many issues require information that cannot be shared in public reviews or involve model- or policy-level changes that cannot be deployed quickly.
-As a result, developers tend to reserve detailed engagement for issues that are actionable in the short term and redirect users elsewhere when deeper diagnostic context is required.
-These patterns highlight a broader platform-level limitation: while app stores surface user feedback effectively, they do not provide the depth of interaction required to investigate or clarify Gen-AI behaviours.
-There is therefore an opportunity for platform owners to introduce more structured, built-in support mechanisms that complement review sections and enable richer, privacy-preserving communication channels.
-Such improvements would better support both developers and users as Gen-AI apps become more complex and more central to mobile ecosystems.
-Summary of RQ 2 From our analysis of user reviews across diverse Gen-AI app categories, we find that users are increasingly experimenting with prompts, offering feedback, and forming emotional connections with non-human agents.
-They expect fast, creative, and context-aware responses, along with personalization and transparency.
-For developers, this means moving beyond baseline functionality to support user control (e.g., ability to disable Gen-AI features) and differentiated value compared with AI tools (e.g., ChatGPT).
-Our developer-response analysis reinforces this point: developers most often engage with actionable, user-reported issues and struggle to address Gen-AI failures directly within the review platform.
-This suggests that better support channels will be essential for meeting rising user expectations.
-6. RQ3: What temporal trends emerge in user feedback on Gen-AI topics?
-To understand how users perceive Gen-AI features over time, we explore the temporal dynamics of user feedback and ratings in app reviews.
-Our goal is to uncover how user opinions about Gen-AI topics shift as Gen-AI capabilities mature and become more integrated into mobile apps.
-6.1. Experiment Setup
-We track the evolution of user feedback over time by conducting a two-level trend analysis.
-First, we conduct a high-level comparison of overall trends between Gen-AI and non-Gen-AI topics by examining how 1) average ratings and 2) the percentage of reviews evolve over time.
-This perspective helps highlight the increasing user attention to Gen-AI features.
-Second, we perform a more fine-grained, category-level analysis of Gen-AI topics to uncover detailed temporal patterns.
-We track the evolution of individual Gen-AI topic categories using clustering techniques and qualitative review analysis.
-Based on the observed rating dynamics, we select three representative topic categories for deeper investigation: Content Quality (decreasing trend), AI Performance (stable trend), and Content Policy & Censorship (increasing trend).
-The following steps outline how we conduct the analysis:
-Step 1: data preparation.
-We group reviews according to their topic category $(T)$ and time period $(P)$.
-Specifically, for every topic category $T$, we organize reviews into half-year intervals (i.e., every 6 months) based on their submission dates.
-Then, for each topic category $T$, we construct two time series across all time periods $P$: 1) $Avg_{(T,P)}$ as the average rating of reviews in category $T$ posted during period $P$, and 2) $Pr_{(T,P)}$ as the percentage of reviews in period $P$ that belong to category $T$.
-We exclude the topic category Accessibility & Inclusivity because it only contains 96 reviews in total, which are not enough to conduct a time series analysis.
-Step 2: clustering methodology.
-To identify patterns in how Gen-AI topics evolve over time, we cluster the topics based on their trends.
-For each time series $Avg_{(T,P)}$ and $Pr_{(T,P)}$, we extract slope and standard deviation as features following the three-step process listed below:
-  1. 1.
-    Features extraction.
-Feature selection in clustering involves identifying the most relevant features to uncover patterns in unlabeled data (Dash_Koot_2009).
-Feature selection enhances interpretability, aligns with research goals  (Kreienkamp_Agostini_Monden_Epstude_DeJonge_Bringmann_2025), and enables grouping of topics with similar temporal trends (e.g., increasing, fluctuating, or stable).
-For each topic category $(T)$, we compute the slope $m$ and the standard deviation $\sigma$ for both $Avg_{(T,P)}$ and $Pr_{(T,P)}$.
-Selecting these features supports more meaningful clustering outcomes by capturing both the directional trend $m$ and variability $\sigma$ of each topic’s trend.
-It further enables the grouping of topics that exhibit similar temporal behaviours, such as consistently increasing or decreasing and fluctuating or stable patterns.
-  2. 2.
-    Clustering algorithm.
-K-Means clustering is a partitioning method that groups unlabeled data into ($k$) clusters by minimizing within-cluster variance (Warren_Liao_2005).
-We apply K-Means clustering to group topics based on their slope $m$ and standard deviation $\sigma$ values.
-We select K-Means because it is efficient, flexible, and widely used for clustering in data mining tasks (Ikotun_Ezugwu_Abualigah_Abuhaija_Heming_2023; 10.1145/3721125).
-K-Means is suitable for low-dimensional feature spaces (Ikotun_Ezugwu_Abualigah_Abuhaija_Heming_2023) and hence enables us to identify generalizable trend patterns across different topics and apps.
-  3. 3.
-    Optimal number of clusters.
-To determine the optimal number of clusters ($k$), we compute Silhouette Scores, one of the most widely adopted methods for evaluating clustering quality (Ashari_DwiNugroho_Baraku_NovriYanda_Liwardana_2023), across multiple values of $k$.
-Based on the silhouette results, we select the $k$ value closest to 1.
-The optimal silhouette scores correspond to kk=3 for both the average rating trend series $Avg_{(T,P)}$ and the review percentage trend series $Pr_{(T,P)}$.
-We complement this with qualitative validation through manual inspection, confirming that kk=3 yields the most interpretable and distinct groupings.
-Step 3: manual trend investigation.
-To interpret the identified clusters, we perform qualitative analysis on statistically significant random review samples from key time periods.
-For each selected topic category, we identify three representative intervals, typically from the beginning, a period of fluctuation, and the end of the timeline.
-From each interval, we extract a statistically representative random sample of reviews using a 95% confidence level and a 10% margin of error.
-We manually code the sampled reviews to identify recurring user concerns, behavioural patterns, and shifts in expectations.
-This process provides qualitative context that enhances our understanding of the observed quantitative trends.
-6.2. Findings
-Observation 3.1: Gen-AI topics consistently receive higher user ratings than non-Gen-AI topics, with an average difference of 0.8 rating stars over time.
-Specifically, the average rating for Gen-AI topics is 4.0, compared to 3.2 for non-Gen-AI topics.
-As illustrated in Figure 19, the percentage of reviews discussing Gen-AI topics increases over time.
-The percentage of Gen-AI reviews increases by 25%, from 18% in 2022-H1 to 43% in 2024-H2, reflecting a growing user focus on Gen-AI features.
-This growth indicates that users are increasingly evaluating apps through the lens of their Gen-AI features.
-As AI becomes a central part of the user experience, developers can no longer treat it as an add-on; users expect it to perform reliably and meaningfully enhance the app’s core value.
-The rising share of Gen-AI reviews therefore signals that user satisfaction and engagement increasingly depend on how well these features are designed, integrated, and maintained.
-Despite this surge in volume, the average rating for Gen-AI topics remains stable over time, as shown in Figure 20 with only a slight change of ± 0.2 stars across the observed period.
-Upon further investigation, we find that this stability is the result of diverging trends across individual topic categories.
-As illustrated in Figure 21, some topic categories (e.g., Content Policy & Censorship) exhibit increasing trends, others (e.g., Content Quality) decline, while several topics (e.g., AI Performance) remain stable.
-These opposing directional trends counterbalance each other when aggregated, leading to the observed overall stability in Gen-AI topic ratings.
-The following observations examine these diverging topic-specific trajectories in greater detail.
-Figure 19. Evolution of the Percentage of Gen-AI Reviews (Full dataset considered).
-Figure 20. Evolution of the Average Ratings of Gen-AI vs Non-Gen-AI Topics (Full dataset considered).
-Gen-AI topics analysis.
-As established in our experimental setup, $k=3$ yields the most interpretable and well-separated clusters.
-These clusters reflect three general trends: increasing, decreasing, and stable as shown in Figure 21 and 22 for the average ratings and the percentage of reviews, respectively.
-To deepen our understanding of the observed rating trends, we perform a manual trend investigation, i.e., a qualitative analysis, on three representative topic categories, each corresponding to a distinct trend direction.
-For the Decreasing trend, we select Content Quality due to the unexpected nature of its decline, particularly in light of ongoing advancements in AI capabilities as shown in Figure 21.
-For the Stable trend, we examine AI Performance, where we had anticipated an upward trend reflecting improvements in system performance.
-For the Increasing trend, we analyze Content Policy & Censorship to investigate how user perceptions of policy changes have evolved and to uncover the factors contributing to rising satisfaction in this area.
-Figure 21. Trend Clusters for Average Ratings of Gen-AI Topics (Full dataset considered).
-Figure 22. Trend Clusters for Percentage of Gen-AI Topics (Full dataset considered).
-Observation 3.2: The Content Quality topic category shows a decline in rating over time, despite improvements in Gen-AI technology.
-As shown in Figure 21, the average rating for Content Quality decreases from 4.5 in 2022-H1 to 3.3 in 2024-H2.
-To better understand the unexpected downward trend in Content Quality ratings, we conduct a manual analysis of user reviews across three strategically selected time periods: 2022-H1 (corresponding to the highest average rating), 2022-H2 (marking a period of sharp decline), and 2024-H2 (representing the lowest observed average rating).
-This sampling allows us to investigate user behavior before, during, and after the major shifts in rating averages.
-Based on the manual inspection, we identify two key user-related factors driving the decrease as follows.
-Figure 23. Examples of reviews discussing Content Quality at different time periods.
-Observation 3.3: Over time, the Content Quality reviews become less tolerant of flaws and more likely to penalize them.
-Excitement for emerging Gen-AI technologies appears to influence early adopters’ evaluation criteria, making them less likely to penalize imperfections.
-In our analysis of the 2022-H1 sample, 96% of the reviews are positive, with only 5% of them including constructive feedback despite assigning a perfect 5-star rating.
-For example, the user writing review 1 as shown in Figure 23 awards the app the highest rating while reporting that the AI generated low-quality facial images.
-Such reviews reflect a pattern in which users prioritize innovation and novelty over technical quality, demonstrating a higher tolerance for flaws during the early adoption phase.
-Over time, users become less tolerant of imperfections.
-The percentage of positive reviews declines from 96% in 2022-H1 to 59% in 2024-H2, indicating a significant shift in user ratings.
-Moreover, positive reviews become increasingly critical: in 2022-H2, 12% of all reviews offer constructive feedback while still assigning a high rating (4 or 5 stars), rising to 18% in 2024-H2.
-This trend suggests that users’ expectations have heightened.
-More notably, an increasing number of users begin deducting rating star points for flaws over time.
-In 2022-H2, only 2% of constructive reviews assign a 5-star rating, while 10% assign a 4-star rating.
-By 2024-H2, this trend intensifies: 15% of constructive reviewers assign a 4-star rating, whereas only 4% assign a 5-star rating.
-For instance, review 2 in Figure 23 expresses the need for improvement in graphics and gives a 4-star rating.
-This shift presents a growing criticality among users.
-While early reviews tend to overlook shortcomings in favour of innovation, later reviews are more discerning, penalizing points even when expressing generally positive sentiment.
-Observation 3.4: Over time, the Content Quality reviews are more experimental and demanding.
-Initial reviews frequently convey surprise and satisfaction with relatively modest capabilities, indicating limited expectations during the early adoption phase.
-For example, review 3 in Figure 23 from 2022-H1 expresses surprise at the high quality of the content, despite initially having low expectations.
-Such statements suggest that early positive ratings are often driven by low expectations of the novel technology rather than its actual performance.
-As Gen-AI technologies become more widespread, user expectations rise accordingly.
-Reviews begin to reflect more ambitious and exploratory usage patterns, with users increasingly testing the system’s capabilities through complex prompts.
-When the app fails to meet these heightened demands, users express dissatisfaction.
-For example, review 4 in Figure 23 in 2022-H2, reports success in simple requests but horrific results with complex requests, ultimately giving the app a poor 1-star rating.
-This evolution in review content presents a shift from initial awe to critical evaluation, as users become more experimental and less tolerant of limitations.
-We refer to users who review the app in 2022 as early adopters because 2022 is the year ChatGPT was launched (Akhtar_2024), and we refer to users in 2023 and 2024 as mainstream users as they started using the app as it rose in popularity.
-Early adopters are typically motivated by innovation and are more tolerant of uncertainties and imperfections in emerging technologies than mainstream users (Paulino2024).
-Hence, the decline in the average rating of Content Quality can be explained by the change in observed characteristics between early adopters and mainstream users.
-Observation 3.5: The AI Performance topic exhibits a stable trend in average user ratings over time, characterized by a slight increase in 2023-H1, and followed by a modest decline in 2024-H1.
-To account for this overall stability, we analyze fluctuations at the app-category level.
-We identify a compensatory pattern revealing that improvements in user-perceived AI Performance in some categories are offset by declines in others, producing the flat global average.
-However, this apparent stability conceals divergent user experiences.
-Categories such as Education and Games show clear improvement, reflecting that users in these domains perceive AI as directly enhancing learning and task performance.
-In contrast, most other categories (e.g., Productivity, Entertainment, Books & Reference) show gradual declines.
-This may suggest that developers in these areas are integrating AI features prematurely, before the technology is fully optimized for their use cases.
-When users accustomed to fast and accurate AI systems encounter slower or less reliable integrations, their satisfaction drops.
-These declines therefore point to gaps in seamless integration and contextual adaptation rather than outright failure of AI Performance.
-For developers, this indicates that AI performance has become a key competitive factor.
-Users now expect fast, accurate, and context-aware AI behavior.
-When AI integrations fall short or introduce lag, inconsistency, or irrelevant responses, users quickly express dissatisfaction.
-Sustaining trust therefore depends on ensuring that AI features operate seamlessly within the app’s core experience rather than as add-ons.
-Figure 24. Examples of reviews discussing AI Performance at different time periods.
-Observation 3.6: Advances in image generation capabilities contribute to the slight increase in the average rating of the AI Performance topic in 2023-H1.
-In 2022-H1, negative reviews of image generation apps constitute 9% of reviews in our sample and account for 47% of all negative reviews during that period.
-These users express frustration with the visual output.
-For example, review 1 in Figure 24 expresses frustration over the performance of the image generator.
-By 2023-H1, these issues appear to be addressed, as the percentage of negative reviews of image generation apps drops to 5%, and their share of total negative reviews decreases to 29%.
-This reduction in dissatisfaction likely suggests improvements in the AI Performance of image generation, which helps lift the overall AI Performance rating during this period.
-Observation 3.7: Underperformance in Productivity apps leads to a slight decline in the average rating of the AI Performance topic in 2024.
-By 2024-H1, Productivity apps emerge as the leading source of performance-related complaints, accounting for 8% of all reviews in our sample and 44% of all negative reviews.
-Unlike creative tools, which are often evaluated in terms of novelty or expressiveness, Productivity apps are frequently judged against well-established, non-AI alternatives that offer straightforward functionality.
-Consequently, users express notable frustration with AI-driven features that are perceived as unnecessarily complex or unreliable.
-For example, review 2 in Figure 24 prefers a regular assistant over an AI assistant because the AI assistant complicates simple tasks, and review 3 in Figure 24 prefers a calculator over an AI due to accuracy issues.
-These comments highlight that Gen-AI integration can complicate tasks that would otherwise be simple.
-Rather than enhancing the user experience, AI features may undermine usability if they are poorly aligned with user expectations or fail to deliver practical value.
-Some reviews explicitly question the necessity of AI features.
-Review 2 in Figure 24 describes abandoning their AI assistant after it offered a detailed classification of fan technologies instead of simply turning on a fan.
-This indicates that forced integration may backfire, especially when it disrupts core functionality.
-Observation 3.8: The stability in the average rating of the AI Performance topic is driven by offsetting trends across app types.
-Some app categories show improvement (e.g., image generation), while others show decline (e.g., Productivity).
-These shifts tend to cancel each other out at the global level.
-In 2023, rising satisfaction with creative apps contributes to a mild increase in performance ratings.
-In 2024, dissatisfaction in Productivity apps is balanced by relatively stable perceptions in creative and educational apps.
-This compensatory effect explains the flat trend in average performance ratings over time.
-Although the global average for AI Performance ratings remains stable over time, this apparent stability masks meaningful variation across app categories.
-Our analysis reveals divergent trends: categories such as Art & Design, Entertainment, and Books & Reference show moderate declines in average AI Performance ratings (–0.84 to –0.93), while others like Education and Games experience notable increases (+1.39 and +0.96, respectively).
-These opposing shifts suggest that stability at the aggregate level may result from the balancing of improvements and declines across distinct domains.
-This trend may also reflect differences in user expectations.
-Creative tools are often evaluated for imagination and output quality, while Productivity apps are judged more on efficiency and reliability.
-$\Rightarrow$ Implication 7.
-Developers entering a competitive Gen-AI market must ensure that their apps are robust, possess refined capabilities and are capable of meeting the higher expectations of mainstream users.
-Over-promising performance can lead to user dissatisfaction and reputational harm, whereas transparent communication about capabilities and limitations can foster trust.
-AI integration should be purpose-driven and context-aware.
-Raji et al.
-(Raji_Kumar_Horowitz_Selbst_2022) discussed how AI systems are often deployed without sufficient consideration of their actual functionality and utility.
-The authors argue that the assumption of AI’s inherent value can lead to ineffective or even harmful implementations, stressing the need for critical assessment before integration.
-Our findings further compel developers to critically assess whether Gen-AI meaningfully enhances a given task, rather than assuming its value by default.
-Inappropriate or unnecessary use of AI can degrade usability and frustrate users.
-Observation 3.9: The average user ratings for the Content Policy & Censorship topic follow an increasing trend, with fluctuations in 2022-H2 and 2024-H1.
-To understand this trend, we examine the raised user concerns in three random samples of user reviews in 2022-H2, 2023-H2, and 2024-H1.
-We find that content policy complaints discuss three main concerns: 1) ethical AI training practices, 2) the effectiveness of content filters, and 3) false-positive censorship.
-This pattern reflects how content policy concerns were especially prominent in the early stages of Gen-AI deployment, when moderation systems were still limited and users frequently encountered unsafe or poorly filtered outputs.
-Over time, developers appear to have become increasingly aware of these issues and implemented more effective safeguards, leading to gradual improvement in user satisfaction.
-The upward trajectory in ratings thus signals a maturing ecosystem where content safety and moderation mechanisms are continuously refined.
-For developers, this emphasizes the importance of iterative moderation design.
-Continuous refinement of content filters not only improves safety but also builds long-term user trust.
-Figure 25. Examples of reviews discussing Content Policy at different time periods.
-Observation 3.10: Ethical concerns around AI-generated art are more prominent in early samples but disappear in later ones.
-In 2022-H2, 10% of sampled reviews express concerns about the ethics of training AI on copyrighted human art without consent or compensation.
-For example, review 1 in Figure 25 expresses how Gen-AI harms artists whose art is stolen.
-This concern disappears from the random samples of 2023-H2 and 2024-H1.
-This highlights the importance of transparent and ethical data use policies.
-However, the absence of similar comments in later samples suggests either that the user-base shifted or that these concerns were deprioritized in users’ feedback.
-The drop may also indicate user desensitization or limited platform changes aimed at addressing the issue.
-Nevertheless, the ethics of AI content policy should still be addressed by researchers and developers.
-Observation 3.11: Content filters improve over time in blocking unwanted adult content (i.e., the raised concerns drop from 12% to 3%) and illegal content (i.e., the raised concerns drop from 4% to 0%).
-In our early sample, 15% of the studied reviews flag harmful content slipping past moderation, especially involving sexual or predatory material.
-This drops to 5% in 2023-H2 and to 3% by 2024-H1, indicating increased user satisfaction with content safety mechanisms.
-Examples of such reviews can be found in reviews 2, 3, and 4 in Figure 25.
-The proportion of users reporting the unintentional generation of adult content drops from 12% in 2022-H2 to 3% in 2024-H1.
-Reports of illegal or predatory content involving children drops from 4% to 0%.
-These declines reflect measurable improvement in censorship mechanisms for harmful content.
-Observation 3.12: Complaints about false-positive censorship grow over time (from 35% to 59%), especially among users engaging in regular and appropriate use.
-As filters improve at blocking harmful content with time, users increasingly report unnecessary censorship during benign interactions.
-In 2022-H2, 35% of users in our sample request a less strict filter.
-This jumps to 71% in 2023-H2 and settles at 59% in 2024-H1.
-While the proportion of users requesting more freedom to generate explicit or violent content decreases slightly (from 21% in 2022-H2 to 14% in 2024-H1), more users begin to complain about filters interfering with regular usage.
-For example, 6% of reviews in 2022-H2 flag such issues, rising to 30% in 2024-H1.
-Review 5 in Figure 25 shows an example of reviews discussing false-positive censorship, where the user reports receiving a guidelines warning for writing a non-explicit role-play scene.
-This suggests that while filters are more successful in removing harmful content, they are over-correcting and obstructing benign, creative, or functional uses.
-Over-correction can hinder the Gen-AI user experience, particularly in role-play and storytelling apps.
-Observation 3.13: Despite an overall positive trend, over-correction in censorship filters causes a minor ratings setback in 2024.
-The average rating for Content Policy and Censorship follows an increasing trend over time, reflecting growing user trust in content moderation systems.
-As previously discussed, this trend is supported by a substantial decline in complaints about harmful content.
-Correspondingly, the share of positive reviews that discuss censorship rises from 4% in 2022-H2 to 51% in 2023-H2.
-This rise indicates that while many users enjoy the app enough to leave a high rating, they still express frustration with the limitations imposed by the censorship filters.
-However, this progress introduces a new challenge of false-positive filtering in 2024-H1.
-As previously discussed, while moderation systems are successful in blocking inappropriate content, they begin flagging benign prompts as violations.
-This over-correction results in a moderate decline in the share of positive censorship-related reviews (20% in 2024-H1) and contributes to a slight drop in the average rating.
-Nevertheless, this dip does not reverse the overall gains.
-Ratings in 2024-H1 remain higher than in 2022-H2, suggesting that while usability concerns emerge, users still view the improvements in content safety as a net positive.
-This also highlights an important reality for developers; progress in content moderation is unlikely to be perfectly linear.
-Minor setbacks and fluctuations are to be expected as systems evolve and adapt to user needs.
-$\Rightarrow$ Implication 8.
-Developers must strike a careful balance between safety and creative freedom.
-While stricter filters reduce the occurrence of disturbing or illegal content, they increasingly disrupt legitimate use.
-To avoid over-correction, developers should implement modular moderation systems with customizable filters or toggles that can be adjusted by personal preference.
-Clear content policy guidelines and transparent communication about AI behaviour are also essential.
-Additionally, ethical considerations around AI training data should not be ignored by developers or researchers, even if fewer users express concerns.
-These considerations remain essential for maintaining long-term trust and ensuring platform responsibility.
-Summary of RQ 3 Our analysis reveals that Gen-AI topics evolve differently over time with increasing, stable, or decreasing trends.
-Content Quality ratings steadily decline, not due to technological regression, but due to increased user expectations.
-In contrast, AI Performance shows a stable rating trend, driven by a balancing effect between improving and declining domains.
-Lastly, Content Policy & Censorship follows an overall positive trend, with improved content filters increasing user ratings.
-These findings reflect a maturing user base that is more experienced, discerning, and vocal.
-Our user base of Gen-AI apps raise the bar for what constitutes a high-quality Gen-AI experience.
-7. Discussion
-In this section, we reflect on the broader implications of our findings and provide targeted recommendations for six key stakeholder groups: 1) users, 2) platform owners, 3) developers, 4) testing tool designers, 5) requirements engineers, and 6) policymakers and regulators.
+Summary of Developer Reply Developer responses are largely generic and template-based, with limited substantive issue resolution and frequent redirection to external support channels.
+Responsiveness varies by category and topic, prioritizing actionable issues, and is lower for Gen-AI reviews than for non-Gen-AI reviews despite similar response times.
+8. Implications
+In this section, we reflect on the broader implications of our findings and provide targeted recommendations for six key stakeholder groups: users, platform owners, developers, testing tool designers, requirements engineers, and policymakers and regulators.
 8.1. Implications for Users
-Our analysis reveals that users are not passive recipients of Gen-AI technology but are increasingly engaged, expressive, and collaborative (Observation 3.3).
-Reviewers provide constructive feedback to co-create better experiences with developers.
-As Gen-AI becomes more embedded in users’ lives, particularly in the emotional or educational contexts, users come to expect empathy, personalization, and reliability (Implications 1, 3, and 5).
-This shift highlights the importance of treating users not merely as consumers, but as stakeholders who help shape and refine Gen-AI apps.
-Furthermore, our findings highlight the need for further research into the evolving relationship between users and Gen-AI conversational agents.
-Users increasingly describe these systems as trusted companions or emotional supports.
-Understanding the psychological and social bonds users form will be crucial for guiding the ethical design, deployment, and regulation of AI in emotionally sensitive contexts (Implication 5).
+Our findings suggest that users are not passive recipients of Gen-AI technology but active participants who shape system behavior through feedback and interaction.
+This evolving role highlights several key actions for users:
+Engage responsibly with AI as both a tool and an emotional companion.
+Gen-AI systems increasingly support accessibility, wellbeing, and companionship (Opportunity 1), but our findings also show that user relationships can evolve toward emotional dependence and addiction (Finding 3.4).
+While these systems can provide meaningful support, users should remain aware of the risks of over-reliance and recognize the limitations of AI as a substitute for human relationships.
+Maintaining balanced usage and critical awareness is essential, particularly in emotionally sensitive contexts.
+Develop informed expectations of AI capabilities and limitations.
+Users should be aware that Gen-AI systems may produce incorrect, inconsistent, or unexpected outputs, particularly for complex tasks.
+As highlighted in earlier findings, many frustrations stem from mismatched expectations of system capabilities.
+Understanding these limitations can help users interact more effectively with AI systems and interpret outputs more critically (Challenge 1).
+Be mindful of content moderation policies and platform differences.
+Content moderation practices vary significantly across Gen-AI apps and can affect the type of content users are able to generate.
+Perceived bias or restrictions may result from underlying safety mechanisms.
+Being aware of these differences enables users to make informed choices about which platforms align with their expectations regarding bias, privacy, freedom of expression, and content access (Challenge 2 and Finding 3.5).
 8.2. Implications for Platform Owners
-Platform owners must bear responsibility for ensuring that Gen-AI apps comply with ethical standards (Observation 3.10).
-As AI capabilities evolve rapidly, app store operators should implement both technical tools and regulatory frameworks that evaluate whether submitted Gen-AI apps meet baseline criteria for safety, transparency, and responsible data usage.
-This may include requiring apps to disclose whether AI is used, clarifying how outputs are generated, and verifying that training data sources are ethically obtained.
-By introducing certification processes or audit mechanisms for Gen-AI features (particularly in apps that claim educational, therapeutic, or professional utility), platforms can promote accountability and mitigate harm, fostering greater trust among users and developers alike.
-App platforms play a critical role in shaping the quality of user feedback.
-Our results show that a large proportion of user reviews are non-informative (Observation 1.2), which limits their value for developers and researchers.
-Future research could explore how redesigning review interfaces might encourage more meaningful feedback.
-One possible approach is to prompt users to answer specific questions or rate distinct topics, such as usability, design, or content accuracy, separately from the overall rating.
-Beyond improving review quality, our findings also reveal structural limitations in how app stores currently support developer–user interaction.
-As shown in our developer response analysis (Implication 6), review threads in Gen-AI apps function primarily as an initial point of contact rather than a space for resolving AI-specific failures.
-Many issues require diagnostic information that cannot be shared publicly or involve model- and policy-level changes that cannot be deployed quickly.
-Platform owners can play an important role by offering richer, privacy-preserving support channels that complement review sections and allow developers to investigate Gen-AI issues more effectively.
-Strengthening these communication pathways would improve user experience and reduce the friction developers face when addressing complex AI behaviours.
+Our findings highlight the critical role of platform owners in ensuring the safe deployment of Gen-AI apps, improving feedback quality, and supporting effective developer-user interaction.
+Enforce transparency and accountability in Gen-AI systems.
+Platform owners should ensure that Gen-AI apps comply with ethical and safety standards as capabilities evolve.
+This includes requiring developers to disclose the use of AI, explain how outputs are generated, and verify that training data sources are ethically obtained (Challenge 2, Finding 3.5).
+App stores could further support this by providing a certification badge indicating that a Gen-AI app meets AI regulatory standards, including safeguards against over-reliance, addiction, and inappropriate content (Opportunity 1, Finding 3.4).
+Introducing certification processes or audit mechanisms is important, especially for apps involving education, wellbeing, or young users, as they can improve accountability and reduce potential harm.
+Improve the quality and structure of user feedback.
+Our results show that a large proportion of user reviews are non-informative (Finding 1.2), limiting their usefulness for developers and researchers.
+Platform owners can address this by redesigning review interfaces to encourage more structured and meaningful input.
+For example, prompting users to evaluate specific aspects such as usability, content quality, or accuracy, rather than relying solely on free-text reviews, can lead to more actionable feedback.
+Enhance support channels for complex Gen-AI issues.
+Our developer response analysis reveals that review threads often serve only as an initial point of contact rather than a space for resolving AI-specific issues.
+Many Gen-AI-related problems require additional context (e.g., prompts, outputs, system logs) or involve model- and policy-level changes that cannot be addressed through public replies.
+Platform owners should provide richer, privacy-preserving support channels that enable more effective communication between users and developers.
+Strengthening these mechanisms can reduce friction in addressing complex AI behaviors and improve overall user experience.
 8.3. Implications for Developers
-Developers face growing pressure to meet user expectations that are increasingly shaped by leading models, such as ChatGPT (Implication 4).
-Users demand fast, accurate, and context-aware outputs across a wide range of use cases, including education, professional assistance, and emotional support (Implications 3 and 5).
-Our findings highlight several critical areas for improvement: fine-tuning censorship filters to avoid false positives (Observation 3.9; Implication 2), supporting accessibility and inclusivity concerns (Implication 5), and enabling more personalized and flexible user experiences (Implication 3).
-Developers should also consider offering modular features and customization options, such as adjustable censorship settings, adjustable AI functionalities, and adaptive content formats.
-Importantly, developers must think critically about when and how AI is used, ensuring that integration adds meaningful value, as not every problem or task requires an AI solution (Observation 3.7; Implication 7).
+Our findings highlight increasing pressure on developers to meet evolving user expectations shaped by leading Gen-AI apps, while ensuring that AI integration remains meaningful, reliable, and aligned with user needs.
+Design for collaborative and evolving user needs.
+Users increasingly engage with Gen-AI systems as collaborative partners that support creativity, ideation, and problem-solving (Opportunity 2).
+Rather than passively consuming outputs, users iteratively refine prompts and incorporate AI suggestions into their workflows.
+At the same time, many use cases emerge organically through user experimentation rather than predefined developer intent (Opportunity 3).
+Developers should therefore monitor how users adapt AI systems in practice and design flexible, extensible systems that support co-creation, iterative interaction, and evolving use cases.
+For example, as users increasingly engage with AI, they begin to form emotional bonds; developers must move beyond optimizing conversational quality and towards designing responsible, safe emotional AI systems.
+This includes incorporating safeguards such as usage awareness features, intervention mechanisms for excessive use, and improved moderation of harmful conversational behaviors, particularly for users relying on AI for emotional support (Finding 3.4).
+Ensure purposeful and value-driven integration of Gen-AI.
+As user expectations are shaped by powerful systems such as ChatGPT, developers face growing pressure to deliver fast, accurate, and context-aware outputs across diverse domains.
+However, our findings show that not all features benefit from AI integration (Challenge 3).
+Developers must carefully evaluate when Gen-AI meaningfully enhances the user experience and when simpler solutions are more appropriate.
+Effective integration requires balancing innovation with reliability, ensuring that AI capabilities provide clear value rather than introducing unnecessary complexity.
+Balance moderation and user expectations.
+Developers must address challenges related to managing user expectations and system limitations (Challenge 1) and balancing content moderation with creative freedom (Challenge 2).
+Developers should provide clearer feedback on system limitations and offer customization options that support diverse user needs.
+Content policy should also incorporate greater transparency in data usage and policy communication.
+Developers should clearly communicate how user data is handled and how moderation decisions are made, helping users better understand system behavior and build trust (Finding 3.5).
 8.4. Implications for Testing Tool Designers
-Our LLM-based framework achieves 91% accuracy in topic extraction and assignment (Observation 1.4), demonstrating the effectiveness of prompt design and data pre-filtering (Observations 1.1 and 1.3).
+Our LLM-based framework achieves 91% accuracy in topic extraction and assignment (Finding 1.4), demonstrating the effectiveness of prompt design and data filtering.
 These results point to new opportunities in automated review analysis, particularly in testing and quality assurance workflows.
 Designers of testing tools should consider incorporating LLMs for top topic extraction and assignment.
-Our methods can help uncover nuanced user feedback across languages, accessibility contexts, and time, significantly enhancing product refinement cycles.
+Our methods can help uncover nuanced user feedback across languages, accessibility contexts, and time, significantly enhancing product refinement cycles (Finding 2.3).
 8.5. Implications for Requirements Engineers
-Requirements engineers can benefit from the structured insights our topic analysis framework provides.
-The ability to track how user expectations evolve by topic over time can inform roadmap planning and feature prioritization (Observation 3.3).
-Our analysis also reveals that users clearly articulate feature-specific frustrations and preferences, providing actionable guidance for interface design, feature retention, and refinement.
-Engineers should leverage such review data to validate or revise assumptions about user needs, especially in light of shifting demographics and rising expectations for Gen-AI-powered functionality.
-Beyond Gen-AI apps, the same framework and analytical tools can be applied to user reviews in other domains, enabling engineers to systematically identify evolving requirements and usability concerns across a wide range of software systems.
+Our findings suggest extending requirements engineering processes to capture emerging quality attributes and design considerations specific to Gen-AI systems.
+In particular, the identified opportunities highlight the need to explicitly model user-centered and interaction-driven requirements.
+For example, AI for Accessibility and Wellbeing (Opportunity 1) requires engineers to account for the emotional impact of Gen-AI systems and incorporate safeguards that mitigate risks such as dependency and user vulnerability.
+Similarly, AI as a Collaborative Creative Tool (Opportunity 2) calls for requirements that support co-creative interaction, treating users as active collaborators rather than passive consumers.
+The diversity of user-driven applications observed in AI Versatility (Opportunity 3) further emphasizes the need for flexible and adaptive requirements that accommodate evolving and unanticipated use cases.
+At the same time, the identified challenges point to critical gaps in current requirements practices.
+Managing User Expectations and AI Limitations (Challenge 1) highlights the importance of specifying realistic system capabilities and continuously validating them against user expectations.
+Balancing Content Moderation and Creative Freedom (Challenge 2) introduces the need to formalize requirements around moderation policies, data safety, and privacy while preserving user autonomy.
+Finally, Strategic Integration of Gen-AI Features (Challenge 3) suggests that requirements engineers must critically evaluate when Gen-AI functionality is appropriate, ensuring that AI is integrated where it provides clear user value rather than as a default solution.
+While our study focuses on Gen-AI apps, the SARA framework may be adapted to other domains to support requirements analysis using user-generated data.
 8.6. Implications for Policy Makers and Regulators
-The emergence of Gen-AI apps raises important questions about ethics, safety, and user protection.
-While some early users voiced concerns about data ethics and copyright in training data, these concerns faded in later samples (Observation 3.10), which may indicate normalization rather than resolution.
-Policy makers must ensure that ethical standards around content ownership, training transparency, and accountability are formalized and enforced (Implication 2).
-Furthermore, content moderation must strike a balance between safety and usability.
-Regulators should develop protocols to mitigate the harm caused by unregulated AI generations (Observation 3.11; Implication 2), particularly as Gen-AI becomes increasingly pervasive in sensitive domains such as education and emotional support.
+The emergence of Gen-AI apps raises important questions about ethics, safety, and user protection (Finding 3.5).
+While some early users voiced concerns about data ethics and copyright in training data, these concerns faded in later samples, which may indicate normalization rather than resolution.
+Policy makers must ensure that ethical standards around content ownership, training transparency, and accountability are formalized and enforced.
+Furthermore, content moderation must strike a balance between safety and usability (Challenge 2).
+Regulators should develop protocols to mitigate the harm caused by unregulated AI generations, particularly as Gen-AI becomes increasingly pervasive in sensitive domains such as education and emotional support (Opportunity 1, and Finding 3.4)
 9. Threats to Validity
@@
 Second, our methodology relies on LLMs to automatically filter and classify reviews.
-To control for potential biases introduced by the LLM-based automation, we manually annotate samples and our results demonstrate high agreement, i.e., 91% accuracy, between the LLM-generated classifications and human annotations, supporting the reliability of using LLMs for these tasks.
+To control for potential biases introduced by the LLM-based automation, we manually annotate samples, and our results demonstrate high agreement, i.e., 91% accuracy, between the LLM-generated classifications and human annotations, supporting the reliability of using LLMs for these tasks.
 Additionally, we measure inter-rater reliability to confirm the consistency and trustworthiness of the automated classifications.
 Internal threats to validity concern the soundness of the procedures used to collect, analyze, and interpret the data.
-A key internal threat relates to the uncertainty surrounding the timeline for introducing the Gen-AI features app.
-While we manually verify that all selected apps currently support Gen-AI functionality, we cannot determine the exact date of integration.
-To reduce the risk of analyzing reviews unrelated to Gen-AI, we exclude all reviews prior to 2022, i.e., the year ChatGPT was released.
-While this filtering improves temporal alignment, it cannot fully guarantee that all remaining reviews pertain to post-Gen-AI functionality.
+A potential threat arises from uncertainty in identifying the exact timeline of Gen-AI feature integration for each app.
+To mitigate this, we adopt a conservative, evidence-based approach by manually examining historical release notes and online sources to determine the earliest documented integration date.
+While this approach may miss undocumented or informally introduced features, it ensures that all included data reflects confirmed Gen-AI functionality.
+As such, our integration dates represent the earliest verifiable evidence rather than the precise deployment time, prioritizing reliability and reducing the risk of including pre-Gen-AI reviews.
+Additionally, a potential threat arises from ambiguity in determining whether an app genuinely incorporates Gen-AI functionality.
+In particular, some apps reference “AI” without providing generative capabilities, which may lead to misclassification.
+To mitigate this risk, we adopt a conservative coding strategy in which a mere mention of AI (e.g., “AI Song”) is not considered sufficient evidence of Gen-AI integration unless it is explicitly supported by documented generative functionality (e.g., “AI Song Generator”).
+This approach prioritizes precision in identifying Gen-AI apps, reducing the likelihood of including apps that do not provide true generative capabilities.
 External threats to validity concern the ability to generalize the results.
-Our analysis focuses on publicly available user reviews from the Google Play Store, which may not fully reflect user experiences on other platforms, such as the Apple App Store.
-This single-platform scope may limit the generalizability of results to other ecosystems with different user bases or review practices.
-We take several steps to enhance the generalizability of our results.
-First, we include a diverse set of apps across multiple categories to ensure a wide range of topics and app functionalities.
-Second, although the selected apps may vary in implementation details from one platform to another, many of the Gen-AI topics discussed (i.e., Content Quality, Content Policy & Censorship, and Accessibility & Inclusivity) are prevalent across platforms and commonly adopted by competing apps within and across categories.
-Consequently, the insights derived from our analysis of user perceptions and feedback are likely to be transferable to other apps that implement similar Gen-AI capabilities.
-Developers of comparable apps, including those not represented in our dataset or operating on different platforms, may benefit from these findings.
-In particular, the results can help them better understand user concerns, inform feature design decisions, and mitigate risks associated with Gen-AI integration.
+Our dataset is constructed using systematic keyword-based searches on the Google Play Store followed by manual validation.
+Google Play Store search results are capped, which prevents enumerating all Gen-AI apps for a given keyword.
+Accordingly, our goal is representativeness rather than exhaustiveness.
+By querying multiple taxonomy-derived keywords and retaining only apps with verified generative functionality, we obtain a diverse sample covering a wide range of Gen-AI capabilities and usage contexts.
+While some apps may not be included, our findings are expected to generalize to apps implementing similar Gen-AI functionalities and design choices.
+Additionally, user perceptions and behaviors may differ across platforms.
+To mitigate this, we conduct a cross-platform comparison between Google Play Store and App Store reviews, observing consistent topic patterns across platforms.
+While this strengthens the generalizability of our findings, differences in platform policies, user demographics, and review mechanisms may still influence the results.
 10. Conclusion
-The rapid integration of Gen-AI features into mobile apps has transformed how users interact with technology.
-In this study, we present a user-centered analysis of Gen-AI mobile apps by examining over 676,066 reviews from 173 apps on the Google Play Store.
-Using our proposed four-phase SARA methodology and prompt-based LLM techniques, we uncover actionable insights into how users experience and evaluate Gen-AI functionalities.
-Our results show that LLMs can reliably extract user-discussed topics with high accuracy, i.e., 91%, when supported by five-shot prompting and data cleaning.
-We identify the 10 most frequently discussed topics revealing user behavior and challenges that inform actionable insights for developers and researchers.
-Furthermore, our findings suggest an evolution in user expectations, with a shift in user behavior, such as shifting from early adopters who were more forgiving of limitations to a mainstream user base that demands higher expectations.
-Future work will explore cross-platform comparisons and further investigate how expectations vary across different user profiles (e.g., casual vs.
-expert users) as Gen-AI tools continue and mature.
-Data Availability. We provide a replication package, including manually labeled data, and Python scripts to replicate the analyses at omitted.
+Gen-AI is rapidly transforming mobile apps, yet understanding how users perceive and interact with these systems remains limited.
+In this paper, we present SARA, a framework for systematically analyzing user reviews of Gen-AI apps, and apply it to over one million reviews across multiple app categories.
+Our findings reveal that user discussions of Gen-AI apps surface key opportunities (i.e., AI for Accessibility and Wellbeing, AI as a Collaborative Creative Tool, and AI Versatility) alongside challenges (i.e., Managing User Expectations and AI Limitations, Balancing Content Moderation and Creative Freedom, and Strategic Integration of Gen-AI Features).
+Together, these insights inform implications across stakeholders: encouraging users towards more informed and safe interactions; guiding platform owners in enforcing transparency and accountability in Gen-AI systems; helping developers in developing Gen-AI systems for collaborative and evolving user needs; inspiring testing tool designers to incorporate LLMs in automated review analysis; enabling requirements engineers to capture emerging socio-technical requirements; and informing policymakers and regulators in addressing risks and safeguards.
+We further observe that user engagement evolves over time, from initial novelty-driven interactions to more mature use, where users increasingly identify risks such as dependency and emotional reliance, while expressing reduced concern for issues such as data policy and privacy.
+This work contributes both a scalable analysis framework and empirical insights into user perceptions of Gen-AI systems.
+These findings inform the design of more effective and user-centered Gen-AI apps and highlight the need to incorporate socio-technical considerations into requirements engineering processes.
+Future work can build on these findings by examining human-AI collaboration in greater depth, particularly focusing on how users interact with Gen-AI systems as co-creative partners and how such interactions shape user expectations, outcomes, and long-term engagement.
+Data Availability. We provide a replication package, including manually labeled data, and Python scripts to replicate the analyses at https://github.com/Safwat-UofT/Gen-AI_User_Reviews.
 Acknowledgements.
 We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [RGPIN-2021-03969].
+Appendix A Gen-AI App Categories and Keywords
+We derive our keyword set from the Gen-AI taxonomy proposed by Gozalo-Brizuela and Garrido-Merchán (Gozalo-Brizuela and Merchan, 2024).
+We exclude categories that are not relevant to mobile applications (i.e., biotech, brain, drug discovery, and other), and adapt the remaining categories into keyword-based queries for app store search.
+Table 11 presents the resulting categories and keywords used in our data collection process.
+Some keywords (e.g.,“Image Editing” and “Business”) are too broad, so we append “Generative AI” for all keywords when searching the Google Play Store.
+Table 11. Categories and keywords for app search in the Google Play Store
+Category Keywords
+Text Conversational AI, Text-to-Science, Text-to-Author Simulation, Text-to-Medical Advice, Text-to-Itinerary, Doc-to-Text
+Image Image Editing, Artistic Images, Realistic Images
+Video Text-to-Video, Video Production, Image-to-Video
+3D Text-to-3D
+Code and Software Text-to-Code, Text-to-Multilingual Code
+Speech Text-to-Speech, AI Understanding, Speech-to-Text
+Business Business
+Marketing Marketing, Advertisement, Accounting
+Gaming Video game Creation
+Music Music Generator
+Appendix B Prompts
+B.1. Filtering Prompt
+Figure 11. Template of the filtering prompt used to filter out non-informative reviews.
+B.2. Topic Extraction Prompt
+Figure 12.
+Template of the topic extraction prompts for (E1: 0-shot, E2: 3-shot, and E3: 5-shot) used to extract the top topics from a large sample of stage $x$ and app category $y$.
+“Insert category name” is where we insert the name of the app category $y$.
+“Insert sample of reviews” is where we list our large sample of reviews, separated by a new line escape.
+B.3. Topic Assignment Prompt
+Figure 13.
+Template of the P5: topic assignment prompt used to assign topics to a list of reviews.
+We insert the set of top topics $T(x,y,n)$, five few-shot examples, and the small sample $S_{\text{small}}(x,y)$.
+Appendix C Topic Categories
+Table 12.
+Summary of Gen-AI topic categories, including the Gen-AI Reviews Count (GRC), the average Gen-AI rating (AGR), the number of apps, and the number of app categories in which each topic category appears (Full dataset considered).
+Gen-AI Topic Category GRC AGR # Apps # App Categories
+AI Performance 158,873 (15%) 4.2 142 9
+Utility & Use Cases 118,405 (11%) 4.8 58 4
+Content Quality 113,714 (11%) 3.7 142 8
+Creative Potential 87,383 (8%) 4.8 96 5
+Content Policy & Censorship 22,371 (2%) 2.4 25 4
+Features & Functionality 20,200 (2%) 3.6 34 4
+Emotional Connection 14,652 (1%) 4.8 9 1
+Comparison to Other Apps 1,225 (0.1%) 4.3 17 2
+Table 13.
+Summary of non-Gen-AI topic categories, including the non-Gen-AI Review Count (NRC), the average non-Gen-AI rating (ANR), the number of apps, and the number of app categories in which each topic category appears (Full dataset considered).
+Non-Gen-AI Topic Category NRC ANR # Apps # App Categories
+Monetization Methods & Structure 149,255 (14%) 2.2 160 9
+Technical Difficulties 126,377 (12%) 2.1 167 9
+Other 111,297 (11%) 3.7 162 9
+User Interface & Experience 58,885 (6%) 4.3 145 9
+Updates & Evolution 35,489 (3%) 2.6 101 8
+Customer Support 9,752 (1%) 2.1 116 6
+Features & Functionality 2,387 (0.2%) 4.3 12 2